### v1

# Parsimonious Compactly Supported Covariance Models in the Gauss Hypergeometric Class: Identifiability, Reparameterizations, and Asymptotic Properties

Abstract
We study the covariance model belonging to the Gauss hypergeometric ($\mathcal{GH}$) class, a highly flexible and compactly supported correlation model (kernel). This class includes the well-known Generalized Wendland ($\mathcal{GW}$) and Matérn ($\mathcal{MT}$) kernels as special cases. First, we provide necessary and sufficient conditions for the validity of the $\mathcal{GH}$ model. We then demonstrate that this family of models suffers from identifiability issues under both increasing and fixed-domain asymptotics.
To address this problem, we propose a parsimonious version of the model that adheres to an optimality criterion. This approach results in a new class of compactly supported kernels, which can be seen as an improvement over the $\mathcal{GW}$ model. Additionally, we show that two specific compact support reparameterizations allow us to recover the $\mathcal{MT}$ model, highlighting the advantages and disadvantages of each reparameterization.
[…]
Keywords: Fixed domain asymptotics, Gaussian random fields, Generalized Wendland model, Integral range, Sparse matrices.
1 Introduction
Gaussian processes, or Gaussian random fields (RFs), are widely used in various disciplines such as spatial statistics (Cressie,, 1993; Stein,, 1999; Banerjee et al.,, 2004), machine learning (Rasmussen and Williams,, 2006), numerical analysis (Wendland,, 2004), and computer experiments (Santner et al.,, 2003), among others. The Gaussian assumption implies that the finite-dimensional distributions of the RF are fully determined by the mean value and the covariance function. In particular, covariance functions (also known as kernels in the machine learning literature) are positive semidefinite functions that play a fundamental role in modeling relationships between data points in Gaussian RFs, kernel methods, and other machine learning algorithms.
For RFs defined in Euclidean spaces, covariance functions that depend solely on the distance between any pair of points within the domain are called isotropic or radially symmetric. This paper operates under this assumption, which greatly simplifies the discussion. However, isotropic covariances serve as the building blocks for more complex models, such as anisotropic, nonstationary (Paciorek and Schervish,, 2006), multivariate (Gneiting et al.,, 2010), or spatiotemporal extensions (Porcu et al.,, 2020), among others.
To date, there exists a rich catalog of parametric isotropic covariance functions (see, for instance, Rasmussen and Williams, (2006) or Chilès and Delfiner, (2012)), among which the Matérn ($\mathcal{MT}$ hereafter) model is by far the most widely used (see Porcu et al., (2024) for a comprehensive review). It has played a central role in spatial statistics for decades and, more recently, in other disciplines such as numerical analysis, approximation theory, and machine learning. A key advantage of the $\mathcal{MT}$ model is its continuous parameterization of the differentiability of the sample paths of the associated Gaussian RF through a positive smoothness parameter. The larger the smoothness parameter, the higher the level of differentiability of the sample paths. .
While the $\mathcal{MT}$ model remains widely used, particularly in the stochastic partial differential equations approach and its generalizations to non-Euclidean domains (Lindgren et al.,, 2011; Bolin et al.,, 2024), enhancements to the $\mathcal{MT}$ model have been proposed in recent years. These covariance models retain the benefits of the Matérn model while offering additional features (Porcu et al.,, 2024).
Among these, the generalized Wendland ($\mathcal{GW}$ hereafter) (Gneiting,, 2002; Bevilacqua et al.,, 2019) has gained popularity, as it allows for a continuous parameterization of the differentiability of the sample paths of the underlying Gaussian RF, similar to the $\mathcal{MT}$ model, while also being compactly supported. This is a desirable feature from a computational perspective, as it produces sparse covariance matrices (Duff et al.,, 1989; Davis,, 2006).
Several connections between the $\mathcal{MT}$ and $\mathcal{GW}$ models have been established. For instance, Bevilacqua et al., (2019) show that, under specific conditions, these models lead to equivalent Gaussian measures. Additionally, Bevilacqua et al., (2022) demonstrate that the $\mathcal{MT}$ model is a special limit case of a compactly supported reparameterization of the $\mathcal{GW}$ model. Estimation and prediction using $\mathcal{MT}$ and $\mathcal{GW}$ models have been extensively studied under both increasing and fixed-domain asymptotics (Mardia and Marshall,, 1984; Zhang,, 2004; Kaufman and Shaby,, 2013; Bevilacqua et al.,, 2019).
The focus of this paper is the study of the Gauss hypergeometric ($\mathcal{GH}$) covariance, a highly flexible kernel recently proposed by Emery and Alegría, (2022). These authors demonstrate that the $\mathcal{GH}$ model encompasses most well-known covariance models as special or asymptotic cases, including both the $\mathcal{MT}$ and $\mathcal{GW}$ models. In addition, they provide sufficient conditions for the validity of the model. In Porcu et al., (2024), the Gauss hypergeometric ($\mathcal{GH}$) kernel appears to be one of the most promising generalizations of the $\mathcal{MT}$ model. However, the estimation of the $\mathcal{GH}$ covariance model, particularly under fixed-domain asymptotics, has not yet been addressed.
This paper presents several contributions related to the $\mathcal{GH}$ covariance model. First, we provide necessary and sufficient conditions for the validity of the $\mathcal{GH}$ model, offering a complete characterization of the kernel. This result ensures that the model is well-defined and applicable to a wide range of scenarios.
Second, we show that this family of models suffers from identifiability issues under both increasing and fixed-domain asymptotics. These findings highlight a critical limitation of the $\mathcal{GH}$ model in practical applications, particularly when attempting to infer parameters from data. To address this problem, we propose a parsimonious version of the model that adheres to an optimality criterion. This criterion is based on the maximization of the isotropic integral range (Lantuéjoul,, 2002), which can be seen as a measure of the flexibility of the correlation model. For instance, if $d=1$, the maximum integral range corresponds to the maximum area under the unidimensional correlation function.
Using this approach, we derive a new class of compactly supported kernels, which we call the hypergeometric ($\mathcal{H}$ hereafter) model that can be seen as an improvement over the $\mathcal{GW}$ model. The model can also be seen as a generalization of the Euclid’s hat model proposed in Gneiting, (1999), which includes the celebrated spherical models widely used in geostatistics (Chilès and Delfiner,, 2012). Additionally, we show that two specific compact support reparameterizations allow us to recover the $\mathcal{MT}$ model as a special limit case, highlighting the advantages and disadvantages of each reparameterization.
Finally, for a fixed smoothness parameter, we establish strong consistency and the asymptotic distribution of the maximum likelihood (ML) estimator of the microergodic parameter associated with the $\mathcal{H}$ class.
[…]
The $\mathcal{H}$ model has been implemented in the GeoModels package (Bevilacqua et al.,, 2025) for the open-source R statistical environment.
The remainder of this paper is organized as follows. Section 2 provides background material on the $\mathcal{MT}$ and $\mathcal{GW}$ covariance models. Section 3 introduces the $\mathcal{GH}$ model and highlights the identifiability issue. Section 4 first provides necessary and sufficient conditions for the $\mathcal{GH}$ model. Then, following the criterion of maximal integral range, the $\mathcal{H}$ model is proposed. In Section 5, some closed-form expressions for the $\mathcal{H}$ model that do not depend on the computation of special functions are presented. In addition, two compactly supported reparameterizations that include the $\mathcal{MT}$ model as a special case are introduced. In Section 6, we provide the asymptotic properties of the ML estimator of the microergodic parameter. In Section 7, we report a simulation study that explores the finite-sample properties of the ML method when estimating the parameters of the proposed (reparameterized) $\mathcal{H}$ model. In Section 8, we present the analysis of a climate data set. Concluding remarks are provided in Section 9. All proofs have been deferred to Appendix A.
2 Background: the Matérn and Generalized Wendland correlation models
We denote $\{Z(\boldsymbol{s}):\boldsymbol{s}\in D\}$ a zero-mean Gaussian RF on a bounded set $D$ of $\mathbb{R}^{d}$, with stationary isotropic correlation function $C:\mathbb{R}^{d}\to\mathbb{R}$. We consider the class $\Phi_{d}$ of continuous mappings $\phi:[0,\infty)\to\mathbb{R}$ with $\sigma^{2}:=\phi(0)=1$, such that
${\rm corr}\left(Z(\boldsymbol{s}),Z(\boldsymbol{s}^{\prime})\right)=C(\boldsymbol{h})=\phi(\|\boldsymbol{h}\|),$
[…]
Schoenberg, (1938) characterized the class $\Phi_{d}$ as scale mixtures of the characteristic functions of random vectors uniformly distributed on the spherical shell of $\mathbb{R}^{d}$, with any finite and non-negative measure, $F$:
$\phi(x)=\int_{0}^{\infty}\Omega_{d}(xr)F(\text{d}r),\qquad x\geq 0,$
with $\Omega_{d}(x)=\Gamma(d/2)(2/x)^{d/2-1}J_{d/2-1}(x)$ and $J_{\nu}$ the Bessel function of the first kind of order $\nu$. The class $\Phi_{d}$ is nested, with the inclusion relation $\Phi_{1}\supset\Phi_{2}\supset\ldots\supset\Phi_{\infty}$ being strict, and where $\Phi_{\infty}:=\bigcap_{d\geq 1}\Phi_{d}$ is the class of continuous mappings $\phi$, the radial version of which is positive semidefinite on any $d$-dimensional Euclidean space.
Fourier transforms of radial versions of members of $\Phi_{d}$, for a given $d$, have a simple expression, as reported in Yaglom, (1987) or Stein, (1999). For a member $\phi$ of the class $\Phi_{d}$ such that $\phi(\|\cdot\|)$ is absolutely integrable in $\mathbb{R}^{d}$, we define its isotropic spectral density as
$\widehat{\phi}(z)=\frac{z^{1-d/2}}{(2\pi)^{\frac{d}{2}}}\int_{0}^{\infty}u^{d/2}J_{d/2-1}(uz)\phi(u){\rm d}u,\qquad z\geq 0.$ (1)
In the following we review two important examples:
  1. 1.
    The Matérn (${\cal MT}$) family
    The Matérn model is a two-parameter globally supported correlation function defined as (Matérn,, 1986):
    ${\cal MT}_{\nu,\alpha}(x)=\begin{cases}1,\quad x=0\\ \frac{2^{1-\nu}}{\Gamma(\nu)}\left(\frac{x}{\alpha}\right)^{\nu}{\cal K}_{\nu}\left(\frac{x}{\alpha}\right),\qquad x>0,\end{cases}$ (2)
    where ${\cal K}_{\nu}$ is the modified Bessel function of the second kind of order $\nu$ (Abramowitz and Stegun,, 1970, 9.6.22). The associated spectral density, $\widehat{{\cal MT}}_{\alpha,\nu}$, is given by
    $\widehat{{\cal MT}}_{\nu,\alpha}(z)=\frac{\Gamma(\nu+d/2)}{\pi^{d/2}\Gamma(\nu)}\frac{\alpha^{d}}{(1+\alpha^{2}z^{2})^{\nu+d/2}},\qquad z\geq 0.$ (3)
    Arguments in Stein, (1999) show that ${\cal MT}_{\nu,\alpha}$ belongs to $\Phi_{\infty}$ when both the smoothness parameter $\nu$ and the scale parameter $\alpha$ are strictly positive.
    The importance of the Matérn model lies in the parameter $\nu$, which controls the differentiability (in the mean square sense) of the associated Gaussian RF. Specifically, for any integer $l=0,1,\ldots$, a Gaussian RF with a Matérn covariance function is $l$-times mean square differentiable, in any direction, if and only if $\nu>l$. When $\nu=l+1/2$, the Matérn covariance function simplifies into the product of an exponential covariance and a polynomial of order $l$, given by:
    ${\cal MT}_{l+1/2,\alpha}(x)=\exp\left(-\frac{x}{\alpha}\right)\sum_{i=0}^{l}\frac{(l+i)!}{2l!}\binom{l}{i}\left(\frac{2x}{\alpha}\right)^{l-i},\quad l=0,1,\ldots$
  2. 2.
    The Generalized Wendland (${\cal GW}$) family
    The ${\cal GW}$ model is a three-parameters compactly supported correlation function defined as (Gneiting,, 2002; Bevilacqua et al.,, 2019):
    $\begin{split}{\cal GW}_{\kappa,\mu,\beta}(x)&=\frac{\Gamma(\kappa)\Gamma(2\kappa+\mu+1)}{\Gamma(2\kappa)\Gamma(\kappa+\mu+1)2^{\mu+1}}\left(1-\frac{x^{2}}{\beta^{2}}\right)^{\kappa+\mu}_{+}\\ &\times{}_{2}F_{1}\left(\frac{\mu}{2},\frac{\mu+1}{2};\kappa+\mu+1;\left(1-\frac{x^{2}}{\beta^{2}}\right)_{+}\right),\quad x\geq 0,\end{split}$ (4)
    with $\beta>0$ a compact support parameter, $\kappa>-\frac{1}{2}$ a smoothness parameter and $\mu$ a shape parameter, and where $(\cdot)_{+}$ stands for the positive part function and ${}_{2}F_{1}$ for the Gauss hypergeometric function (Abramowitz and Stegun,, 1970, 15.1.1). The associated spectral density is given by:
    $\widehat{{\cal GW}}_{\kappa,\mu,\beta}(z)=L\beta^{d}\mathstrut_{1}F_{2}\Big{(}\lambda;\lambda+\frac{\mu}{2},\lambda+\frac{\mu}{2}+\frac{1}{2};-\frac{(z\beta)^{2}}{4}\Big{)},\quad z>0,$ (5)
    where ${}_{1}F_{2}$ is a generalized hypergeometric function, $\lambda=(d+1)/2+\kappa$ and $L=\frac{2^{-d}\pi^{-\frac{d}{2}}\Gamma(\mu+2\nu+1)\Gamma(2\nu+d)\Gamma(\nu)}{\Gamma\left(\nu+{d}/{2}\right)\Gamma(\mu+2\nu+d+1)\Gamma(2\nu)}$ is a normalization constant.
    For a given $\kappa>-\frac{1}{2}$, a necessary and sufficient condition for ${\cal GW}_{\kappa,\mu,\beta}\in\Phi_{d}$ is (Bevilacqua et al.,, 2024)
    $\mu\geq\mu_{\min}(\kappa,d):=\begin{cases}\frac{\sqrt{8\kappa+9}-1}{2}\text{ if $d=1$ and $-\frac{1}{2}<\kappa<0$}\\ \kappa+\frac{d+1}{2}\text{ otherwise.}\end{cases}$ (6)
    The ${\cal GW}_{\kappa,\mu,\beta}$ model, as in the ${\cal MT}_{\nu,\alpha}$ case, allows parameterizing in a continuous fashion the mean squared differentiability of the underlying Gaussian RF and its associated sample paths through the smoothness parameter $\kappa$. Specifically, for any integer $l=1,2\ldots$, the sample paths of the ${\cal GW}_{\kappa,\mu,\beta}$ model are $l$ times differentiable, in any direction, if and only if $\kappa>l-0.5$ and for $0\leq\kappa\leq 0.5$ they are not differentiable.
    When $\kappa=k\in\mathbb{N}$, the Generalized Wendland correlation simplifies into the product of an Askey (truncated power) correlation with a polynomial of degree $k$:
    ${\cal GW}_{k,\mu,\beta}=\left(1-\frac{x}{\beta}\right)^{\mu+k}_{+}P_{k}(x;\mu,\beta),\quad k=0,1,2,\ldots$ (7)
    where $P_{k}(h,\mu,\beta)$ is a polynomial of degree $k$ that can be expressed as:
    $P_{k}(x;\mu,\beta)=\frac{2^{k}k!}{(2k)!}\sum_{j=0}^{k}c_{j,k}(\mu)\left(\frac{x}{\beta}\right)^{k-j}\left(1-\frac{x}{\beta}\right)^{j},$ (8)
    and $c_{j,k}(\mu)$ are suitable constants (Bevilacqua et al.,, 2024).
    Recently, Bevilacqua et al., (2022) considered a compact support reparameterized version of the ${\cal GW}$ model, and proved that
    $\lim_{\mu\to\infty}{\cal GW}_{\kappa,\mu,g(\beta,\mu,\kappa)}(x)={\cal MT}_{\kappa+1/2,\beta}(x),\quad\kappa\geq-0.5,$ (9)
    where $g(\beta,\mu,\kappa)=\beta\left(\frac{\Gamma(\mu+2\kappa+1)}{\Gamma(\mu)}\right)^{\frac{1}{1+2\kappa}}$, with uniform convergence on $[0,\infty)$.
    A simpler compact support reparameterization that leads to the ${\cal MT}$ model as a special limit case has been proposed in Emery et al., (2025):
    $\lim_{\mu\to\infty}{\cal GW}_{\kappa,\mu,\mu\beta}(x)={\cal MT}_{\kappa+1/2,\beta}(x),\quad\kappa\geq-0.5,$ (10)
    again with uniform convergence on $[0,\infty)$.
    As a consequence, the two reparameterized ${\cal GW}$ models can be viewed as a generalization of the ${\cal MT}$ model with an additional parameter that, for given smoothness and spatial dependence parameters, allows switching from the world of flexible compactly supported covariance functions to the world of flexible globally supported covariance functions.
3 The Gauss hypergeometric model and the identifiability problem
The $\mathcal{GH}$ class in Emery and Alegría, (2022) has been defined starting from the following spectral density:
$\widehat{{\cal GH}}_{\delta,\beta,\gamma,a}(z)=\frac{\Gamma(\delta)\Gamma(\beta-\frac{d}{2})\Gamma(\gamma-\frac{d}{2})a^{d}}{2^{d}\pi^{\frac{d}{2}}\Gamma\left(\delta-\frac{d}{2}\right)\Gamma(\beta)\Gamma(\gamma)}\mathstrut_{1}F_{2}\Big{(}\delta;\beta,\gamma;-\frac{(za)^{2}}{4}\Big{)},\quad z\geq 0.$ (11)
Note that (11) has basically the same form of the spectral density of the ${\cal GW}$ model in (5). However, in (11) there is no dependence between the parameters of the $\mathstrut{}_{1}F_{2}$ function as in the ${\cal GW}$ case.
Using some results in Cho and Yun, (2018), Emery and Alegría, (2022) show that, for a given $a>0$, the following conditions:
  1. (a)
    $\delta>\frac{d}{2}$;
  2. (b)
    $2(\beta-\delta)(\gamma-\delta)\geq\delta$;
  3. (c)
    $\beta+\gamma\geq 3\delta+\frac{1}{2}$;
are sufficient for the positivity of the spectral density on $[0,\infty)$. The associated four-parameter isotropic correlation model is given by:
$\begin{split}\mathcal{GH}_{\delta,\beta,\gamma,a}(x)&=\frac{\Gamma(\beta-\frac{d}{2})\Gamma(\gamma-\frac{d}{2})}{\Gamma(\beta-\delta+\gamma-\frac{d}{2})\Gamma(\delta-\frac{d}{2})}\left(1-\frac{x^{2}}{a^{2}}\right)_{+}^{\beta-\delta+\gamma-\frac{d}{2}-1}\\ &\times{}_{2}F_{1}\left(\beta-\delta;\gamma-\delta;\beta-\delta+\gamma-\frac{d}{2};\left(1-\frac{x^{2}}{a^{2}}\right)_{+}\right),\quad x\geq 0.\end{split}$ (12)
Here $a>0$ is a compact support parameter, $\delta>\frac{d}{2}$ is a smoothness parameter, and $\beta$ and $\gamma$ are two additional shape parameters.
As in the ${\cal MT}$ and ${\cal GW}$ cases, the ${\cal GH}$ model allows parameterizing in a continuous fashion the mean squared differentiability of the underlying Gaussian RF and its associated sample paths through the smoothness parameter $\delta$. Furthermore, it is not difficult to see that the ${\cal GH}$ model includes the ${\cal GW}$ and the ${\cal MT}$ models as special cases. For instance,
$\mathcal{GH}_{\delta,\delta+\frac{\mu}{2},\delta+\frac{\mu}{2}+\frac{1}{2},a}(x)={\cal GW}_{\kappa,\mu,a}(x),$ (13)
where $\delta=\frac{d+1}{2}+\kappa$.
Although the ${\cal GH}$ model is highly flexible, it suffers from identifiability problems. This is apparent under an increasing domain scenario, where all the parameters can be estimated, since $\mathcal{GH}_{\delta,\beta,\gamma,a}=\mathcal{GH}_{\delta,\gamma,\beta,a}$ due to the symmetry of the Gauss hypergeometric function: ${}_{2}F_{1}\left(a;b;\cdot;x\right)={}_{2}F_{1}\left(b;a;\cdot;x\right)$. This implies that there is not a one-to-one correspondence between the parametric space and the correlation function.
To study identifiability under fixed-domain asymptotics, the following theorem establishes the equivalence condition for two distinct Gaussian measures with $\mathcal{GH}$ correlation models that share a common smoothness parameter. Hereafter, we denote by $P(\rho_{i})$, $i=0,1$, a zero-mean Gaussian measure with an isotropic covariance function $\rho_{i}$.
Theorem 1.
Let $P(\sigma_{i}^{2}{\cal GH}_{\delta,\beta_{i},\gamma_{i},a_{i}})$, $i=0,1$, be two zero-mean Gaussian measures. If $\frac{d+1}{2}+3\delta\leq\min(\beta_{1}+\gamma_{1},\beta_{0}+\gamma_{0})$, for any bounded infinite set $D\subset\mathbb{R}^{d}$, $d=1,2,3$, the Gaussian measures $P(\sigma_{1}^{2}{\cal GH}_{\delta,\beta_{1},\gamma_{1},a_{1}})$ and $P(\sigma_{0}^{2}{\cal GH}_{\delta,\beta_{0},\gamma_{0},a_{0}})$ are equivalent on the paths of $Z(\boldsymbol{s}),\boldsymbol{s}\in D$, if and only if
$\frac{\sigma_{0}^{2}L(\beta_{0},\gamma_{0})}{a_{0}^{2\delta-d}}=\frac{\sigma_{1}^{2}L(\beta_{1},\gamma_{1})}{a_{1}^{2\delta-d}},$ (14)
where $L(\beta_{i},\gamma_{i})=\frac{2^{2\delta-d}\Gamma(\beta_{i}-d/2)\Gamma(\gamma_{i}-d/2)}{\Gamma(\gamma_{i}-\delta)\Gamma(\beta_{i}-\delta)}$ with $i=0,1.$
An immediate consequence of Theorem 1 is that the microergodic parameter of the $\mathcal{GH}_{\delta,\beta,\gamma,a}$ model is given by $\frac{\sigma^{2}L(\beta,\gamma)}{a^{2\delta-d}}$. Since $L(\beta,\gamma)=L(\gamma,\beta)$, it is clear that there is not a unique correspondence between the parametric space and the microergodic parameter, which implies that the microergodic parameter is not identifiable.
Summarizing our results, the $\mathcal{GH}$ correlation model is highly flexible, including the $\mathcal{GW}$ and the $\mathcal{MT}$ models as special cases. However, it is not identifiable under both increasing and fixed-domain scenarios. To fix this issue, in the next section, we will assume that the second shape parameter ($\gamma$) is a function of the first shape parameter ($\beta$).
4 Fixing the identifiability problem: the Hypergeometric correlation model
In this Section we first provide necessary and sufficient conditions for the validity of the ${\cal GH}$ model. Then we solve the identifiability problem by considering a parsimonious version that responds to an optimality criterion. With this goal in mind, we consider a ${\cal GW}$-type reparameterization of the $\mathcal{GH}_{\delta,\beta,\gamma,a}$ model, that is, we set
$\delta=\frac{d+1}{2}+\kappa,\quad\beta=\delta+\frac{\mu}{2},\quad\gamma=\beta+l,$ (15)
with $l\geq 0$. As outlined in (13), this parameterization includes the ${\cal GW}$ model as a special case when $l=\frac{1}{2}$.
The following theorem can be viewed as an improvement of Theorem 1 in Emery and Alegría, (2022) and provides necessary and sufficient conditions for the validity of the $\mathcal{GH}_{\delta,\delta+\frac{\mu}{2},\delta+\frac{\mu}{2}+l,a}$ model.
Theorem 2.
Let $d$ be a fixed dimension, $\delta=\frac{d+1}{2}+\kappa$ and $\kappa>-\frac{1}{2}$ a smoothness parameter. Then $\mathcal{GH}_{\delta,\delta+\frac{\mu}{2},\delta+\frac{\mu}{2}+l,a}\in\Phi_{d}$ if and only if
  1. 1.
    $\mu\geq\delta-l+\frac{1}{2}\quad$ if $\quad 0\leq l\leq\frac{d}{2}+\kappa$
  2. 2.
    $\mu\geq\sqrt{2\kappa+l^{2}+d+1}-l\quad$ if $\quad l>\frac{d}{2}+\kappa$.
It should be stressed that necessary and sufficient conditions can also be found for $l<0$. They are symmetric with respect to the case $l\geq 0$ (see the proof of Theorem 2 in Appendix A). However, to solve the issue of identifiability, hereafter we consider the condition $l\geq 0$ as in Theorem 2.
If $l=\frac{1}{2}$ then the sufficient and necessary conditions of the ${\cal GW}_{\nu,\mu,a}$ model in (6) are obtained as special case. It is apparent that, to solve the identifiability problem, $l$ should be fixed (as in the ${\cal GW}$ model) or should depend on the other parameters and/or the dimension, so an interesting question is: which is the “best” $l\geq 0$ in the $\mathcal{GH}_{\delta,\delta+\frac{\mu}{2},\delta+\frac{\mu}{2}+l,a}$ family?
This question can be responded by establishing an optimality criterion. Here we consider the criterion of maximum integral range that can be attained by a compactly supported correlation for a given smoothness and compact support. For an isotropic correlation $\phi\in\Phi_{d}$, the integral range is defined as
$A=\int_{\mathbb{R}^{d}}\phi(\|\boldsymbol{h}\|)\text{d}\boldsymbol{h}=\frac{\pi^{\frac{d}{2}}}{2^{d-1}\Gamma(\frac{d}{2})}\int_{\mathbb{R}^{+}}x^{d-1}\phi(x)\text{d}x.$ (16)
$A$ is the isotropic version of the integral range proposed in Lantuéjoul, (1991) for stationary correlation functions. This measure has a close relation with the ergodicity of the RF, with the definition of an effective number of independent observations, and with the definition of long-range dependence (Yaglom,, 1987; Matheron,, 1989). In particular, following Lim and Teo, (2009), a RF with isotropic correlation $\phi$ is long-range dependent if $A=\infty$ and short-range dependent when $A<\infty$. If $\phi(\|\cdot\|)$ is absolutely integrable in $\mathbb{R}^{d}$ (in particular, if it is compactly supported), then $A=\widehat{\phi}(0)$. Hristopoulos, (2020) interprets $A^{1/d}$ as an average measure of the distance over which two RF values are correlated.
When $d=1$, the integral range corresponds to the area under the unidimensional correlation function. A parametric correlation model that can cover a wide range of areas, from 00 to a specific upper bound, can be considered a flexible model (see Figure  1 and Theorem 6 hereunder). As a consequence, this can serve as a criterion for selecting between two different parametric correlation models, particularly when comparing two compactly supported correlation models that share the same smoothness and compact support parameters. In such cases, the correlation model that achieves a larger integral range is preferred.
For a parametric correlation model $\phi_{\boldsymbol{\theta}}\in\Phi_{d}$, we will denote its associated integral range as $A_{\boldsymbol{\theta},d}$. For instance, using formula 6.561.16 of Gradshteyn and Ryzhik, (2007), for the ${\cal MT}_{\nu,\alpha}$ correlation one has:
$A_{\nu,\alpha,d}=\frac{\alpha^{d}\pi^{d/2}\Gamma\left(\nu+\frac{d}{2}\right)}{\Gamma(\nu)}.$ (17)
The following theorem provides an expression for the integral range associated with the $\mathcal{GH}_{\delta,\delta+\frac{\mu}{2},\delta+\frac{\mu}{2}+l,a}$ correlation model.
Theorem 3.
The integral range of $\mathcal{GH}_{\delta,\delta+\frac{\mu}{2},\delta+\frac{\mu}{2}+l,a}$ is given by:
$A_{\kappa,\mu,l,a,d}=a^{d}\left[\frac{\pi^{d/2}\Gamma(\frac{d+1}{2}+\kappa)\Gamma\left(\frac{\mu+1}{2}+\kappa\right)\Gamma\left(\frac{\mu+1}{2}+\kappa+l\right)}{2^{d}\Gamma(\kappa+\frac{1}{2})\Gamma\left(\frac{\mu+1+d}{2}+\kappa\right)\Gamma\left(\frac{\mu+1+d}{2}+\kappa+l\right)}\right].$ (18)
Note that, as expected, the compact support $a$ raised to power $d$ represents a scale factor in the integral range computation. In addition, using asymptotic properties of the Gamma function, it can be shown that $A_{\kappa,\mu,l,a,d}\sim\mu^{-d}$ as $\mu\to\infty$ and, as a consequence,
${\lim_{\mu\to\infty}}A_{\kappa,\mu,l,a,d}=0.$ (19)
This implies that, as the shape parameter $\mu$ increases, the integral range approaches zero, regardless of the compact support and smoothness parameters. The following theorem establishes that, for a given $l\geq 0$, the maximum integral range is obtained by fixing $\mu$ on the lower bound of its parametric space.
Theorem 4.
For an arbitrary $l\geq 0$, the maximum of $A_{\kappa,\mu,l,a,d}$ is attained when
  1. 1.
    $\mu=\mu_{1}(l):=\delta-l+\frac{1}{2}\quad$ if $\quad 0\leq l\leq\frac{d}{2}+\kappa$
  2. 2.
    $\mu=\mu_{2}(l):=\sqrt{2\kappa+l^{2}+d+1}-l\quad$ if $\quad l>\frac{d}{2}+\kappa$.
For instance, let $\kappa\geq 0$. If $l=\frac{1}{2}$, then from (13) $\mathcal{GH}_{\delta,\delta+\frac{\mu}{2},\delta+\frac{\mu+1}{2},a}={\cal GW}_{\kappa,\mu,a}$ and the maximum integral range is attained when $\mu=\mu_{1}(\frac{1}{2})=\delta=\frac{d+1}{2}+\kappa$, which is the lower bound of ${\cal GW}$ model (6).
As an example, Figure 1 (left part) depicts the ${\cal GW}_{0,\mu,1}$ model and the right part depicts the ${\cal GW}_{1,\mu,1}$ model for $\mu=\mu_{1}(\frac{1}{2}),3,10,100$. In these examples $d=1$ and it is apparent that the model with maximum area corresponds to the case $\mu=\mu_{1}(\frac{1}{2})$ (red line) in both examples.
The next step is to find $l\geq 0$ that maximizes $A_{\kappa,\mu_{i}(l),l,a,d}$ for $i=1,2$.
Theorem 5.
The maximum of $A_{\kappa,\mu_{i}(l),l,a,d}$ for $i=1,2$ is attained when $l=d/2+\kappa$.
Refer to caption Refer to caption
Figure 1: Left part: ${\cal GW}_{0,\mu,1}$ for $\mu=\mu_{1}(\frac{1}{2}),3,10,100$ with $\delta=\frac{d+1}{2}+\kappa$ and $\kappa=0$. Right part: the same Figure but with $\kappa=1$. In both cases $d=1$.
As an example, Figure 2 compares $\mathcal{GH}_{\delta,\delta+\frac{\mu_{1}(l)}{2},\delta+\frac{\mu_{1}(l)}{2}+l,1}$ when $l=0,\frac{1}{2},\frac{d}{2}+\kappa$ (black, red and blue lines, respectively). In particular the left part depicts the case $\kappa=0$ and the right part depicts the case $\kappa=1$ ($d=2$ in both cases). It is apparent that the case $l=\frac{d}{2}+\kappa$ (blue line) corresponds to the model with higher integral range and improves the maximum integral range attained by the ${\cal GW}$ models (red line).
Refer to caption Refer to caption
Figure 2: Left part: $\mathcal{GH}_{\delta,\delta+\frac{\mu(l)}{2},\delta+\frac{\mu(l)}{2}+l,1}$ with $\delta=\frac{d+1}{2}+\kappa$, $\kappa=0$, $\mu(l)=\delta-l+\frac{1}{2}$ and $l=0,\frac{1}{2},\frac{d}{2}+\kappa$ (black, red and blue lines respectively). Right part: the same Figure but with $\kappa=1$. The case $l=\frac{1}{2}$ (red line) corresponds to the $\mathcal{GW}_{\kappa,\mu,1}$ model. The blue line corresponds to the correlation model that maximizes the integral range. In both cases $d=2$.
Summarizing our results, the model that can potentially achieve the maximum integral range is $\mathcal{GH}_{\delta,\delta+\frac{\mu}{2},\delta+\frac{\mu}{2}+l,a}$ with $l=\frac{d}{2}+\kappa$ and, in particular, the maximum is achieved when $\mu=\mu_{1}(d/2+\kappa)=1$. Taking into account (19), this implies that
$0\leq A_{\kappa,\mu,l,a,d}\leq a^{d}\left[\frac{\pi^{d/2}\,\Gamma\left(\frac{d+1}{2}+\kappa\right)\,\Gamma\left(1+\kappa\right)\,\Gamma\left(1+\frac{d}{2}+2\kappa\right)}{2^{d}\,\Gamma\left(\kappa+\frac{1}{2}\right)\,\Gamma\left(1+\frac{d}{2}+\kappa\right)\,\Gamma\left(1+d+2\kappa\right)}\right].$
Accounting for (15), the model with maximum integral range (obtained when $l=\frac{d}{2}+\kappa$ and $\mu=1$), is $\mathcal{GH}_{\delta,\delta+\frac{1}{2},2\delta,a}$ and coincides with the so-called Euclid’s hat correlation model (Gneiting,, 1999) when $\kappa=0$ (triangular or tent model if $d=1$, circular model if $d=2$, spherical model if $d=3$, pentaspherical model if $d=5$, see Section 5) and with an upgraded Euclid’s hat when $\kappa\neq 0$ (Emery and Alegría,, 2022), obtained from the Euclid’s hat by a (fractional) montée or upgrading, as defined by Matheron, (1965, I.4.18).
Hereafter, we call the correlation model $\mathcal{GH}_{\delta,\delta+\frac{\mu}{2},\delta+\frac{\mu}{2}+l,a}$ with $l=\frac{d}{2}+\kappa$ as hypergeometric correlation and we denote as
$\mathcal{H}_{\kappa,\mu,a,d}(x):=\mathcal{GH}_{\delta,\delta+\frac{\mu}{2},\delta+\frac{\mu+d}{2}+\kappa,a}(x),\quad x\geq 0,$ (20)
which is defined as
$\begin{split}\mathcal{H}_{\kappa,\mu,a,d}(x)&=\frac{\Gamma(\kappa+\frac{\mu+1}{2})\Gamma(2\kappa+\frac{d+\mu+1}{2})}{\Gamma(\mu+\frac{d+1}{2}+2\kappa)\Gamma(\kappa+\frac{1}{2})}\left(1-\frac{x^{2}}{a^{2}}\right)_{+}^{\mu+\frac{d-1}{2}+2\kappa}\\ &\times{}_{2}F_{1}\left(\frac{\mu}{2};\frac{\mu+d}{2}+\kappa;\mu+\frac{d+1}{2}+2\kappa;\left(1-\frac{x^{2}}{a^{2}}\right)_{+}\right),\quad x\geq 0.\end{split}$ (21)
Necessary and sufficient validity conditions for $\mathcal{H}_{\kappa,\mu,a,d}$ can be easily obtained from Theorem 2 by setting $l=\frac{d}{2}+\kappa$. Specifically, given the smoothness parameter $\kappa>-\frac{1}{2}$, $\mathcal{H}_{\kappa,\mu,a,d}\in\Phi_{d}$ if and only if $\mu\geq 1$. Interestingly, the conditions are much simpler than the conditions (6) of the $\mathcal{GW}_{\kappa,\mu,a}$ model. In particular, the condition on $\mu$ does not depend on the dimension. The price to pay is that, unlike the $\mathcal{GW}_{\kappa,\mu,a}$ model, the expression of $\mathcal{H}_{\kappa,\mu,a,d}$ depends on the dimension $d$.
The following theorem proves that the ${\cal H}_{\kappa,1,a,d}$ subclass of hypergeometric correlation functions with maximum integral range allows recovering any member of the ${\cal H}_{\kappa,\mu,a,d}$ class, as well as great part of the $\mathcal{GH}_{\delta,\delta+\frac{\mu}{2},\delta+\frac{\mu}{2}+l,a}$ class, by scale mixtures. The chosen optimality criterion (maximizing the integral range) therefore coincides with the greatest flexibility from a modeling viewpoint.
Theorem 6.
Let $a>0$, $\kappa>-\frac{1}{2}$, $\delta=\frac{d+1}{2}+\kappa$, $\mu>1$ and $l\geq 0$ such that $\mu+2l>1+d+2\kappa$, so that $\mathcal{GH}_{\delta,\delta+\frac{\mu}{2},\delta+\frac{\mu}{2}+l,a}\in\Phi_{d}$ owing to Theorem 2. Then, $\mathcal{GH}_{\delta,\delta+\frac{\mu}{2},\delta+\frac{\mu}{2}+l,a}$ is a beta-type scale mixture of hypergeometric correlation functions of the form $\mathcal{H}_{\kappa,1,b,d}$ with $0<b\leq a$:
$\begin{split}&\mathcal{GH}_{\delta,\delta+\frac{\mu}{2},\delta+\frac{\mu}{2}+l,a}(x)=\frac{\Gamma(\delta+\frac{\mu-d}{2})\Gamma(\delta+\frac{\mu-d}{2}+l)}{\Gamma(\delta+\frac{1-d}{2})\Gamma(\delta+\frac{1}{2}+\kappa)\Gamma(\frac{\mu-1}{2})\Gamma(\frac{\mu-1-d}{2}-\kappa+l)}\\ &\times\int_{0}^{1}\int_{0}^{1}u^{\delta-\frac{d+1}{2}}(1-u)^{\frac{\mu-3}{2}}v^{\delta-\frac{1}{2}+\kappa}(1-v)^{\frac{\mu-d-3}{2}-\kappa+l}\mathcal{H}_{\kappa,1,a\sqrt{uv},d}(x){\rm d}u{\rm d}v,\quad x\geq 0.\end{split}$ (22)
In particular, for $l=\frac{d}{2}+\kappa$, one obtains
$\begin{split}\mathcal{H}_{\kappa,\mu,a,d}(x)&=\frac{\Gamma(\delta+\frac{\mu-d}{2})\Gamma(\delta+\frac{\mu}{2}+\kappa)}{\Gamma(\delta+\frac{1-d}{2})\Gamma(\delta+\frac{1}{2}+\kappa)\Gamma^{2}(\frac{\mu-1}{2})}\\ &\times\int_{0}^{1}\int_{0}^{1}u^{\delta-\frac{d+1}{2}}(1-u)^{\frac{\mu-3}{2}}v^{\delta-\frac{1}{2}+\kappa}(1-v)^{\frac{\mu-3}{2}}\mathcal{H}_{\kappa,1,a\sqrt{uv},d}(x){\rm d}u{\rm d}v,\quad x\geq 0.\end{split}$
Clearly, the ${\cal H}_{\kappa,\mu,a,d}$ model allows parameterizing in a continuous fashion the mean squared differentiability of the underlying Gaussian RF and its associated sample paths through the smoothness parameter $\kappa$ in the same way as the ${\cal GW}_{\kappa,\mu,\beta}$ model. Figure 3 depicts two realizations of zero-mean and unit-variance Gaussian RFs with correlations ${\cal H}_{\kappa,4,0.2,2}$ for $\kappa=0,1,2$. The realizations were constructed via the Cholesky decomposition of the covariance matrix.
Refer to caption Refer to caption
Figure 3: Two Gaussian RF realizations with ${\cal H}_{\kappa,4,0.2,2}$ correlation model, for $\kappa=0,1$ (from left to right).
5 Some special cases and bridges to the Matérn model
The computation of the $\mathcal{H}_{\kappa,\mu,a,d}$ correlation function requires an efficient evaluation of the Gauss hypergeometric function, which is feasible (Johansson,, 2019) but could be cumbersome to (spatial) statisticians or machine learning practitioners used to handle closed-form parametric correlation models.
However, similarly to the $\mathcal{MT}$ and $\mathcal{GW}$ models, in some special cases, the computation can be considerably simplified. The following theorem considers the case when $\kappa$ is a non-negative integer, $d=1$ and $\mu\geq 1$.
Theorem 7.
If $\kappa=k$ is a non-negative integer, $d=1$ and $\mu\geq 1$, then:
$\begin{split}\mathcal{H}_{k,\mu,a,1}(x)&=\frac{\sqrt{\pi}\Gamma\left(\frac{\mu}{2}+2k+1\right)}{\Gamma\left(k+0.5\right)}\left(1-\frac{x}{a}\right)^{\mu+2k}\left(\frac{x}{a}\right)^{k}\\ &\times\sum\limits^{k}_{n=0}\frac{(n+k)!}{2^{2n}n!\Gamma(n+\frac{\mu}{2}+k+1)(k-n)!}\left(\frac{(1-\frac{x}{a})^{2}}{\frac{x}{a}}\right)^{n}.\end{split}$
Unfortunately, general closed forms for any $d>1$, $\mu\geq 1$ and $\kappa$ a non-negative integer, are difficult to obtain. Some special cases can, however, be obtained using closed forms of the Gauss hypergeometric function. For instance, the following theorem provides a closed form expression when $d$ is an odd integer, $\mu\geq 1$ and $\kappa=0$.
Theorem 8.
Let $d$ an odd integer and $\mu\geq 1$. If $\kappa=0$ then
$\mathcal{H}_{0,\mu,a,d}(x)=C\left(1-\frac{x}{a}\right)_{+}^{\mu+\frac{d-1}{2}}\sum\limits^{\frac{d-1}{2}}_{n=0}\frac{\left(\frac{1-d}{2}\right)_{n}\left(\frac{d-1}{2}+1\right)_{n}\left(1-\frac{x}{a}\right)^{n}_{+}}{2^{n}n!\Gamma\left(n+\frac{d-1}{2}+\mu+1\right)},$
with $C=\frac{2^{\mu+\frac{d-1}{2}}\Gamma\left(\frac{\mu+1}{2}\right)\Gamma\left(\frac{\mu+d+1}{2}\right)}{\sqrt{\pi}}$.
Additional closed-form expressions, including the spherical models, for some fixed $\mu$ and $\kappa$ are given in Appendix B.
The ${\cal MT}$ model can be obtained as a special limit case of the proposed $\mathcal{H}_{\kappa,\mu,a,d}$ model after two suitable compact support reparameterizations. To obtain the first parameterization, the following theorem provides sufficient conditions of equivalence between two Gaussian measures with covariances $\sigma_{0}^{2}{\cal MT}_{\nu,\alpha}$ and $\sigma_{1}^{2}\mathcal{H}_{\kappa,\mu,a,d}$, respectively. The proof can be found in Appendix A.
Theorem 9.
Let $P(\sigma_{0}^{2}{\cal MT}_{\nu,\alpha})$ and $P(\sigma_{1}^{2}\mathcal{H}_{\kappa,\mu,a,d})$ be two zero-mean Gaussian measures. If $\nu=\kappa+\frac{1}{2}$ with $\nu>-\frac{1}{2}$ and
$\frac{\sigma_{0}^{2}}{\alpha^{2\nu}}=\frac{\sigma_{1}^{2}}{a^{2\kappa+1}}B(\kappa,\mu,d),$ (23)
then for any bounded infinite set $D\subset\mathbb{R}^{d}$, $d=1,2,3$, $P(\sigma_{0}^{2}{\cal MT}_{\nu,\alpha})\equiv P(\sigma_{1}^{2}\mathcal{H}_{\kappa,\mu,a,d})$ on the paths of $Z(\boldsymbol{s}),\boldsymbol{s}\in D$, where
$B(\kappa,\mu,d)=\frac{2^{2\kappa+1}\Gamma(\frac{\mu+1}{2}+\kappa)\Gamma(\frac{\mu+d+1}{2}+2\kappa)}{\Gamma(\frac{\mu}{2})\Gamma(\frac{\mu+d}{2}+\kappa)}.$ (24)
Note that the equivalence condition (23) sharing a common variance parameter and taking into account the relation between the smoothness parameters $\nu=\kappa+\frac{1}{2}$, can be written as
$a=\alpha B(\kappa,\mu,d)^{\frac{1}{1+2\kappa}}:=b(\alpha,\mu,\kappa,d).$
If we use the previous equation as a compact support reparameterization in the ${\cal H}$ model, then, using the same arguments as in Bevilacqua et al., (2022), it can be shown that
$\lim_{\mu\to\infty}{\cal H}_{\kappa,\mu,b(\alpha,\mu,\kappa,d),d}(x)={\cal MT}_{\kappa+\frac{1}{2},\alpha}(x),\quad\kappa>-\frac{1}{2},$ (25)
with uniform convergence on $[0,\infty)$. In addition, following the same arguments in Emery et al., (2025), it can be shown that
$\lim_{\mu\to\infty}{\cal H}_{\kappa,\mu,\mu\alpha,d}(x)={\cal MT}_{\kappa+\frac{1}{2},\alpha}(x),\quad\kappa\geq-\frac{1}{2},$ (26)
with uniform convergence on $[0,\infty)$. As a consequence, the two reparameterized versions of the ${\cal H}$ model can be viewed as generalizations of the ${\cal MT}$ model with an additional parameter $\mu$ that, for given smoothness and spatial dependence parameters, allows to obtain compactly ($\mu<\infty$) or globally ($\mu=\infty$) supported covariance functions. Thus, under this parameterization, $\mu$ can be fixed by the user, particularly when dealing with large datasets, to achieve sparse covariance matrices and obtain computational benefits.
An illustration of (25) can be found in Figure 4 where the ${\cal H}_{\kappa,\mu,b(\alpha,\mu,\kappa,d),d}$ model with increasing values of $\mu$ is compared with the ${\cal MT}_{\kappa+\frac{1}{2},\alpha}$ model when $d=2$.
Refer to caption Refer to caption
Figure 4: The ${\cal H}_{\kappa,\mu,b(\alpha,\mu,\kappa,2),2}$ for $\mu=1,2,10$ when $\kappa=0$ (left) and $\kappa=1$ (right). The red line is the ${\cal MT}_{\kappa+\frac{1}{2},\alpha}$ correlation model ($\alpha=0.2$ in both cases).
It should be emphasized that, under this compact support reparameterization, the role of the shape parameter $\mu$ changes significantly. Let us assume, without loss of generality, that the scale dependence parameter $\alpha$ is equal to $1$. Under this assumption, the parameter $\mu$ (together with the smoothness parameter and the dimension $d$) determines the compact support. For large values of $\mu$, the compact support becomes larger, and both the ${\cal H}_{\kappa,\mu,b(1,\mu,\kappa,d),d}$ and ${\cal H}_{\kappa,\mu,\mu,d}$ models approach the ${\cal MT}_{\kappa+\frac{1}{2},1}$ model. In this case, the maximum integral range is the one associated with the ${\cal MT}$ model in Equation (17).
Conversely, as $\mu$ approaches its lower bound, the compact support shrinks, and a potential smaller compact support results in a more flexible model. If we compare the compact support of $\mathcal{H}_{\kappa,\mu,b(1,\mu,\kappa,d),d}$ with that of $\mathcal{GW}_{\kappa,\mu,g(1,\mu,\kappa)}$ (see Equation 9) and given that $\mu$ is lower bounded by $1$ for the $\mathcal{H}$ model and by $\mu_{\min}(\kappa,d)$ for the $\mathcal{GW}$ model, it can be shown that
$b(1,1,\kappa,d)\leq g(1,\mu_{\min}(\kappa,d),\kappa)$
when $d\geq 2$ or $\kappa\geq 0$ and $b(1,1,\kappa,d)>g(1,\mu_{\min}(\kappa,d),\kappa)$ when $d=1$ and $-\frac{1}{2}<\kappa<0$. This implies that the $\mathcal{H}_{\kappa,\mu,b(1,\mu,\kappa,d),d}$ model achieves a smaller compact support compared to the $\mathcal{GW}_{\kappa,\mu,g(1,\mu,\kappa)}$ model for a significant portion of the parameter space. As a consequence, the $\mathcal{H}_{\kappa,\mu,b(1,\mu,\kappa,d),d}$ model is more flexible from this perspective.
On the other hand, if we compare the smaller compact supports of $\mathcal{H}_{\kappa,\mu,\mu,d}$ and $\mathcal{GW}_{\kappa,\mu,\mu}$, then we reach the same conclusions since $1\leq\mu_{\min}(\kappa,d)$ when $d\geq 2$ or $\kappa\geq 0$ and $1>\mu_{\min}(\kappa,d)$ when $d=1$ and $-\frac{1}{2}<\kappa<0$.
We also compare the smaller compact supports that can be achieved by the ${\cal H}_{\kappa,\mu,b(1,\mu,\kappa,d),d}$ and ${\cal H}_{\kappa,\mu,\mu,d}$ models, which are $b(1,1,\kappa,d)$ and $1$, respectively. Figure 5 illustrates the compact support $b(1,1,\kappa,d)$ for $-0.5<\kappa<1.5$ and $d=1,2,3,4,5$, compared to the reference value $1$. It can be observed that only for a small region of the smoothness parameter space and $d=1,2,3,4$, we obtain $b(1,1,\kappa,d)<1$.
Figure 5: The smaller compact support $b(1,1,\kappa,d)$ that can be achieved by the model ${\cal H}_{\kappa,\mu,b(1,\mu,\kappa,d),d}$ for $d=1,2,3,4,5$ and $-0.5<\kappa<1.5$. The red line is the smaller compact support that can be achieved by the ${\cal H}_{\kappa,\mu,\mu,d}$ model.
Summarizing our results, when comparing the two reparameterized versions of the $\mathcal{H}$ model with the analogous versions of the $\mathcal{GW}$ model that share the same smoothness parameter $\kappa$ and converge to the same $\mathcal{MT}$ model as the shape parameter $\mu$ becomes very large, it turns out that the reparameterized $\mathcal{H}$ models can achieve smaller compact supports. In addition the ${\cal H}_{\kappa,\mu,\mu,d}$ model can achieve smaller compact supports with respect to the ${\cal H}_{\kappa,\mu,b(1,\mu,\kappa,d),d}$ model.
6 Asymptotic properties of the maximum likelihood estimator of the Hypergeometric model
When considering the asymptotic properties of the ML estimator for the parameters of the $\sigma^{2}\mathcal{H}_{\kappa,\mu,a,d}$ model, two different scenarios can be examined. Under increasing domain asymptotics, the parameters $(\sigma^{2},\kappa,\mu,a)$ can be estimated consistently. Furthermore, the asymptotic distribution of the ML estimator is Gaussian, with a covariance matrix equal to the inverse of the Fisher information matrix (Mardia and Marshall,, 1984).
Under fixed-domain, using the results of Theorem 1 and taking into account the parameterization (15) with $l=\frac{d}{2}+\kappa$, the equivalence condition for the two Gaussian measures $P(\sigma_{i}^{2}{\cal H}_{\kappa,\mu_{i},a_{i},d})$, $i=0,1$, sharing a common smoothness parameter, when $d=1,2,3$ and provided that $\min(\mu_{0},\mu_{1})\geq 1+\frac{d}{2}$, reads:
$\frac{\sigma_{0}^{2}B(\kappa,\mu_{0},d)}{a_{0}^{2\kappa+1}}=\frac{\sigma_{1}^{2}B(\kappa,\mu_{1},d)}{a_{1}^{2\kappa+1}},$ (27)
As a consequence only the microergodic parameter $\sigma^{2}B(\kappa,\mu,d)/a^{2\kappa+1}$ can be estimated consistently. On the other hand, assuming a fixed known shape parameter $\mu$, the microergodic parameter is given by $\sigma^{2}/a^{2\kappa+1}$.
In what follows, we provide the asymptotic properties of ML estimator of the microergodic parameter $\sigma^{2}/a^{2\kappa+1}$. In particular, we will show that the microergodic parameter can be estimated consistently, and then we will assess the asymptotic distribution of the ML estimator.
Let $D$ be a bounded subset of $\mathbb{R}^{d}$ and let $\boldsymbol{Z}_{n}=(Z(\boldsymbol{s}_{1}),\ldots,Z(\boldsymbol{s}_{n}))^{\prime}$ be a finite realization of $Z(\boldsymbol{s})$, $\boldsymbol{s}\in D$, a zero-mean stationary Gaussian RF with a given parametric covariance function $\sigma^{2}\phi(\|\cdot\|;\boldsymbol{\tau})$, with $\sigma^{2}>0$, $\boldsymbol{\tau}$ a parameter vector and $\phi$ a member of the class $\Phi_{d}$, with $\phi(0;\boldsymbol{\tau})=1$.
We then write $R_{n}(\boldsymbol{\tau})=[\phi(\|\boldsymbol{s}_{i}-\boldsymbol{s}_{j}\|;\boldsymbol{\tau})]_{i,j=1}^{n}$ for the associated correlation matrix. The Gaussian log-likelihood function is defined as:
$\mathcal{L}_{n}(\sigma^{2},\boldsymbol{\tau})=-\frac{1}{2}\left(n\log(2\pi\sigma^{2})+\log(|R_{n}(\boldsymbol{\tau})|)+\frac{1}{\sigma^{2}}\boldsymbol{Z}_{n}^{\prime}R_{n}(\boldsymbol{\tau})^{-1}\boldsymbol{Z}_{n}\right).$ (28)
Under the $\cal H$ model, the Gaussian log-likelihood is obtained with $\phi(\cdot;\boldsymbol{\tau})\equiv\mathcal{H}_{\kappa,\mu,a,d}$ with $\boldsymbol{\tau}=(\kappa,\mu,a)^{\prime}$. Since in what follows $\mu$ and $\kappa$ are assumed known and fixed, for notation convenience, we write $\tau=a$.
Let $\hat{\sigma}^{2}_{n}$ and $\hat{a}_{n}$ be the ML estimator obtained maximizing $\mathcal{L}_{n}(\sigma^{2},a)$ for a fixed $\mu$ and $\kappa$. We consider two types of estimators. The first maximizes (28) with respect to $\sigma^{2}$ for a fixed arbitrary compact support $a>0$, obtaining the following estimator
$\hat{\sigma}_{n}^{2}(a)=\operatorname*{arg\,max}_{\sigma^{2}}\mathcal{L}_{n}(\sigma^{2},a)=\boldsymbol{Z}_{n}^{\prime}R_{n}(a)^{-1}\boldsymbol{Z}_{n}/n.$ (29)
Here $R_{n}(a)$ is the correlation matrix coming from the $\mathcal{H}_{\kappa,\mu,a,d}$ model. The following result establishes strong consistency and asymptotic distribution of the ML estimator $\hat{\sigma}_{n}^{2}(a)/a^{(1+2\kappa)}$ of the microergodic parameter. We omit the proof, as it follows the same steps as Theorem 8 in Bevilacqua et al., (2019).
Theorem 10.
Let $Z(\boldsymbol{s})$, $\boldsymbol{s}\in D\subset\mathbb{R}^{d}$, $d=1,2,3$, be a zero-mean Gaussian RF with $\sigma^{2}\mathcal{H}_{\kappa,\mu,a,d}$ covariance model with $\mu\geq 1+d/2$. Suppose $(\sigma_{0}^{2},a_{0})\in(0,\infty)\times(0,\infty)$. For a fixed $a>0$, let $\hat{\sigma}_{n}^{2}(a)$ as defined in Equation (29). Then, as $n\to\infty$,
  1. 1.
    $\hat{\sigma}_{n}^{2}(a)/a^{1+2\kappa}\stackrel{{\scriptstyle a.s}}{{\longrightarrow}}\sigma_{0}^{2}(a_{0})/a_{0}^{1+2\kappa}$ and
  2. 2.
    $\sqrt{n}(\hat{\sigma}_{n}^{2}(a)/a^{1+2\kappa}-\sigma_{0}^{2}(a_{0})/a_{0}^{1+2\kappa})\stackrel{{\scriptstyle\mathcal{D}}}{{\longrightarrow}}\mathcal{N}(0,2(\sigma_{0}^{2}(a_{0})/a_{0}^{1+2\kappa})^{2})$.
The second type of estimation technique considers the joint maximization of (28) with respect to $(\sigma^{2},a)\in(0,\infty)\times I$ where $I=[a_{L},a_{U}]$ and $0<a_{L}<a_{U}<\infty$. The solution of this optimization problem is given by $(\hat{\sigma}_{n}^{2}(\hat{a}_{n}),\hat{a}_{n})$ where
$\hat{\sigma}_{n}^{2}(\hat{a}_{n})=\boldsymbol{Z}_{n}^{\prime}R_{n}(\hat{a}_{n})^{-1}\boldsymbol{Z}_{n}/n$
and $\hat{a}_{n}=\operatorname*{arg\,max}_{a\in I}\mathcal{PL}_{n}(a)$. Here $\mathcal{PL}_{n}(a)$ is the profile log-likelihood:
$\mathcal{PL}_{n}(a)=-\frac{1}{2}\left(\log(2\pi)+n\log(\hat{\sigma}_{n}^{2}(a))+\log|R_{n}(a)|+n\right).$ (30)
In order to establish strong consistency and asymptotic distribution of the sequence of random variables $\hat{\sigma}_{n}^{2}(\hat{a}_{n})/\hat{a}_{n}^{1+2\kappa}$, we use the following lemma that establishes the monotone behavior of $\hat{\sigma}_{n}^{2}(a)/a^{1+2\kappa}$ when viewed as a function of $a\in I$ under specific conditions on the $\mu$ parameter. The proof can be found in Appendix A.
Lemma 1.
Let $S_{n}=\{\boldsymbol{s}_{1},\ldots,\boldsymbol{s}_{n}\in D\subset\mathbb{R}^{d}\}$ denote any set of distinct locations. For any $a_{1}<a_{2}\in I$ and for each $n$, $\hat{\sigma}_{n}^{2}(a_{1})/a_{1}^{1+2\kappa}\leq\hat{\sigma}_{n}^{2}(a_{2})/a_{2}^{1+2\kappa}$ if and only if $\mu\geq 4$.
Finally, the following theorem establishes the asymptotic properties of the ML estimator $\hat{\sigma}_{n}^{2}(\hat{a}_{n})/\hat{a}_{n}^{1+2\kappa}$. We omit the proof, as it follows the same steps as Theorem 9 in Bevilacqua et al., (2019) using the results in Lemma 1.
Theorem 11.
Let $Z(\boldsymbol{s})$, $\boldsymbol{s}\in D\subset\mathbb{R}^{d}$, $d=1,2,3$, be a zero-mean Gaussian RF with $\sigma^{2}\mathcal{H}_{\kappa,\mu,a,d}$ covariance model with $\mu\geq 4$. Suppose $(\sigma_{0}^{2},a_{0})\in(0,\infty)\times I$ where $I=[a_{L},a_{U}]$ with $0<a_{L}<a_{U}<\infty$. Let $(\hat{\sigma}_{n}^{2},\hat{a}_{n})^{\prime}$ maximize (28) over $(0,\infty)\times I$. Then as $n\to\infty$,
  1. 1.
    $\hat{\sigma}_{n}^{2}(\hat{a}_{n})/\hat{a}_{n}^{1+2\kappa}\stackrel{{\scriptstyle a.s}}{{\longrightarrow}}\sigma_{0}^{2}(a_{0})/a_{0}^{1+2\kappa}$ and
  2. 2.
    $\sqrt{n}(\hat{\sigma}_{n}^{2}(\hat{a}_{n})/\hat{a}_{n}^{1+2\kappa}-\sigma_{0}^{2}(a_{0})/a_{0}^{1+2\kappa})\stackrel{{\scriptstyle\mathcal{D}}}{{\longrightarrow}}\mathcal{N}(0,2(\sigma_{0}^{2}(a_{0})/a_{0}^{1+2\kappa})^{2})$.
We conclude this section with some comments on the reparameterized models $\sigma^{2}\mathcal{H}_{\kappa,\mu,b(\alpha,\mu,\kappa,d),d}$ and $\sigma^{2}\mathcal{H}_{\kappa,\mu,\alpha\mu,d}$. In particular, assuming $\mu$ unknown and using (27), for the latter model the associated microergodic parameter is given by $\sigma^{2}B(\kappa,\mu,d)/(\alpha\mu)^{2\kappa+1},$ and for the former model, it is given by $\sigma^{2}/\alpha^{2\kappa+1}.$ Notably, the microergodic parameter in the second case does not depend on $\mu$. This implies that, for the $\sigma^{2}\mathcal{H}_{\kappa,\mu,b(\alpha,\mu,\kappa,d),d}$ model, we have two microergodic parameters, $\sigma^{2}/\alpha^{2\kappa+1}$ and $\mu$, that can be estimated consistently under fixed-domain asymptotics. As a consequence, from an inferential point of view, the model $\sigma^{2}\mathcal{H}_{\kappa,\mu,b(\alpha,\mu,\kappa,d),d}$ should be preferred over the $\sigma^{2}\mathcal{H}_{\kappa,\mu,\alpha\mu,d}$ model. However, particularly for large datasets, the $\mu$ parameter should be fixed, as it allows the construction of sparse covariance matrices. In this case, the microergodic parameter is given by $\sigma^{2}/\alpha^{2\kappa+1}$ for both parameterizations.
7 A simulation study
In this section, we perform a simulation study on the estimation of the $\sigma^{2}\mathcal{H}_{\kappa,\mu,a,d}$ family. We simulate 500 realizations, using Cholesky decomposition, of a zero-mean Gaussian RF with covariance $\sigma^{2}\mathcal{H}_{\kappa,\mu,a,d}$ with $d=2$, setting $\sigma^{2}=1$, $\kappa=0,1,2$, and $\mu=4,6,8$. We have considered a regular grid with increments $0.025$ over $[0,1]^{2}$, that is $1681$ grid points. Then the grid points have been perturbed, adding a uniform random value on $[-0.01,0.01]$ to each coordinate. We assume the parameters $\kappa$ and $\mu$ known and we focus on the ML estimation of $\sigma^{2}$ and $a$ and the microergodic parameter $\sigma^{2}/a^{2\kappa+1}$ under both increasing and fixed domain scenarios. Since the increasing as well as the fixed-domain frameworks can be mimicked by fixing the number of location sites over a given spatial domain and decreasing or increasing the spatial dependence (Zhang and Zimmerman,, 2005; Kaufman and Shaby,, 2013), we set the compact support $a=0.1,0.25,0.5$. We consider $a=0.1$ as an increasing domain scenario and $a=0.5$ as a fixed domain scenario.
Table 1 reports biases and standard deviations (in parentheses) of the 500 ML parameter estimates, namely $\widehat{\sigma}_{i}^{2}$, $\widehat{a}_{i}$, and $\widehat{\sigma}_{i}^{2}/\widehat{a}_{i}^{2\kappa+1}$ for $i=1,\ldots,500$, $\kappa=0,1,2$, $\mu=4,6,8$ and $a=0.1,0.25,0.5$.
Table 1: Bias and standard deviation (in parenthesis) of ML estimates $\widehat{\sigma}_{i}^{2}$, $\widehat{a}_{i}$, and $\widehat{\sigma}_{i}^{2}/\widehat{a}_{i}^{2\kappa+1}$ for $i=1,\ldots,500$ when estimating a zero-mean Gaussian RF with covariance $\sigma^{2}\mathcal{H}_{\kappa,\mu,a,2}$ when $\kappa=0,1,2$, $\mu=4,6,8$, $a=0.1,0.25,0.5$ and $\sigma^{2}=1$.
$\kappa=0$ $\kappa=1$ $\kappa=2$
$\mu=4$ $\mu=6$ $\mu=8$ $\mu=4$ $\mu=6$ $\mu=8$ $\mu=4$ $\mu=6$ $\mu=8$
$\widehat{\sigma}^{2}$ $a=0.1$ $0.0037$ $0.0025$ $0.0018$ $0.0060$ $0.0039$ $0.0028$ $0.0053$ $0.0038$ $0.0029$
$(0.0422)$ $(0.0382)$ $(0.0368)$ $(0.0485)$ $(0.0427)$ $(0.0398)$ $(0.0475)$ $(0.0432)$ $(0.0407)$
$a=0.25$ $0.0030$ $0.0047$ $0.0046$ $0.0085$ $0.0098$ $0.0095$ $0.0100$ $0.0116$ $0.0111$
$(0.0808)$ $(0.0622)$ $(0.0522)$ $(0.1082)$ $(0.0870)$ $(0.0737)$ $(0.1044)$ $(0.088)$ $(0.0773)$
$a=0.5$ $-0.0023$ $0.0024$ $0.0045$ $0.0029$ $0.0079$ $0.0101$ $0.0052$ $0.0091$ $0.0116$
$(0.1446)$ $(0.1103)$ $(0.0900)$ $(0.1921)$ $(0.1592)$ $(0.135)$ $(0.1892)$ $(0.1599)$ $(0.1397)$
$\widehat{a}$ $a=0.1$ $0.0006$ $0.0007$ $0.0007$ $0.0004$ $0.0004$ $0.0004$ $0.0002$ $0.0003$ $0.0003$
$(0.0049)$ $(0.0055)$ $(0.0068)$ $(0.0024)$ $(0.0028)$ $(0.0033)$ $(0.0017)$ $(0.0021)$ $(0.0026)$
$a=0.25$ $0.0007$ $0.0014$ $0.0016$ $0.0005$ $0.0008$ $0.0009$ $0.0004$ $0.0006$ $0.0007$
$(0.0214)$ $(0.0176)$ $(0.0157)$ $(0.0099)$ $(0.0086)$ $(0.0079)$ $(0.0059)$ $(0.0054)$ $(0.0052)$
$a=0.5$ $-0.0015$ $0.0010$ $0.0023$ $-0.0016$ $-0.0001$ $0.0008$ $-0.0009$ $-0.0001$ $0.0005$
$(0.0737)$ $(0.0578)$ $(0.0488)$ $(0.0329)$ $(0.0280)$ $(0.0245)$ $(0.0196)$ $(0.0171)$ $(0.0154)$
$\frac{\widehat{\sigma}^{2}}{\widehat{a}^{2\kappa+1}}$ $a=0.1$ $-0.0150$ $-0.0219$ $-0.0151$ $-3.9684$ $-4.4120$ $-3.7659$ $-496.06$ $-486.01$ $-346.94$
$(0.4370)$ $(0.5362)$ $(0.6792)$ $(51.3695)$ $(68.6746)$ $(88.9434)$ $(6397.8)$ $(8930.7)$ $(11628.7)$
$a=0.25$ $0.0045$ $0.0015$ $-0.0017$ $0.0228$ $-0.0470$ $-0.1218$ $-0.2553$ $-1.7130$ $-3.0083$
$(0.1499)$ $(0.1600)$ $(0.1708)$ $(2.5102)$ $(2.7402)$ $(2.9959)$ $(41.521)$ $(46.056)$ $(51.067)$
$a=0.5$ $0.0034$ $0.0033$ $0.0026$ $0.0113$ $0.0099$ $0.0065$ $0.0399$ $0.0319$ $0.0150$
$(0.0719)$ $(0.0740)$ $(0.0765)$ $(0.2923)$ $(0.3050)$ $(0.3187)$ $(1.1922)$ $(1.2484)$ $(1.3091)$
Overall, the ML estimation of $\sigma^{2}$ and $a$ is approximately unbiased, and the standard deviations increase when switching from the increasing domain setting ($a=0.1$) to the fixed-domain setting ($a=0.5$), as expected. Another general pattern that can be observed is that, for each $\kappa$, increasing the shape parameter $\mu$ results in a lower standard deviation.
For the microergodic parameter, it can be observed that under the increasing domain scenario ($a=0.1$), the ML estimation can be significantly unbiased but with a high standard deviation, particularly when $\kappa$ increases. When moving to the fixed-domain scenario ($a=0.5$), both the bias and the standard deviation are significantly reduced, as expected. In addition, the standard deviation of the microergodic parameter increases as the smoothness parameter $\kappa$ increases.
Figure 6 depicts the boxplots of the ML estimates of $\sigma^{2}$, $a$, and the microergodic parameter $\sigma^{2}/a^{2\kappa+1}$ for $\mu=4,6,8$ and $a=0.1$ (on the left) and $a=0.5$ (on the right) when $\kappa=0$. The Gaussian distribution of the estimates for $\sigma^{2}$ and $a$ is more evident under the increasing domain scenario, as expected. On the other hand, when switching to the fixed-domain scenario, the distribution of the ML estimates of the microergodic parameter becomes more centered around the true values, and the distribution appears Gaussian, as expected.
Refer to caption Refer to caption
($a=0.1$) ($a=0.5$)
Figure 6: Left part: Boxplots of ML estimates of $\sigma^{2}$, $a$ and the microergodic parameter $\frac{\sigma^{2}}{a^{2\kappa+1}}$ (from left to right) when estimating the covariance model $\sigma^{2}\mathcal{H}_{\kappa,\mu,a,d}$ with $\sigma^{2}=1$, for $\mu=4,6,8$ and $a=0.1$ (increasing domain scenario). Right part: the same Figure with $a=0.5$ (fixed domain scenario). In both cases $\kappa=0$.
8 A real data application
We consider data from WorldClim (www.worldclim.org), a global database of high spatial resolution weather and climate data for the years 1970–2000, which can be downloaded using the R package geodata (Hijmans et al.,, 2024). Specifically, we analyze mean temperature data for September observed at 725,330 georeferenced locations across Madagascar. To make ML estimation feasible, we select a random sample of 3,500 locations and perform the analysis using a UTM projection.
Following Li and Zhang, (2011), we first detrend the data using splines to remove the cyclic patterns of both variables along the two coordinates. We then treat the residuals $y(\boldsymbol{s}_{i})$, $i=1,\ldots,3500$, as realizations from a zero-mean RF $\{Y(\boldsymbol{s}):\boldsymbol{s}\in D\}$. Figure 7, from left to right, depicts the colored map, the boxplot, and the empirical semivariogram of the residuals. The graphics suggest that a weakly stationary RF with a marginal distribution capable of modeling heavy tails is potentially an appropriate model for our data.
Refer to caption Refer to caption Refer to caption
Figure 7: From left to right: spatial map of the residuals for the Madagascar mean temperature data, boxplot, and empirical semivariogram of the data.
Let $\{Z(\boldsymbol{s}):\boldsymbol{s}\in D\}$ be a standard Gaussian RF with an isotropic parametric correlation function. In our analysis, we consider two specifications for $Y(\boldsymbol{s})$:
  • a zero-mean Gaussian RF, i.e., $Y(\boldsymbol{s})=\sigma Z(\boldsymbol{s})$, and
[…]
${\cal MT}_{\kappa,\beta}$ ${\cal GW}_{\kappa,\mu,\beta}$ ${\cal H}_{\kappa,\mu,\beta,2}$ ${\cal MT}_{\kappa,\beta}$ ${\cal GW}_{\kappa,\mu,\beta}$ ${\cal H}_{\kappa,\mu,\beta,2}$
$\sigma^{2}$ $1.2054$ $1.2066$ $1.2061$ $0.6998$ $0.6986$ $0.6984$
$(0.0553)$ $(0.0541)$ $(0.0540)$ $(0.0418)$ $(0.0403)$ $(0.0403)$
$\kappa$ $0.6079$ $0.0305$ $0.0313$ $0.5767$ $-0.0036$ $-0.0043$
$(0.0293)$ $(0.0307)$ $(0.0312)$ $(0.0265)$ $(0.0237)$ $(0.0237)$
$\beta$ $16.0229$ $85.8058$ $87.9751$ $18.5405$ $76.0127$ $77.6552$
$(1.3104)$ $(20.4440)$ $(20.9644)$ $(1.6183)$ $(12.0024)$ $(10.5911)$
$\mu$ $4.2150$ $3.8914$ $3.1346$ $2.8085$
$(1.3021)$ $(1.3026)$ $(0.6548)$ $(0.5764)$
$h$ $0.1885$ $0.1899$ $0.1898$
$(0.0172)$ $(0.0170)$ $0.0170$
loglik $-3900.01$ $-3896.88$ $-3896.94$ $-3689.88$ $-3683.81$ $-3683.94$
AIC $7806$ $7802$ $7802$ $7388$ $7378$ $7378$
RMSE $0.6593$ $0.6588$ $0.6588$ $0.6586$ $0.6581$ $0.6581$
MAE $0.4068$ $0.4065$ $0.4065$ $0.4073$ $0.4071$ $0.4070$
[…]
$Time$ $7.22$ $0.04$ $0.04$ $6.26$ $0.03$ $0.03$
[…]
Refer to caption Refer to caption
Refer to caption Refer to caption
[…]
9 Concluding Remarks
In this paper, we study the Gauss hypergeometric covariance model, a highly flexible kernel introduced by Emery and Alegría, (2022), which generalizes both the Matérn and Generalized Wendland models. We establish necessary and sufficient conditions for the validity of the model and highlight that the model suffers from identifiability issues under both increasing and fixed domain asymptotics. To address the identifiability problem, we introduce a new class of compactly supported kernels that satisfy a specific optimality criterion. In particular, the proposed model achieves the maximum integral range within the class of Gauss hypergeometric correlations. From this perspective, the proposed model can be regarded as an enhancement of the Generalized Wendland model.
The proposed model has the potential to outperform the Matérn family in terms of model fitting, prediction performance, and computational efficiency, as demonstrated in the real-data application.
Additionally, the proposed model can be easily implemented, as efficient numerical computation of the Gaussian hypergeometric function is available in various libraries, such as the GNU Scientific Library (Gough,, 2009), and in major statistical software packages, including R, MATLAB, and Python. In particular, the R package GeoModels (Bevilacqua et al.,, 2025), used in this paper for numerical experiments and applications, computes the proposed model using the Python implementation of the Gaussian hypergeometric function in the SciPy library (Virtanen et al.,, 2020).
From an inferential perspective, the results on the asymptotic properties of the maximum likelihood estimator of the microergodic parameter could be generalized to include a nugget effect, as in Tang et al., (2021). Additionally, the estimation of the smoothness parameter could, in principle, be studied following the ideas in Loh and Sun, (2023). In a Bayesian setup, the fixed-domain asymptotic posterior distribution of the microergodic parameter could be investigated following Li, (2022). Finally, the proposed model can be generalized in various directions, such as space-time, multivariate, or non-stationary contexts. These topics are left for future research.
Acknowledgement
We would like to thank Professor Igor Kondrashuk for his help in partially proving Lemma 1. This work was funded by the National Agency for Research and Development of Chile, under grants ANID FONDECYT 1240308 (M. Bevilacqua), ANID project Data Observatory Foundation DO210001 (M. Bevilacqua), MATH-AMSUD1167 22-MATH-06 (AMSUD220041) (M. Bevilacqua), ANID FONDECYT 1250862 (C. Caamaño-Carrillo) and ANID PIA AFB230001 (X. Emery).
Appendix A Proofs
A.1 Proof of Theorem 1
Proof.
Let $P(\rho_{i})$, $i=0,1$, be two zero-mean Gaussian measures with isotropic covariance function $\rho_{i}$ and associated spectral density $\widehat{\rho}_{i}$, $i=0,1$, as defined through (1). Using results in Skorokhod and Yadrenko, (1973) and Ibragimov and Rozanov, (1978), Stein, (2004) has shown that, if for some $a>0$, $\widehat{\rho}_{0}(z)z^{a}$ is bounded away from 0 and $\infty$ as $z\to\infty$, and for some finite and positive $c$,
$\int_{c}^{\infty}z^{d-1}\;\left\{\frac{\widehat{\rho}_{1}(z)-\widehat{\rho}_{0}(z)}{\widehat{\rho}_{0}(z)}\right\}^{2}\;{\rm d}z<\infty,$ (31)
then for any bounded subset $D\subset\mathbb{R}^{d}$, $P(\rho_{0})$ and $P(\rho_{1})$ are equivalent Gaussian measures on the paths of $Z(\boldsymbol{s}),\boldsymbol{s}\in D$. Then (31) is the key tool to prove Theorem 1.
We first need the results in Lemma 2 hereunder. They follow from Zastavnyi, (2006), Chernih and Hubbert, (2014), and from standard properties of Fourier transforms and give asymptotic expansion of the spectral density of the $\mathcal{GH}$ model. Their proofs are thus omitted. For two given functions $g_{1}(x)$ and $g_{2}(x)$, with $g_{1}(x)\asymp g_{2}(x)$ we mean that there exist two constants $c$ and $C$ such that $0<c<C<\infty$ and $c|g_{2}(x)|\leq|g_{1}(x)|\leq C|g_{2}(x)|$ for each $x$.
Lemma 2.
Let $\widehat{{\cal GH}}_{\delta,\beta,\gamma,a}(z)$ be the function defined at Equation (11). Then, for $\delta>d/2$, $2(\beta-\delta)(\gamma-\delta)\geq\delta$ and $12(\beta+\gamma)\geq 6\delta+$ 1
  1. 1.
    $\widehat{{\cal GH}}_{\delta,\beta,\gamma,a}(z)=\sigma^{2}Ka^{d}\Big{[}c_{3}(za)^{-2\delta}\big{\{}1+\mathcal{O}(z^{-2})\big{\}}$
        $+\,c_{4}(za)^{\frac{1}{2}+\delta-\beta-\gamma}\big{\{}\cos(za+c_{5})+\mathcal{O}(z^{-1})\big{\}}\Big{]}$,  for $z\to\infty$;
  2. 2.
    $\widehat{{\cal GH}}_{\delta,\beta,\gamma,a}(z)\asymp z^{-2\delta}$,  for $z\to\infty,$
where $c_{3}=\frac{2^{2\delta}\Gamma(\beta)\Gamma(\gamma)}{\Gamma(\gamma-\delta)\Gamma(\beta-\delta)}$, $c_{4}=\frac{\pi^{-\frac{1}{2}}\Gamma(\beta)\Gamma(\gamma)}{2^{\frac{1}{2}+\delta-\beta-\gamma}\Gamma(\delta)}$ and $c_{5}=\frac{\pi}{2}(\frac{1}{2}+\delta-\beta-\gamma)$.
We are now ready to prove the Theorem. Let us start with the sufficient part of the assertion. From Lemma  2 (Point 2), we have:
$c_{1}\leq z^{2\delta}\widehat{{\cal GH}}_{\delta,\beta_{i},\gamma_{i},a_{i}}\leq c_{2},\quad i=0,1.$
We then proceed by direct construction, and, using Lemma 2 (Points 1 and 2) and (31) we find that, as $z\to\infty$,
$\Big{|}\frac{\sigma_{1}^{2}\widehat{\mathcal{GH}}_{\delta,\beta_{1},\gamma_{1},a_{1}}(z)-\sigma_{0}^{2}\widehat{\mathcal{GH}}_{\delta,\beta_{0},\gamma_{0},a_{0}}(z)}{\sigma_{0}^{2}\widehat{\mathcal{GH}}_{\delta,\beta_{0},\gamma_{0},a_{0}}(z)}\Big{|}\leq\Big{|}\sigma_{1}^{2}K_{1}a_{1}^{d}\Big{[}c_{3}^{1}(za_{1})^{-2\delta}\big{\{}1+\mathcal{O}(z^{-2})\big{\}}$
$\displaystyle+\,c_{4}^{1}(za_{1})^{\lambda_{1}}\big{\{}\cos(za_{1}+c_{5}^{1})+\mathcal{O}(z^{-1})\big{\}}\Big{]}-\sigma_{0}^{2}K_{0}a_{0}^{d}\Big{[}c_{3}^{0}(za_{0})^{-2\delta}\big{\{}1+\mathcal{O}(z^{-2})\big{\}}$
$+\,c_{4}^{0}(za_{0})^{\lambda_{0}}\big{\{}\cos(za+c_{5}^{0})+\mathcal{O}(z^{-1})\big{\}}\Big{]}\Big{|}\frac{z^{2\delta}}{c_{1}}$
$\leq\frac{1}{c_{1}}\Big{|}\big{[}c_{3}^{1}\sigma_{1}^{2}a_{1}^{d-2\delta}K_{1}-c_{3}^{0}\sigma_{0}^{2}a_{0}^{d-2\delta}K_{0}\big{]}\{1+\mathcal{O}(z^{-2})\}+\big{[}c_{4}^{1}z^{\lambda_{1}+2\delta}\sigma_{1}^{2}a_{1}^{\lambda_{1}+d}K_{1}\cos(a_{1}z-c_{5}^{1})$
$-c_{4}^{0}z^{\lambda_{0}+2\delta}\sigma_{0}^{2}a_{0}^{\lambda_{0}+d}K_{0}\cos(a_{0}z-c_{5}^{0})\big{]}+\mathcal{O}(z^{-1})\{c_{4}^{1}z^{\lambda_{1}+2\delta}\sigma_{1}^{2}a_{1}^{\lambda_{1}+d}K_{1}-c_{4}^{0}z^{\lambda_{0}+2\delta}\sigma_{0}^{2}a_{0}^{\lambda_{0}+d}K_{0}\}\Big{|},$
where $c_{3}^{i}=\frac{2^{2\delta}\Gamma(\beta_{i})\Gamma(\gamma_{i})}{\Gamma(\gamma_{i}-\delta)\Gamma(\beta_{i}-\delta)}$, $c_{4}^{i}=\frac{\pi^{-\frac{1}{2}}\Gamma(\beta_{i})\Gamma(\gamma_{i})}{2^{\frac{1}{2}+\delta-\beta_{i}-\gamma_{i}}\Gamma(\delta)}$,$c_{5}^{i}=\frac{\pi}{2}(\frac{1}{2}+\delta-\beta_{i}-\gamma_{i})$ and $\lambda_{i}=\frac{1}{2}+\delta-\beta_{i}-\gamma_{i}$, $i=0,1$.
Let us now write
$A(z)$ $=\big{[}c_{3}^{1}\sigma_{1}^{2}a_{1}^{d-2\delta}K_{1}-c_{3}^{0}\sigma_{0}^{2}a_{0}^{d-2\delta)}K_{0}\big{]}\{1+\mathcal{O}(z^{-2})\},$
$B(z)$ $=\big{[}c_{4}^{1}z^{\lambda_{1}+2\delta}\sigma_{1}^{2}a_{1}^{\lambda_{1}+d}K_{1}\cos(a_{1}z-c_{5}^{1})-c_{4}^{0}z^{\lambda_{0}+2\delta}\sigma_{0}^{2}a_{0}^{\lambda_{0}+d}K_{0}\cos(a_{0}z-c_{5}^{0})\big{]}$
$D(z)$ $=\mathcal{O}(z^{-1})\{c_{4}^{1}z^{\lambda_{1}+2\delta}\sigma_{1}^{2}a_{1}^{\lambda_{1}+d}K_{1}-c_{4}^{0}z^{\lambda_{0}+2\delta}\sigma_{0}^{2}a_{0}^{\lambda_{0}+d}K_{0}\}.$
Then, for some positive and finite $c$, Condition (31) can be written as
$\int_{c}^{\infty}z^{d-1}\bigg{(}\frac{\sigma_{1}^{2}\widehat{\mathcal{GH}}_{\delta,\beta_{1},\gamma_{1},a_{1}}(z)-\sigma_{0}^{2}\widehat{\mathcal{GH}}_{\delta,\beta_{0},\gamma_{0},a_{0}}(z)}{\sigma_{0}^{2}\widehat{\mathcal{GH}}_{\delta,\beta_{0},\gamma_{0},a_{0}}(z)}\biggr{)}^{2}{\rm d}z<\infty,$ (32)
being equivalent to
$\int_{c}^{\infty}z^{d-1}\big{(}A(z)+B(z)+D(z)\big{)}^{2}{\rm d}z<\infty.$
Now, we have that $A(z)$ is identically equal to zero thanks to our condition. We claim that the partial integral that depends of $B(z)$ and $D(z)$ is finite if $\frac{d+1}{2}+3\delta\leq\min(\beta_{1}+\gamma_{1},\beta_{0}+\gamma_{0})$. In fact, we have, for $z\to\infty$,
$|B(z)|$ $\leq c_{4}^{1}z^{\lambda_{1}+2\delta}\sigma_{1}^{2}a_{1}^{\lambda_{1}+d}K_{1}+c_{4}^{0}z^{\lambda_{0}+2\delta}\sigma_{0}^{2}a_{0}^{\lambda_{0}+d}K_{0}\leq c_{6}\min\{z^{\lambda_{1}+2\delta},z^{\lambda_{0}+2\delta}\}$
and
$|D(z)|$ $\leq\mathcal{O}(z^{-1})\{c_{4}^{1}z^{\lambda_{1}+2\delta}\sigma_{1}^{2}a_{1}^{\lambda_{1}+d}K_{1}+c_{4}^{0}z^{\lambda_{0}+2\delta}\sigma_{0}^{2}a_{0}^{\lambda_{0}+d}K_{0}\}$
$\leq c_{7}\min\{z^{\lambda_{1}+2\delta},z^{\lambda_{0}+2\delta}\},$
with $c_{i}$ being positive and finite constants, $i=6,7$. Thus, if $\frac{\sigma_{0}^{2}L_{0}}{a_{0}^{2\delta-d}}=\frac{\sigma_{1}^{2}L_{1}}{a_{1}^{2\delta-d}}$, we have
$\int_{c}^{\infty}$ $z^{d-1}\left(\frac{\widehat{\mathcal{GH}}_{\delta,\beta_{1},\gamma_{1},a_{1}}(z)-\widehat{\mathcal{GH}}_{\delta,\beta_{0},\gamma_{0},a_{0}}(z)}{\widehat{\mathcal{GH}}_{\delta,\beta_{0},\gamma_{0},a_{0}}(z)}\right)^{2}{\rm d}z$ (33)
$\leq\int_{c}^{\infty}z^{d-1}(B(z)+D(z))^{2}{\rm d}z$
$\leq\int_{c}^{\infty}z^{d-1}\Big{(}B(z)^{2}+D(z)^{2}+2B(z)D(z)\Big{)}{\rm d}z.$
We conclude that (33) is finite if $\frac{d+1}{2}+3\delta\leq\min(\beta_{1}+\gamma_{1},\beta_{0}+\gamma_{0})$. The sufficient part of our claim is thus proved.
The necessary part follows the proof of Zhang, (2004). We suppose that $\frac{d+1}{2}+3\delta\leq\min(\beta_{1}+\gamma_{1},\beta_{0}+\gamma_{0})$, $\frac{\sigma_{0}^{2}L_{0}}{a_{0}^{2\delta-d}}\neq\frac{\sigma_{1}^{2}L_{1}}{a_{1}^{2\delta-d}}$ and let $\sigma_{2}^{2}=\sigma_{0}^{2}L_{0}\left(\frac{a_{1}}{a_{0}}\right)^{2\delta-d}$. Then $\mathcal{GH}_{\delta,\beta_{0},\gamma_{0},a_{0}}$ and $\mathcal{GH}_{\delta,\beta_{1},\gamma_{1},a_{2}}$ induce two equivalent Gaussian measures. This, we need to show that $\mathcal{GH}_{\delta,\beta_{1},\gamma_{1},a_{2}}$ and $\mathcal{GH}_{\delta,\beta_{1},\gamma_{1},a_{1}}$ define two orthogonal Gaussian measures. The rest of the proof follows the same arguments in Zhang, (2004). ∎
A.2 Proof of Theorem 2
Proof.
Let us consider the $\mathcal{GH}_{\delta,\delta+\frac{\mu}{2},\delta+\frac{\mu}{2}+l,a}$ model and, for the moment, let us assume that $l$ is a real number. The sufficient conditions (a), (b), (c) given in Section 3, under the parameterization (15) are:
  1. (a)
    $\kappa>-\frac{1}{2}$;
  2. (b)
    $\mu(\mu+2l)\geq d+2\kappa+1$;
  3. (c)
    $\mu+l\geq\frac{d}{2}+\kappa+1$.
Solving the previous system of inequalities, we derive the following sufficient conditions:
  1. 1.
    For $d=1$
    • a)
      $\mu\geq\sqrt{2\kappa+l^{2}+2}-l$ and $l\leq-\frac{(1+2\kappa)}{2}$
    • b)
      $\mu\geq\delta-(l-\frac{1}{2})$ and $-\frac{(1+2\kappa)}{2}<l\leq\frac{(1+2\kappa)}{2}$.
    • c)
      $\mu\geq\sqrt{2\kappa+l^{2}+2}-l$ and $l>\frac{(1+2\kappa)}{2}$.
  2. 2.
    For $d>1$
    • a)
      $\mu\geq\sqrt{2\kappa+l^{2}+d+1}-l$ and $l<-\frac{(d+2\kappa)}{2}$
    • b)
      $\mu\geq\delta-(l-\frac{1}{2})$ and $-\frac{(d+2\kappa)}{2}\leq l\leq\frac{(d+2\kappa)}{2}$.
    • c)
      $\mu\geq\sqrt{2\kappa+l^{2}+d+1}-l$ and $l>\frac{(d+2\kappa)}{2}$.
Necessary conditions can be obtained following Cho and Yun, (2018, Theorem 6.1), who show that if $\mathstrut{}_{1}F_{2}\left(a;b;c;-\frac{x^{2}}{4}\right)\geq 0$, then $b>a$, $c>a$, $b+c\geq 3a+\frac{1}{2}$. Using the spectral density (11), under the parameterization (15), we have
  1. (a)
    $\mu>0$;
  2. (b)
    $\mu>-2l$;
  3. (c)
    $\mu+l\geq\frac{d}{2}+\kappa+1$.
Solving the system of inequalities (under the restriction $\kappa>-\frac{1}{2}$) for $d\geq 1$, we derive the following necessary conditions:
  1. 1.
    For $d=1$
    • a)
      $\mu>-2l$ and $l\leq-\frac{(3+2\kappa)}{2}$
    • b)
      $\mu\geq\delta-(l-\frac{1}{2})$ and $-\frac{(3+2\kappa)}{2}<l\leq\frac{(3+2\kappa)}{2}$.
    • d)
      $\mu>0$ and $l>\frac{(3+2\kappa)}{2}$.
  2. 2.
    For $d>1$
    • a)
      $\mu>-2l$ and $l<-\frac{(d+2\kappa+2)}{2}$
    • c)
      $\mu\geq\delta-(l-\frac{1}{2})$ and $-\frac{(d+2\kappa+2)}{2}\leq l<\frac{(d+2\kappa+2)}{2}$.
    • d)
      $\mu>0$ and $l\geq\frac{(d+2\kappa+2)}{2}$.
Putting together the necessary and sufficient conditions for the cases $d=1$ and $d>1$, we obtain the following conditions:
  1. 1.
    For $d=1$
    • a)
      $\mu\geq\sqrt{2\kappa+l^{2}+2}-l$ and $l\leq-\frac{(1+2\kappa)}{2}$
    • b)
      $\mu\geq\delta-(l-\frac{1}{2})$ and $-\frac{(1+2\kappa)}{2}<l\leq\frac{(1+2\kappa)}{2}$.
    • c)
      $\mu\geq\sqrt{2\kappa+l^{2}+2}-l$ and $l>\frac{(1+2\kappa)}{2}$.
  2. 2.
    For $d>1$
    • a)
      $\mu\geq\sqrt{2\kappa+l^{2}+d+1}-l$ and $l<-\frac{(d+2\kappa)}{2}$
    • b)
      $\mu\geq\delta-(l-\frac{1}{2})$ and $-\frac{(d+2\kappa)}{2}\leq l\leq\frac{(d+2\kappa)}{2}$.
    • c)
      $\mu\geq\sqrt{2\kappa+l^{2}+d+1}-l$ and $l>\frac{(d+2\kappa)}{2}$.
We observe that the necessary and sufficient conditions are symmetric around $l=0$, since $\mathcal{GH}_{\delta,\beta,\gamma,a}(x)=\mathcal{GH}_{\delta,\gamma,\beta,a}(x)$. Specifically, for $d=1$, conditions a) and c) are symmetric when $l<-\frac{(1+2\kappa)}{2}$ and $l>\frac{(1+2\kappa)}{2}$, respectively, while condition b) is symmetric when $-\frac{(d+2\kappa)}{2}<l\leq 0$ and $0\leq l\leq\frac{(d+2\kappa)}{2}$. These same symmetry patterns are also observed for the case $d>1$.
Using the symmetry of the conditions, we put together the necessary and sufficient conditions only for $l\geq 0$ obtaining the main result, that is for $\kappa>-\frac{1}{2}$ the model is valid if and only if
  1. 1.
    $\mu\geq\delta-(l-\frac{1}{2})$ and $0\leq l\leq\frac{(d+2\kappa)}{2}$
  2. 2.
    $\mu\geq\sqrt{2\kappa+l^{2}+d+1}-l$ and $l>\frac{(d+2\kappa)}{2}$.
A.3 Proof of Theorem 3
Proof.
The result stems from formula 2.21.1.11 of Prudnikov et al., (1990) and the expressions (12) and (16) of the hypergeometric correlation and of the integral range.
A.4 Proof of Theorem 4
Hereafter, without loss of generality consider a unit compact support $a=1$. Let us define $LA_{\kappa,d}(\mu,l):=log(A_{\kappa,\mu,l,d,1})$ where
$A_{\kappa,\mu,l,d,1}=\frac{\pi^{d/2}\Gamma((d+1)/2+\kappa)\Gamma\left(\frac{\mu+1}{2}+\kappa\right)\Gamma\left(\frac{\mu+1}{2}+\kappa+l\right)}{2^{d}\Gamma(\kappa+0.5)\Gamma\left(\frac{\mu+1}{2}+\kappa+d/2\right)\Gamma\left(\frac{\mu+1}{2}+\kappa+l+d/2\right)}.$
is the integral range.
[…]
  1. 1.
    $\mu\geq\delta-l+0.5\quad$ if $\quad 0\leq l\leq d/2+\kappa$
  2. 2.
    $\mu\geq\sqrt{2\kappa+l^{2}+d+1}-l\quad$ if $\quad l>d/2+\kappa$.
This is a maximization problem over a bounded interval. Hereafter, with $\psi(x)=\frac{\text{d}log\Gamma(x)}{\text{d}x}$ we denote the digamma function which is strictly increasing for $x>0$. The fist derivative of $LA_{\kappa,d}(\mu,l)$ with respect to $\mu$ is given by
$\frac{\text{d}LA_{\kappa,d}(\mu,l)}{\text{d}\mu}=0.5[\psi(x)-\psi(x+d/2)+\psi(x+l)-\psi(x+l+d/2)],$
where $x=\frac{\mu+1}{2}+\kappa$. It is a negative function since $\psi(\cdot)$ is strictly increasing and this implies that the function $LA_{\kappa,d}(\mu,l)$ is monotone non-increasing as a function of $\mu$. As a consequence the maximum is achieved by $LA_{\kappa,d}(\mu,l)$ on the lower boundary of region 1) and region 2), that is
  1. 1.
[…]
  2. 2.
[…]
where $\mu_{1}(l)=\delta-l+0.5$ and $\mu_{2}(l)=\sqrt{2\kappa+l^{2}+d+1}-l$. This proves Theorem 4.
A.5 Proof of Theorem 5
We focus on the functions $LA_{\kappa,d}(\mu_{i}(l),l)$ for $i=1,2$ defined on the proof of Theorem 4 and the goal is to find the maximum with respect to $l$.
If $i=1$, the derivative of $LA_{\kappa,d}(\mu_{1}(l),l)$ for $0\leq l\leq d/2+\kappa$ is given by
$\frac{\text{d}LA_{\kappa,d}(\mu_{1}(l),l)}{\text{d}l}=0.5[\psi(x+l/2)-\psi(x-l/2)+\psi(x-l/2+d/2)-\psi(x+l/2+d/2)],$
where $x=\frac{d}{4}+\frac{3\kappa}{2}+1$. It can be shown that this function is non-negative using the fact that the derivative of the digamma function $\frac{\text{d}\psi(x)}{\text{d}x}$ is a decreasing function. This implies that the maximum of $LA_{\kappa,d}(\mu_{1}(l),l)$ is achieved on the upper bound of the parametric space, that is $l=d/2+\kappa$.
Using similar arguments, if $i=2$ it can be shown that the derivative of $LA_{\kappa,d}(\mu_{2}(l),l)$ for $l>d/2+\kappa$ is non-positive. As a consequence the maximum is achieved on the lower bound of the parametric space, that is when $l=d/2+\kappa+\epsilon$ where $\epsilon$ an arbitrary small positive constant.
Finally we can prove that $LA_{\kappa,d}(\mu_{1}(d/2+\kappa),d/2+\kappa))=LA_{\kappa,d}(1,d/2+\kappa)$ is greater than $LA_{\kappa,d}(\mu_{2}(d/2+\kappa+\epsilon),d/2+\kappa+\epsilon))=LA_{\kappa,d}(1-g(\epsilon),d/2+\kappa+\epsilon)$ where $g(\epsilon)$ is a positive constant.
A.6 Proof of Theorem 6
Proof.
Using (11) and Equation (25) in Emery and Alegría, (2022), the Fourier transform of the right-hand side member of (22) is found to be
$\begin{split}&\frac{\Gamma(\delta+\frac{\mu-d}{2})\Gamma(\delta+\frac{\mu-d}{2}+l)\Gamma(\delta)a^{d}}{2^{d}\pi^{\frac{d}{2}}\Gamma\left(\delta-\frac{d}{2}\right)\Gamma(\delta+\frac{1}{2})\Gamma(2\delta)\Gamma(\frac{\mu-1}{2})\Gamma(\frac{\mu-1-d}{2}-\kappa+l)}\\ &\times\int_{0}^{1}\int_{0}^{1}u^{\delta-\frac{1}{2}}(1-u)^{\frac{\mu-3}{2}}v^{d+2\kappa}(1-v)^{\frac{\mu-d-3}{2}-\kappa+l}\mathstrut_{1}F_{2}\Big{(}\delta;\delta+\frac{1}{2},2\delta;-\frac{uv(za)^{2}}{4}\Big{)}{\rm d}u{\rm d}v\\ &=\frac{\Gamma(\delta)\Gamma(\delta+\frac{\mu-d}{2})\Gamma(\delta+\frac{\mu-d}{2}+l)a^{d}}{2^{d}\pi^{\frac{d}{2}}\Gamma\left(\delta-\frac{d}{2}\right)\Gamma(\delta+\frac{\mu}{2})\Gamma(\delta+\frac{\mu}{2}+l)}\mathstrut_{1}F_{2}\Big{(}\delta;\delta+\frac{\mu}{2},\delta+\frac{\mu}{2}+l;-\frac{(za)^{2}}{4}\Big{)}\\ &=\widehat{\mathcal{GH}}_{\delta,\delta+\frac{\mu}{2},\delta+\frac{\mu}{2}+l,a}(z),\quad z\geq 0.\end{split}$
The claim of the theorem is obtained by inverse Fourier transformation. ∎
A.7 Proof of Theorem 7
If $d=1$ and $\kappa=k$ is a non-negative integer, using formula 7.3.1.70 of Prudnikov et al., (1990) and applying the identity
$P_{a}^{b}(x)=P^{b}_{-a-1}(x),$
where $P_{a}^{b}(x)$ denotes the associated Legendre function (of the first kind) with degree $a$ and order $b$ (see Abramowitz and Stegun, (1970, 8.3.1)), one can rewrite (21) as
$\mathcal{H}_{k,\mu,a,1}(x)$ $=$ $\frac{\Gamma(\frac{\mu}{2}+2k+1)}{\Gamma(k+\frac{1}{2})}\sqrt{\pi}\left(1-\frac{x^{2}}{a^{2}}\right)^{\frac{\mu}{2}+k}\left(\frac{x}{a}\right)^{k}P^{-\frac{\mu}{2}-k}_{k}\left(\frac{a^{2}+x^{2}}{2ax}\right).$
The claim of the Theorem then derives from the following result, which is valid for $a$ being a non-negative integer and $z>1$:
$P^{-b}_{a}(z)=\frac{(z-1)^{b/2}}{(z+1)^{b/2}}\sum_{j=0}^{a}\frac{(j+a)!(0.5(z-1))^{j}}{j!\Gamma(j+b+1)(a-j)!}.$ (35)
A.8 Proof of Theorem 8
Proof.
For $\mu\geq 1$ and $\kappa=0$, applying formula 7.3.1.41 from Prudnikov et al., (1990) to Equation (21), one gets
$\mathcal{H}_{0,\mu,a,d}(x)=\frac{\Gamma(\frac{1+\mu}{2})\Gamma(\frac{1+\mu+d}{2})}{\sqrt{\pi}}2^{\mu+\frac{d-1}{2}}\left(1-\frac{x^{2}}{a^{2}}\right)_{+}^{\frac{\mu}{2}+\frac{d-1}{4}}\,P_{\frac{d-1}{2}}^{\frac{1-d}{2}-\mu}\left(\frac{x}{a}\right).$
The claim of the Theorem stems from the fact that, for $m$ being an integer, the associated Legendre function of degree $m$ is a terminating series:
$P_{m}^{\eta}(z)=\left(\frac{1+z}{1-z}\right)^{\frac{\eta}{2}}\sum\limits_{n=0}^{m}\frac{(-m)_{n}(m+1)_{n}}{n!\Gamma(n-\eta+1)}\left(\frac{1-z}{2}\right)^{n},\quad 0<z<1.$ (36)
A.9 Proof of Theorem 9
Proof.
For notation convenience we consider the notation ${\cal GH}_{\delta,\beta,\gamma,a}$ for the $\mathcal{H}_{\kappa,\mu,a,d}$ model, recalling that (see 20) if
$\delta=\frac{d+1}{2}+\kappa,\quad\beta=\delta+\frac{\mu}{2},\quad\gamma=\delta+\frac{\mu}{2}+d/2+\kappa,$ (37)
then $\mathcal{H}_{\kappa,\mu,a,d}=\mathcal{GH}_{\delta,\beta,\gamma,a}.$
In order to prove Theorem 9, as in the proof of Theorem 1, we use (31) to find conditions such that, for some positive and finite $c$,
$\int_{c}^{\infty}z^{d-1}\bigg{(}\frac{\sigma_{1}^{2}\widehat{{\cal GH}}_{\delta,\beta,\gamma,a}(z)-\sigma_{0}^{2}\widehat{{\cal MT}}_{\nu,\alpha}(z)}{\sigma_{0}^{2}\widehat{{\cal MT}}_{\nu,\alpha}(z)}\biggr{)}^{2}{\rm d}z<\infty.$ (38)
It is known that $\sigma_{0}^{2}\widehat{{\cal MT}}_{\nu,\alpha})(z)z^{\tau}$ is bounded away from 0 and $\infty$ as $z\to\infty$ for some $\tau>0$ (Zhang,, 2004). Let $\lambda$ be a parameter such as $\lambda=\frac{1}{2}+\delta-\beta-\gamma$. Using Points 1 of Lemma 2 , we have, as $z\to\infty$,
$\Bigg{|}\frac{\sigma_{1}^{2}\widehat{{\cal GH}}_{\delta,\beta,\gamma,a}(z)-\sigma_{0}^{2}\widehat{{\cal MT}}_{\nu,\alpha}(z)}{\sigma_{0}^{2}\widehat{{\cal MT}}_{\nu,\alpha}(z)}\Bigg{|}$
$=$ $\Bigg{|}\frac{\sigma_{1}^{2}Ka^{d}\Gamma(\nu)}{\Gamma(\nu+d/2)\sigma_{0}^{2}\alpha^{-2\nu}\pi^{-\frac{d}{2}}}\Big{[}c_{3}(za)^{-2\delta}\big{\{}1+\mathcal{O}(z^{-2})\big{\}}+\,c_{4}(za)^{\lambda}\big{\{}\cos(za+c_{5})+\mathcal{O}(z^{-1})\big{\}}\Big{]}(\alpha^{-2}+z^{2})^{\nu+\frac{d}{2}}-1\Bigg{|}$
$=$ $\Bigg{|}\frac{\sigma_{1}^{2}Ka^{d}\Gamma(\nu)\pi^{\frac{d}{2}}}{\Gamma(\nu+d/2)\sigma_{0}^{2}\alpha^{-2\nu}}\Big{[}c_{3}(za)^{-2\delta}\big{\{}1+\mathcal{O}(z^{-2})\big{\}}+\,c_{4}(za)^{\lambda}\big{\{}\cos(za+c_{5})+\mathcal{O}(z^{-1})\big{\}}\Big{]}\frac{(({\alpha z})^{-2}+1)^{\nu+\frac{d}{2}}}{z^{-2\nu-d}}-1\Bigg{|}$
$=$ $\Bigg{|}\frac{\sigma_{1}^{2}Ka^{d}\Gamma(\nu)\pi^{\frac{d}{2}}}{\Gamma(\nu+d/2)\sigma_{0}^{2}\alpha^{-2\nu}}\Big{[}c_{3}(za)^{-2\delta}\big{\{}1+\mathcal{O}(z^{-2})\big{\}}+\,c_{4}(za)^{\lambda}\big{\{}\cos(za+c_{5})+\mathcal{O}(z^{-1})\big{\}}\Big{]}\times$
$\quad\frac{\Big{[}1+(\nu+d/2)(\alpha z)^{-2}+\mathcal{O}(z^{-2})\Big{]}}{z^{-2\nu-d}}-1\Bigg{|}.$
$=$ $\Bigg{|}w_{1}z^{2\nu+d-2\delta}\big{\{}1+\mathcal{O}(z^{-2})\big{\}}\Big{[}1+(\nu+d/2)(\alpha z)^{-2}+\mathcal{O}(z^{-2})\Big{]}$
$\quad+w_{2}z^{\lambda+2\nu+d}\Big{[}1+(\nu+d/2)(\alpha z)^{-2}+\mathcal{O}(z^{-2})\Big{]}\big{\{}\cos(\beta z-c_{5})+\mathcal{O}(z^{-1})\big{\}}-1\Bigg{|},$
where $w_{1}=\frac{\sigma_{1}^{2}Ka^{d-2\delta}\Gamma(\nu)c_{3}}{\Gamma(\nu+d/2)\sigma_{0}^{2}\alpha^{-2\nu}\pi^{-\frac{d}{2}}}$, $w_{2}=w_{1}c_{4}a^{\lambda}/c_{3}$.
Since $z^{2\nu+d}\big{[}(\nu+d/2)(\alpha z)^{-2}+\mathcal{O}(z^{-2})\big{]}=\mathcal{O}(z^{2\nu+d-2})$, we have
$\int_{c}^{\infty}z^{d-1}\Bigg{|}\frac{\sigma_{1}^{2}\widehat{{\cal GH}}_{\delta,\beta,\gamma,a}(z)-\sigma_{0}^{2}\widehat{{\cal MT}}_{\nu,\alpha}(z)}{\sigma_{0}^{2}\widehat{{\cal MT}}_{\nu,\alpha}(z)}\Bigg{|}^{2}{\rm d}z$
$=\int_{c}^{\infty}z^{d-1}\Bigg{|}w_{1}z^{-2\delta}\mathcal{O}(z^{2\nu+d-2})+\big{\{}w_{1}z^{2\nu+d-2\delta}-1\big{\}}+w_{1}z^{-2\delta}$
×$\quad\times\big{\{}\mathcal{O}(z^{2\nu+d-2})+\mathcal{O}(z^{2\nu+d-4})\big{\}}+w_{2}z^{\lambda}\big{\{}\mathcal{O}(z^{2\nu+d-2})+z^{2\nu+d}\big{\}}\big{\{}\cos(\beta z-c_{5}^{\boldsymbol{\varsigma}})$
$\quad+\mathcal{O}(z^{-1})\big{\}}\Bigg{|}^{2}{\rm d}z.$
For assessing the last integral, the following is relevant:
  • (i)
    $w_{1}z^{2\nu+d-2\delta}-1=0$ if $\delta=\nu+\frac{d}{2}$ and $w_{1}=1$.
  • (ii)
    $\int_{c}^{\infty}z^{d-1}\big{(}w_{1}z^{-2\delta}\mathcal{O}(z^{2\nu+d-2})\big{)}^{2}{\rm d}z<\infty$ if $d=1,2,3$ and $2\nu+d=2\delta$.
  • (iii)
    $\int_{c}^{\infty}z^{d-1}\Big{(}w_{1}z^{-2\delta}\big{\{}\mathcal{O}(z^{2\nu+d-2})+\mathcal{O}(z^{2\nu+d-4})\big{\}}\Big{)}^{2}{\rm d}z<\infty$ if $d=1,2,3$ and $\nu=\kappa+\frac{1}{2}$.
  • (iv)
    $\int_{c}^{\infty}z^{d-1}\Big{(}w_{2}z^{\lambda}\big{\{}\mathcal{O}(z^{2\nu+d-2})+z^{2\nu+d}\big{\}}\big{\{}\cos(\beta z-c_{5}^{\boldsymbol{\varsigma}})+\mathcal{O}(z^{-1})\big{\}}\Big{)}^{2}{\rm d}z<\infty$ if $\nu=\kappa+1/2$.
  • (v)
    $\int_{c}^{\infty}z^{d-1}\big{(}w_{1}z^{-2\delta}\mathcal{O}(z^{2\nu+d-2})\big{)}\big{(}w_{1}z^{-2\delta}\mathcal{O}(z^{-2})(\mathcal{O}(z^{2\nu+d-2})+z^{2\nu+d})\big{)}{\rm d}z<\infty$ if $d=1,2,3$ and $\nu=\kappa+1/2$.
  • (vi)
    $\int_{c}^{\infty}z^{d-1}\big{(}w_{1}z^{-2\delta}\mathcal{O}(z^{2\nu+d-2})\big{)}\big{(}w_{2}z^{\lambda}\big{\{}\mathcal{O}(z^{2\nu+d-2})+z^{2\nu+d}\big{\}}\big{\{}\cos(\beta z-c_{5})+\mathcal{O}(z^{-1})\big{\}}\big{)}{\rm d}z<\infty$ if $d=1,2,3$ and $\nu=\kappa+1/2$.
  • (vii)
    $\int_{c}^{\infty}z^{d-1}\big{(}w_{1}z^{-2\delta}\mathcal{O}(z^{-2})\big{\{}\mathcal{O}(z^{2\nu+d-2})+z^{2\nu+d}\big{\}}\big{)}\big{(}w_{2}z^{\lambda}\big{\{}\mathcal{O}(z^{2\nu+d-2})+z^{2\nu+d}\big{\}}$$\times\big{\{}\cos(\beta z-c_{5})+\mathcal{O}(z^{-1})\big{\}}\big{)}{\rm d}z<\infty$ if $d=1,2,3$ and $\nu=\kappa+1/2$.
This allows us to conclude that, for a given $\kappa>-\frac{1}{2}$, if $w_{1}=1$, $\nu=\kappa+1/2$ and $d=1,2,3$ then (38) holds and thus $P(\sigma^{2}_{0}{\cal MT}_{\nu,\alpha})\equiv P(\sigma_{1}^{2}{\cal GH}_{\delta,\beta,\gamma,a})$. Condition $w_{1}=1$ is equivalent to
$\frac{\sigma_{0}^{2}}{\alpha^{2\nu}}=\frac{\sigma_{1}^{2}A}{a_{1}^{2\delta-d}},$
where $A=\frac{2^{2\delta-d}\Gamma(\beta-d/2)\Gamma(\gamma-d/2)}{\Gamma(\gamma-\delta)\Gamma(\beta-\delta)}$ and, taking into account the parameterization (37), the previous condition can be written as
$\frac{\sigma_{0}^{2}}{\alpha^{2\nu}}=\frac{\sigma_{1}^{2}}{a^{2\kappa+1}}B(\kappa,\mu,d),$
where $B(\kappa,\mu,d)$ is defined in (24).
A.10 Proof of Lemma 1
Proof.
The proof follows Kaufman and Shaby, (2013). Let $0<a_{1}<a_{2}$, with $a_{1},a_{2}\in I$. Then, for any $\boldsymbol{Z}_{n}$,
$\hat{\sigma}_{n}^{2}(a_{1})/a_{1}^{1+2\kappa}-\hat{\sigma}_{n}^{2}(a_{2})/a_{2}^{1+2\kappa}=\frac{1}{n}\boldsymbol{Z}_{n}^{\prime}(R_{n}(a_{1})^{-1}a_{1}^{-(1+2\kappa)}-R_{n}(a_{2})^{-1}a_{2}^{-(1+2\kappa)})\boldsymbol{Z}_{n}$
is nonnegative if the matrix $R_{n}(a_{1})^{-1}a_{1}^{-(1+2\kappa)}-R_{n}(a_{2})^{-1}a_{2}^{-(1+2\kappa)}$ is positive semidefinite and this happens if and only if the matrix $B=R_{n}(a_{2})a_{2}^{1+2\kappa}-R_{n}(a_{1})a_{1}^{1+2\kappa}$ with generic element
$B_{ij}=a_{2}^{1+2\kappa}\mathcal{H}_{\kappa,\mu,a_{2},d}(||\boldsymbol{s}_{i}-\boldsymbol{s}_{j}||)-a_{1}^{1+2\kappa}\mathcal{H}_{\kappa,\mu,a_{1},d}(||\boldsymbol{s}_{i}-\boldsymbol{s}_{j}||)$
is positive semidefinite. The last assertion holds if and only if the function
$f(z)=a_{2}^{1+2\kappa}\widehat{\mathcal{H}}_{\kappa,\mu,a_{2},d}(z)-a_{1}^{1+2\kappa}\widehat{\mathcal{H}}_{\kappa,\mu,a_{1},d}(z)\geq 0$ (39)
for all $z>0$. Here ^$\widehat{\mathcal{H}}_{\kappa,\mu,a,d}$ is the spectral density defined in (11) that is $\widehat{{\cal GH}}_{\delta,\beta,\gamma,a}$ under the parameterization (37). Condition (39) is true if and only if the function $f_{1}(t)=t^{1+2\kappa}\widehat{\mathcal{H}}_{\kappa,\mu,t,d}(z)$ is an increasing function with respect to $t>0$ for all $z>0$. Following the same arguments in Faouzi et al., (2020), $f_{1}(t)$ is increasing if and only if $g(t)=t^{1+2\kappa}\widehat{\mathcal{H}}_{\kappa,\mu,1,d}(t)$ is increasing and a sufficient condition is given by $\frac{dg(t)}{dt}>0$. Using contour integral representation of the $\mathstrut{}_{1}F_{2}\Big{(}\cdot;\cdot,\cdot;-t^{2}\Big{)}$ function and setting $\delta=\frac{d+1}{2}+\kappa$, $\beta=\delta+\frac{\mu}{2}$, $\gamma=\beta+\frac{d}{2}+\kappa$, and $M$ a positive constant not depending on $t$, the derivative is given by:
$\frac{dg(t)}{dt}$ $=\frac{d}{dt}\left[Mt^{2\kappa+1+d}\mathstrut_{1}F_{2}\Big{(}\delta;\beta,\gamma;-t^{2}\Big{)}\right]$ (40)
[…]
$=M2\delta t^{2\kappa+d}\mathstrut_{1}F_{2}\left(\delta+1;\beta,\gamma;-t^{2}\right)$
$=M2\delta t^{2\kappa+d}\mathstrut_{1}F_{2}\left(\delta_{1};\beta_{1},\gamma_{1};-t^{2}\right),$
where $\delta_{1}=\frac{d+1}{2}+(k+1)$, $\beta_{1}=\delta_{1}+\frac{\mu-2}{2}$, $\gamma_{1}=\beta_{1}+l$, and $l=\frac{d}{2}+\kappa$.
Since $M2\delta t^{2\kappa+d}$ is positive, $\frac{dg(t)}{dt}>0$ if and only if $\mathstrut{}_{1}F_{2}\left(\delta_{1};\beta_{1},\gamma_{1}+l;-t^{2}\right)>0$ Using point 1 of Theorem 2, this happens if and only if $\mu\geq 4$.
Appendix B Closed-form expressions of the hypergeometric model
In Euclidean spaces of odd dimension. When $d$ is an odd integer, $\kappa=k$ is a non-negative integer and $\mu$ is a positive integer, the hypergeometric model yields a truncated polynomial (Emery and Alegría,, 2022, Section 2.5.3). Specifically, if $\mu$ is odd, then
$\begin{split}\mathcal{H}_{k,\mu,a,d}(x)=\begin{cases}\frac{\Gamma(\frac{1}{2}-k)\Gamma(\frac{\mu+1}{2}+k)}{\Gamma(\frac{1-\mu-d}{2}-2k)}\bigg{[}\sum_{n=0}^{k+\frac{\mu-1}{2}}\frac{(-1)^{n}\Gamma(\frac{1-\mu-d}{2}-2k+n)}{\Gamma(\frac{1}{2}-k+n)\Gamma(\frac{\mu+1}{2}+k-n)n!}\left(\frac{x}{a}\right)^{2n}\\ +\sum_{n=0}^{k-1+\frac{\mu+d}{2}}\frac{(-1)^{n}\Gamma(n+1-\frac{\mu}{2})}{\Gamma(\frac{3}{2}+k+n)\Gamma(\frac{\mu+d}{2}+k-n)n!}\left(\frac{x}{a}\right)^{2n+2k+1}\bigg{]}\text{ if $x<a$}\\ 0\text{ if $x\geq a$,}\end{cases}\end{split}$
while, if $\mu$ is even,
$\begin{split}\mathcal{H}_{k,\mu,a,d}(x)=\begin{cases}\frac{\Gamma(\frac{1}{2}-k)\Gamma(\frac{\mu+d+1}{2}+2k)}{\Gamma(\frac{1-\mu}{2}-k)}\bigg{[}\sum_{n=0}^{2k+\frac{\mu+d-1}{2}}\frac{(-1)^{n}\Gamma(\frac{1-\mu}{2}-k+n)}{\Gamma(\frac{1}{2}-k+n)\Gamma(\frac{\mu+d+1}{2}+2k-n)n!}\left(\frac{x}{a}\right)^{2n}\\ +\sum_{n=0}^{\frac{\mu}{2}-1}\frac{(-1)^{n}\Gamma(n+1-\frac{\mu+d}{2}-k)}{\Gamma(\frac{3}{2}+k+n)\Gamma(\frac{\mu}{2}-n)n!}\left(\frac{x}{a}\right)^{2n+2k+1}\bigg{]}\text{ if $x<a$}\\ 0\text{ if $x\geq a$.}\end{cases}\end{split}$
The cases $k=0$, $\mu=1$ and $d=1$, $3$ and $5$ correspond to the well-known triangular, spherical and pentaspherical models (Matérn,, 1986), which behave linearly near the origin:
$\mathcal{H}_{0,1,a,1}(x)=1-\frac{x}{a},\quad x\leq a,$
$\mathcal{H}_{0,1,a,3}(x)=1-\frac{3x}{2a}+\frac{x^{3}}{2a^{3}},\quad x\leq a,$
$\mathcal{H}_{0,1,a,5}(x)=1-\frac{15x}{8a}+\frac{5x^{3}}{4a^{3}}-\frac{3x^{5}}{8a^{5}},\quad x\leq a,$
whereas the cases $d=3$, $\mu=1$ and $k=1$ or $2$ correspond to the so-called cubic and penta models (Chilès and Delfiner,, 2012), which are associated with Gaussian RFs that are once and twice mean-squared differentiable, respectively:
$\mathcal{H}_{1,1,a,3}(x)=1-\frac{7x^{2}}{a^{2}}+\frac{35x^{3}}{4a^{3}}-\frac{7x^{5}}{2a^{5}}+\frac{3x^{7}}{4a^{7}},\quad x\leq a,$
$\mathcal{H}_{2,1,a,3}(x)=1-\frac{22x^{2}}{3a^{2}}+\frac{33x^{4}}{a^{4}}-\frac{77x^{5}}{2a^{5}}+\frac{33x^{7}}{2a^{7}}-\frac{11x^{9}}{2a^{9}}+\frac{5x^{11}}{6a^{11}},\quad x\leq a.$
In the 2D plane. If $d=2$ and $\mu=1$, the following closed-form expressions are found.
  1. 1.
    When $\kappa=0$, formula 7.3.2.92 of Prudnikov et al., (1990) gives:
    $\mathcal{H}_{0,1,a,2}(x)=\frac{2}{\pi}\left(\arcsin\sqrt{\left(1-\frac{x^{2}}{a^{2}}\right)_{+}}-\frac{x}{a}\left(1-\frac{x^{2}}{a^{2}}\right)_{+}^{\frac{1}{2}}\right),$ (41)
    which corresponds to the circular correlation model (Chilès and Delfiner,, 2012, 2.46), which behaves linearly near the origin.
  2. 2.
    When $\kappa=\frac{1}{2}$, formula 7.3.2.99 of Prudnikov et al., (1990) gives:
    $\mathcal{H}_{\frac{1}{2},1,a,2}(x)=\left(2-\frac{x^{2}}{a^{2}}\right)\operatorname{arctanh}\sqrt{\left(1-\frac{x^{2}}{a^{2}}\right)_{+}}-\sqrt{\left(1-\frac{x^{2}}{a^{2}}\right)_{+}},$
    which is the montée (upgrading) (Matheron,, 1965, I.4.18) of order $1$ of the spherical correlation model in $\mathbb{R}^{3}$.
  3. 3.
    When $\kappa=1$, one obtains (Weisstein,, 2025):
    $\mathcal{H}_{1,1,a,2}(x)$ $=$ $\frac{2}{3\pi}\Bigg{\{}\frac{x}{a}\left(1-\frac{x^{2}}{a^{2}}\right)_{+}^{\frac{1}{2}}\left[15-4\left(1-\frac{x^{2}}{a^{2}}\right)_{+}\left(3-\frac{x^{2}}{a^{2}}\right)\right]$ (42)
    $+3\left[6\left(1-\frac{x^{2}}{a^{2}}\right)_{+}-5\right]\arcsin\sqrt{\left(1-\frac{x^{2}}{a^{2}}\right)_{+}}\Bigg{\}},$
    which is the montée of order $2$ of the Euclid’s hat model in $\mathbb{R}^{4}$ and is associated with a Gaussian RF that is mean-squared differentiable.