### v2

# Parsimonious Compactly Supported Covariance Models in the Gauss Hypergeometric Class: Identifiability, Reparameterizations, and Asymptotic Properties

Abstract
We study covariance functions in the Gauss hypergeometric ($\mathcal{GH}$) class, a flexible family that encompasses the Generalized Wendland ($\mathcal{GW}$) and Matérn ($\mathcal{MT}$) models. We derive sharp validity conditions, providing a complete characterization of the admissible parameter space, and show that the model exhibits structural identifiability issues under both increasing- and fixed-domain asymptotics.
To resolve this issue, we introduce a parsimonious compactly supported subclass selected via a maximum integral range criterion. The resulting hypergeometric model can be viewed as a structural refinement of the $\mathcal{GW}$ family and admits compact-support reparameterizations that recover the $\mathcal{MT}$ model as a limit case.
We further establish strong consistency and asymptotic normality of the maximum likelihood estimator of the associated microergodic parameter under fixed-domain asymptotics. Simulation experiments and a real-data application to climate data illustrate the finite-sample behavior and practical performance of the proposed model.
Keywords: Fixed domain asymptotics, Gaussian random fields, Generalized Wendland model, Integral range, Sparse matrices.
1 Introduction
Gaussian random fields (RFs) are widely used in spatial statistics, machine learning, numerical analysis, and computer experiments (Cressie, 1993; Stein, 1999; Banerjee et al., 2004; Rasmussen and Williams, 2006; Wendland, 2004; Santner et al., 2003). Under Gaussianity, the dependence structure is fully characterized by the covariance function, whose specification is central for modeling, prediction, and uncertainty quantification.
For RFs on Euclidean spaces, isotropic covariance (kernel) functions— depending only on Euclidean distance—provide a tractable and widely used modeling framework. Among parametric isotropic models, the Matérn ($\mathcal{MT}$) class is arguably the most prominent (Porcu et al., 2024), owing to its smoothness parameter, which continuously controls the mean-square differentiability of the sample paths. A compactly supported alternative is the generalized Wendland ($\mathcal{GW}$) family (Gneiting, 2002; Bevilacqua et al., 2019), which also allows for continuous control of smoothness while yielding sparse covariance matrices and substantial computational gains (Duff et al., 1989; Davis, 2006). Connections between $\mathcal{MT}$ and $\mathcal{GW}$ models have been extensively studied (Bevilacqua et al., 2019, 2022).
We focus on the Gauss hypergeometric ($\mathcal{GH}$) covariance model introduced by Emery and Alegría (2022). This highly flexible class includes both $\mathcal{MT}$ and $\mathcal{GW}$ as special or asymptotic cases and has been recognized as a prominent extension of the Matérn family (Porcu et al., 2024). While sufficient validity conditions were derived in Emery and Alegría (2022), a complete characterization of the parameter space and a detailed inferential analysis remain open.
This paper provides several theoretical and methodological contributions. First, we refine the validity conditions of the $\mathcal{GH}$ model. We show that the sufficient conditions of Emery and Alegría (2022) are also necessary in the parameter range where the two sets of conditions agree. Second, we clarify that the $\mathcal{GH}$ family is not identifiable in the classical parametric sense. The parameter-to-correlation mapping is not injective, and this structural non-identifiability must be distinguished from the well-known inferential limitations under fixed-domain asymptotics, where Gaussian measures may remain equivalent on bounded domains. In such settings, only microergodic parameter combinations are consistently estimable.
To resolve the structural non-identifiability, we introduce an optimality-driven parsimonious subclass of $\mathcal{GH}$. For fixed smoothness and compact support, we select the member that maximizes the isotropic integral range, that is, the total mass of the correlation over $\mathbb{R}^{d}$. Since the integral range coincides with the spectral density at zero frequency, this criterion selects the admissible model with maximal low-frequency spectral power, i.e., the strongest large-scale dependence compatible with the imposed constraints. The integral range provides a global and threshold-free measure of dependence (Vanmarcke, 1983; Lantuéjoul, 2002; Hristopulos and Zukovic, 2011); in dimension $d=1$, it reduces to the area under the correlation function. The proposed selection principle is intrinsic to the model class and does not depend on sampling design or loss functions.
The resulting hypergeometric ($\mathcal{H}$) model constitutes the main contribution of this paper. It provides a canonical and parsimonious representative of the $\mathcal{GH}$ class, preserving compact support, continuous smoothness parameterization, and modeling flexibility. Notably, the $\mathcal{H}$ model satisfies simple and explicit necessary and sufficient validity conditions, reducing to a single lower bound on the shape parameter that is independent of the spatial dimension. In addition, it admits a beta-type scale-mixture representation and may be interpreted as a generalization of the Euclid’s hat model (Gneiting, 1999), encompassing the classical spherical covariance functions widely used in geostatistics (Chilès and Delfiner, 2012).
We further show that a compact-support reparameterization of the $\mathcal{H}$ model yields the Matérn model as a special limit case, thereby extending to the hypergeometric framework the Wendland-to-Matérn connection proved in Bevilacqua et al. (2022).
A simulation study investigates finite-sample behavior under increasing- and fixed-domain scenarios, and a real-data analysis of climate data illustrates that the proposed $\mathcal{H}$ model can outperform the $\mathcal{MT}$ model in terms of model fit, prediction accuracy, and computational efficiency for both Gaussian and Tukey-$h$ RFs (Xu and Genton, 2017).
The $\mathcal{H}$ model is implemented in the GeoModels package (Bevilacqua et al., 2026). The remainder of the paper is organized as follows. Section 2 reviews the $\mathcal{MT}$ and $\mathcal{GW}$ models. Section 3 introduces the $\mathcal{GH}$ model and discusses identifiability. Section 4 provides validity conditions and derives the $\mathcal{H}$ subclass. Section 5 presents closed-form expressions and compact-support reparameterizations. Section 6 establishes asymptotic properties of the microergodic estimator. Sections 7 and 8 report simulation and real-data results. All proofs are collected in Appendix A.
2 Background: the Matérn and Generalized Wendland correlation models
We denote $\{Z(\boldsymbol{s}):\boldsymbol{s}\in D\}$ a zero-mean Gaussian RF on a bounded set $D$ of $\mathbb{R}^{d}$, with stationary isotropic correlation function $C:\mathbb{R}^{d}\to\mathbb{R}$. We consider the class $\Phi_{d}$ of continuous mappings $\phi:[0,\infty)\to\mathbb{R}$ with $\sigma^{2}:=\phi(0)=1$, such that
${\rm corr}\left(Z(\boldsymbol{s}),Z(\boldsymbol{s}^{\prime})\right)=C(\boldsymbol{h})=\phi(\|\boldsymbol{h}\|),$
with $\boldsymbol{h}=\boldsymbol{s}^{\prime}-\boldsymbol{s}$ and $\boldsymbol{s},\boldsymbol{s}^{\prime}\in D$, and $\|\cdot\|$ denoting the Euclidean norm. Gaussian RFs with such covariance functions are called weakly stationary and isotropic.
Schoenberg (1938) characterized the class $\Phi_{d}$ as scale mixtures of the characteristic functions of random vectors uniformly distributed on the spherical shell of $\mathbb{R}^{d}$, with any finite and nonnegative measure, $F$:
$\phi(x)=\int_{0}^{\infty}\Omega_{d}(xr)F(\text{d}r),\qquad x\geq 0,$
with $\Omega_{d}(x)=\Gamma(d/2)(2/x)^{d/2-1}J_{d/2-1}(x)$ and $J_{\nu}$ the Bessel function of the first kind of order $\nu$. The class $\Phi_{d}$ is nested, with the inclusion relation $\Phi_{1}\supset\Phi_{2}\supset\ldots\supset\Phi_{\infty}$ being strict, and where $\Phi_{\infty}:=\bigcap_{d\geq 1}\Phi_{d}$ is the class of continuous mappings $\phi$, the radial version of which is positive semidefinite on any $d$-dimensional Euclidean space.
Fourier transforms of radial versions of members of $\Phi_{d}$, for a given $d$, have a simple expression, as reported in Yaglom (1987) or Stein (1999). For a member $\phi$ of the class $\Phi_{d}$ such that $\phi(\|\cdot\|)$ is absolutely integrable in $\mathbb{R}^{d}$, we define its isotropic spectral density as
$\widehat{\phi}(z)=\frac{z^{1-d/2}}{(2\pi)^{\frac{d}{2}}}\int_{0}^{\infty}u^{d/2}J_{d/2-1}(uz)\phi(u){\rm d}u,\qquad z\geq 0.$ (1)
In the following we review two important examples:
  1. 1.
    The Matérn ($\cal MT$}) family
    The Matérn model is a two-parameter globally supported correlation function defined as (Matérn, 1986):
    ${\cal MT}_{\nu,\alpha}(x)=\frac{2^{1-\nu}}{\Gamma(\nu)}\left(\frac{x}{\alpha}\right)^{\nu}{\cal K}_{\nu}\left(\frac{x}{\alpha}\right),\qquad x>0,$ (2)
    where ${\cal K}_{\nu}$ is the modified Bessel function of the second kind of order $\nu$ (Abramowitz and Stegun, 1970, 9.6.22). The associated spectral density, $\widehat{{\cal MT}}_{\alpha,\nu}$, is given by
    $\widehat{{\cal MT}}_{\nu,\alpha}(z)=\frac{\Gamma(\nu+d/2)}{\pi^{d/2}\Gamma(\nu)}\frac{\alpha^{d}}{(1+\alpha^{2}z^{2})^{\nu+d/2}},\qquad z\geq 0.$ (3)
    Arguments in Stein (1999) show that ${\cal MT}_{\nu,\alpha}$ belongs to $\Phi_{\infty}$ when both the smoothness parameter $\nu$ and the scale parameter $\alpha$ are strictly positive.
    A key feature of the Matérn model is the smoothness parameter $\nu$, which determines the mean-square differentiability of the sample paths of the associated Gaussian RF. In particular, for any integer $l=0,1,\ldots$, the sample paths are $l$-times mean-square differentiable (in any direction) if and only if $\nu>l$. In the special case $\nu=l+\tfrac{1}{2}$, the correlation function reduces to an exponential term multiplied by a polynomial of degree $l$.
  2. 2.
    The Generalized Wendland ($\cal GW$}) family
    The $\cal GW$} model is a three-parameters compactly supported correlation function defined as (Gneiting, 2002; Bevilacqua et al., 2019):
    $\begin{split}{\cal GW}_{\kappa,\mu,\beta}(x)&=\frac{\Gamma(\kappa)\Gamma(2\kappa+\mu+1)}{\Gamma(2\kappa)\Gamma(\kappa+\mu+1)2^{\mu+1}}\left(1-\frac{x^{2}}{\beta^{2}}\right)^{\kappa+\mu}_{+}\\ &\times{}_{2}F_{1}\left(\frac{\mu}{2},\frac{\mu+1}{2};\kappa+\mu+1;\left(1-\frac{x^{2}}{\beta^{2}}\right)_{+}\right),\quad x\geq 0,\end{split}$ (4)
    with $\beta>0$ a compact support parameter, $\kappa>-\frac{1}{2}$ a smoothness parameter and $\mu$ a shape parameter, and where $(\cdot)_{+}$ stands for the positive part function and $_{2}F_{1}$ for the Gauss hypergeometric function (Abramowitz and Stegun, 1970, 15.1.1). The associated spectral density is given by:
    $\widehat{{\cal GW}}_{\kappa,\mu,\beta}(z)=L\beta^{d}\mathstrut_{1}F_{2}\Big(\lambda;\lambda+\frac{\mu}{2},\lambda+\frac{\mu}{2}+\frac{1}{2};-\frac{(z\beta)^{2}}{4}\Big),\quad z>0,$ (5)
    where $_{1}F_{2}$ is a generalized hypergeometric function, $\lambda=(d+1)/2+\kappa$ and $L=\frac{2^{-d}\pi^{-\frac{d}{2}}\Gamma(\mu+2\nu+1)\Gamma(2\nu+d)\Gamma(\nu)}{\Gamma\left(\nu+{d}/{2}\right)\Gamma(\mu+2\nu+d+1)\Gamma(2\nu)}$ is a normalization constant.
    For a given $\kappa>-\frac{1}{2}$, the validity conditions for $\mathcal{GW}_{\kappa,\mu,\beta}\in\Phi_{d}$ follow as a special case of Theorem 2 by setting $l=\frac{1}{2}$:
    • (A)
      $\mathcal{GW}_{\kappa,\mu,\beta}\in\Phi_{d}$ if and only if $\mu\geq\kappa+\frac{d+1}{2}$, whenever $\kappa\geq 0$ or $d\geq 2$;
    • (B)
      $\mathcal{GW}_{\kappa,\mu,\beta}\in\Phi_{1}$ if $\mu\geq\frac{\sqrt{8\kappa+9}-1}{2}$, and only if $\mu\geq\kappa+1$, for $\frac{1}{2}<\kappa<0$.
    Since $\kappa+1<\frac{\sqrt{8\kappa+9}-1}{2}$ for all $\frac{1}{2}<\kappa<0$, a gap exists between the necessary and sufficient conditions. This partially corrects Bevilacqua et al. (2024), which claimed in case (B) that $\mu\geq\frac{\sqrt{8\kappa+9}-1}{2}$ is both necessary and sufficient for $d=1$. Whether this lower bound is also necessary remains an open problem.
    The ${\cal GW}_{\kappa,\mu,\beta}$ model, as in the ${\cal MT}_{\nu,\alpha}$ case, allows parameterizing in a continuous fashion the mean squared differentiability of the underlying Gaussian RF and its associated sample paths through the smoothness parameter $\kappa$. Specifically, for any integer $l=1,2\ldots$, the sample paths of the ${\cal GW}_{\kappa,\mu,\beta}$ model are $l$ times differentiable, in any direction, if and only if $\kappa>l-0.5$ and for $0\leq\kappa\leq 0.5$ they are not differentiable.
    When $\kappa=k\in\mathbb{N}$, the Generalized Wendland correlation simplifies into the product of an Askey (truncated power) correlation with a polynomial of degree $k$.
    Recently, Bevilacqua et al. (2022) considered a compact support reparameterized version of the $\cal GW$} model, and proved that
    $\lim_{\mu\to\infty}{\cal GW}_{\kappa,\mu,g(\beta,\mu,\kappa)}(x)={\cal MT}_{\kappa+1/2,\beta}(x),\quad\kappa\geq-0.5,$ (6)
    where $g(\beta,\mu,\kappa)=\beta\left(\frac{\Gamma(\mu+2\kappa+1)}{\Gamma(\mu)}\right)^{\frac{1}{1+2\kappa}}$, with uniform convergence on $0,\infty)$.
    As a consequence, the reparameterized $\cal GW$} model can be viewed as a generalization of the $\cal MT$} model with an additional parameter that, for given smoothness and spatial dependence parameters, allows switching from the world of flexible compactly supported covariance functions to the world of flexible globally supported covariance functions.
3 The Gauss hypergeometric model and the identifiability problem
The $\mathcal{GH}$ class in Emery and Alegría (2022) has been defined starting from the following spectral density:
$\widehat{{\cal GH}}_{\delta,\beta,\gamma,a}(z)=\frac{\Gamma(\delta)\Gamma(\beta-\frac{d}{2})\Gamma(\gamma-\frac{d}{2})a^{d}}{2^{d}\pi^{\frac{d}{2}}\Gamma\left(\delta-\frac{d}{2}\right)\Gamma(\beta)\Gamma(\gamma)}\mathstrut_{1}F_{2}\Big(\delta;\beta,\gamma;-\frac{(za)^{2}}{4}\Big),\quad z\geq 0.$ (7)
Note that (7) has basically the same form of the spectral density of the $\cal GW$} model in (5). However, in (7) there is no dependence between the parameters of the $\mathstrut{}_{1}F_{2}$ function as in the $\cal GW$} case.
Using some results in Cho and Yun (2018), Emery and Alegría (2022) show that, for a given $a>0$, the following conditions:
  1. (a)
    $\delta>\frac{d}{2}$;
  2. (b)
    $2(\beta-\delta)(\gamma-\delta)\geq\delta$;
  3. (c)
    $\beta+\gamma\geq 3\delta+\frac{1}{2}$;
are sufficient for the positivity of the spectral density on $0,\infty)$. The associated four-parameter isotropic correlation model is given by:
$\begin{split}\mathcal{GH}_{\delta,\beta,\gamma,a}(x)&=\frac{\Gamma(\beta-\frac{d}{2})\Gamma(\gamma-\frac{d}{2})}{\Gamma(\beta-\delta+\gamma-\frac{d}{2})\Gamma(\delta-\frac{d}{2})}\left(1-\frac{x^{2}}{a^{2}}\right)_{+}^{\beta-\delta+\gamma-\frac{d}{2}-1}\\ &\times{}_{2}F_{1}\left(\beta-\delta;\gamma-\delta;\beta-\delta+\gamma-\frac{d}{2};\left(1-\frac{x^{2}}{a^{2}}\right)_{+}\right),\quad x\geq 0.\end{split}$ (8)
Here $a>0$ is a compact support parameter, $\delta>\frac{d}{2}$ is a smoothness parameter, and $\beta$ and $\gamma$ are two additional shape parameters.
As in the $\cal MT$} and $\cal GW$} cases, the $\cal G$H} model allows parameterizing in a continuous fashion the mean squared differentiability of the underlying Gaussian RF and its associated sample paths through the smoothness parameter $\delta$.
Furthermore, it is not difficult to see that the $\cal G$H} model includes the $\cal GW$} and the $\cal MT$} models as special cases. For instance,
$\mathcal{GH}_{\delta,\delta+\frac{\mu}{2},\delta+\frac{\mu}{2}+\frac{1}{2},a}(x)={\cal GW}_{\kappa,\mu,a}(x),$ (9)
where $\delta=\frac{d+1}{2}+\kappa$.
Although the $\cal G$H} family is highly flexible, it is not identifiable in the usual parametric sense. Indeed, even under increasing-domain asymptotics—where one would in principle expect all parameters to be estimable—the parameter-to-correlation map is not injective. In particular,
$\mathcal{GH}_{\delta,\beta,\gamma,a}=\mathcal{GH}_{\delta,\gamma,\beta,a},$
because the Gauss hypergeometric function is symmetric in its first two arguments:
$_{2}F_{1}(a,b;\cdot\,;x)={}_{2}F_{1}(b,a;\cdot\,;x).$
Hence distinct parameter values (obtained by swapping $\beta$ and $\gamma$) induce exactly the same correlation function. This implies that additional constraints are required.
A related inferential limitation arises under fixed-domain (infill) asymptotics, where distinct parameter values may induce equivalent Gaussian measures. The following theorem establishes the equivalence condition for two distinct Gaussian measures with $\mathcal{GH}$ correlation models that share a common smoothness parameter. Hereafter, we denote by $P(\rho_{i})$, $i=0,1$, a zero-mean Gaussian measure with an isotropic covariance function $\rho_{i}$.
Theorem 1.
Let $P(\sigma_{i}^{2}{\cal GH}_{\delta,\beta_{i},\gamma_{i},a_{i}})$, $i=0,1$, be two zero-mean Gaussian measures and let $\delta>\frac{d}{2}$. If $\frac{d+1}{2}+3\delta<\min(\beta_{1}+\gamma_{1},\beta_{0}+\gamma_{0}),$ for any bounded infinite set $D\subset\mathbb{R}^{d}$, $d=1,2,3$, the Gaussian measures $P(\sigma_{1}^{2}{\cal GH}_{\delta,\beta_{1},\gamma_{1},a_{1}})$ and $P(\sigma_{0}^{2}{\cal GH}_{\delta,\beta_{0},\gamma_{0},a_{0}})$ are equivalent on the paths of $Z(\boldsymbol{s})$, $\boldsymbol{s}\in D$, if and only if
$\frac{\sigma_{0}^{2}L(\beta_{0},\gamma_{0})}{a_{0}^{2\delta-d}}=\frac{\sigma_{1}^{2}L(\beta_{1},\gamma_{1})}{a_{1}^{2\delta-d}},$ (10)
where $L(\beta_{i},\gamma_{i})=\dfrac{2^{2\delta-d}\Gamma(\beta_{i}-d/2)\Gamma(\gamma_{i}-d/2)}{\Gamma(\gamma_{i}-\delta)\Gamma(\beta_{i}-\delta)}$, $i=0,1$.
The restriction to $d=1,2,3$ is typical when studying equivalence of Gaussian measures for isotropic models on $\mathbb{R}^{d}$, and is not specific to the $\cal G$H} class. Indeed, the same dimensional restriction appears in the equivalence results for the $\cal MT$} model (Zhang, 2004), and stems from the integrability conditions required by the criteria of Stein (2004).
An immediate consequence of Theorem 1 is that, for fixed $\delta$, the microergodic parameter is given by
$\frac{\sigma^{2}L(\beta,\gamma)}{a^{2\delta-d}}.$
Because $L(\beta,\gamma)=L(\gamma,\beta)$, this quantity does not distinguish $(\beta,\gamma)$ from $(\gamma,\beta)$. Thus, while the microergodic parameter is the relevant identifiable feature under fixed-domain asymptotics, the individual shape parameters $\beta$ and $\gamma$ are not identifiable without an additional constraint.
Summarizing, the $\mathcal{GH}$ correlation model is highly flexible and includes the $\mathcal{GW}$ and $\mathcal{MT}$ models as special cases. However, it is not identifiable in the usual parametric sense because swapping $(\beta,\gamma)$ leaves the correlation unchanged. This issue can be observed under both increasing and fixed domain scenario. To this end, in the next Section we assume that the second shape parameter ($\gamma$) is a function of the first ($\beta$).
4 Fixing the identifiability problem: the Hypergeometric correlation model
In this section we first refine the validity conditions of the $\mathcal{GH}$ model, clarifying the relationship between necessary and sufficient constraints. We then address the identifiability issue discussed in the previous section by introducing a parsimonious subclass selected according to an explicit optimality criterion.
To this end, we adopt a $\cal GW$}-type reparameterization of the $\mathcal{GH}_{\delta,\beta,\gamma,a}$ model and set
$\delta=\frac{d+1}{2}+\kappa,\qquad\beta=\delta+\frac{\mu}{2},\qquad\gamma=\beta+l.$ (11)
As noted in (9), this parametrization includes the $\cal GW$} model as the special case $l=\frac{1}{2}$.
The following theorem improves Theorem 1 in Emery and Alegría (2022) by providing sharp validity conditions for $\mathcal{GH}_{\delta,\delta+\frac{\mu}{2},\delta+\frac{\mu}{2}+l,a}$.
Theorem 2.
Let $d$ be fixed, $\kappa>-\frac{1}{2}$, $\delta=\frac{d+1}{2}+\kappa$, $a>0$, and $l\geq 0$. Then:
  1. 1.
    If $0\leq l\leq\frac{d}{2}+\kappa$, then
    $\mathcal{GH}_{\delta,\delta+\frac{\mu}{2},\delta+\frac{\mu}{2}+l,a}\in\Phi_{d}\quad\text{if and only if}\quad\mu\geq\delta-l+\frac{1}{2}.$
  2. 2.
    If $l>\frac{d}{2}+\kappa$, then
    $\mathcal{GH}_{\delta,\delta+\frac{\mu}{2},\delta+\frac{\mu}{2}+l,a}\in\Phi_{d}\quad\text{if}\quad\mu\geq\sqrt{2\kappa+l^{2}+d+1}-l,$
    and only if
    $\mu\geq\delta-l+\frac{1}{2}.$
For $l=\frac{1}{2}$, the necessary and sufficient validity conditions for $\mathcal{GW}_{\kappa,\mu,a}\in\Phi_{d}$ discussed in Section 2 are recovered as a special case. Specifically, since $l=\frac{1}{2}\leq\frac{d}{2}+\kappa$ if and only if $\kappa\geq\frac{1}{2}-\frac{d}{2}$, which holds for all $d\geq 2$ and $\kappa>-\frac{1}{2}$, or for $d=1$ and $\kappa\geq 0$, Regime 1 of Theorem 2 gives the necessary and sufficient condition
$\mu\geq\frac{d+1}{2}+\kappa,$
For $d=1$ and $\frac{1}{2}<\kappa<0$, instead $l=\frac{1}{2}>\frac{d}{2}+\kappa$, so Regime 2 applies, yielding the sufficient condition $\mu\geq 0.5(\sqrt{8\kappa+9}-1)$ and the necessary condition $\mu\geq\kappa+1$, consistently with the discussion in Section 2.
As discussed in Section 3, the model $\mathcal{GH}_{\delta,\beta,\gamma,a}$ is invariant under the interchange $\beta\leftrightarrow\gamma$, corresponding to $l\mapsto-l$ in the model $\mathcal{GH}_{\delta,\delta+\frac{\mu}{2},\delta+\frac{\mu}{2}+l,a}$. The admissible region is therefore symmetric with respect to $l=0$, and restricting attention to $l\geq 0$ entails no loss of generality.
Translating Theorem 2 back to the original parameterization $\mathcal{GH}_{\delta,\beta,\gamma,a}$, the following equivalent conditions hold for $\delta>d/2$ and $a>0$:
  1. 1.
    If $0\leq\gamma-\beta\leq\delta-\frac{1}{2}$, then
    $\mathcal{GH}_{\delta,\beta,\gamma,a}\in\Phi_{d}\quad\text{if and only if}\quad\beta+\gamma\geq 3\delta+\frac{1}{2}.$
  2. 2.
    If $\gamma-\beta>\delta-\frac{1}{2}$, then
    $\mathcal{GH}_{\delta,\beta,\gamma,a}\in\Phi_{d}\quad\text{if}\quad\beta+\gamma\geq 2\delta+\sqrt{(\gamma-\beta)^{2}+2\delta},$
    and only if
    $\beta+\gamma\geq 3\delta+\frac{1}{2}.$
In particular, in Regime 1 ($0\leq\gamma-\beta\leq\delta-\frac{1}{2}$), condition (c) of Section 3 implies condition (b), so that (a) and (c) alone characterize validity. Hence, within this regime, the sufficient conditions of Emery and Alegría (2022) reduce to (a) and (c) and are therefore also necessary. In Regime 2 ($\gamma-\beta>\delta-\frac{1}{2}$), whether the gap between sufficiency and necessity can be closed remains an open question.
We now turn to the structural identifiability issue for the $\mathcal{GH}_{\delta,\delta+\frac{\mu}{2},\delta+\frac{\mu}{2}+l,a}$ model. To resolve it, the parameter $l$ must either be fixed (as in the $\cal GW$} model) or linked to the remaining parameters and/or to the spatial dimension. This naturally raises the question: which choice of $l\geq 0$ is optimal ?
We address this question by introducing an explicit optimality criterion that selects a single parsimonious representative once smoothness and compact support are fixed. We adopt the criterion of maximum integral range attainable by a compactly supported correlation with given smoothness and support.
For an isotropic correlation $\phi\in\Phi_{d}$, the integral range is
$A=\int_{\mathbb{R}^{d}}\phi(\|\boldsymbol{h}\|)\,\mathrm{d}\boldsymbol{h}=\frac{\pi^{\frac{d}{2}}}{2^{d-1}\Gamma(\frac{d}{2})}\int_{0}^{\infty}x^{d-1}\phi(x)\,\mathrm{d}x.$ (12)
This quantity, originally introduced by Lantuéjoul (1991), provides a global and threshold-free measure of dependence (Yaglom, 1987; Matheron, 1989; Hristopulos and Zukovic, 2011). If $\phi(\|\cdot\|)$ is absolutely integrable in $\mathbb{R}^{d}$ (in particular, if it is compactly supported), then $A=\widehat{\phi}(0)$, so that the integral range coincides with the spectral density at zero frequency. When $d=1$, AA reduces to the area under the correlation function.
For compactly supported correlations, maximizing $A$ therefore selects, among all admissible models with fixed smoothness and support, the one with maximal low-frequency spectral power, that is, the strongest large-scale dependence compatible with the imposed constraints. While alternative optimality principles could be considered, they would typically depend on the sampling design or on a specific loss function. In contrast, the integral range criterion is intrinsic to the model class and yields a reproducible, one-to-one specification within the family.
For a parametric model $\phi_{\boldsymbol{\theta}}\in\Phi_{d}$, we denote the associated integral range by $A_{\boldsymbol{\theta},d}$. For example, for the Matérn correlation one has
$A_{\nu,\alpha,d}=\frac{\alpha^{d}\pi^{d/2}\Gamma\!\left(\nu+\frac{d}{2}\right)}{\Gamma(\nu)}.$ (13)
The following theorem provides the integral range of $\mathcal{GH}_{\delta,\delta+\frac{\mu}{2},\delta+\frac{\mu}{2}+l,a}$.
Theorem 3.
The integral range of $\mathcal{GH}_{\delta,\delta+\frac{\mu}{2},\delta+\frac{\mu}{2}+l,a}$ is given by:
$A_{\kappa,\mu,l,a,d}=a^{d}\left[\frac{\pi^{d/2}\Gamma(\frac{d+1}{2}+\kappa)\Gamma\left(\frac{\mu+1}{2}+\kappa\right)\Gamma\left(\frac{\mu+1}{2}+\kappa+l\right)}{2^{d}\Gamma(\kappa+\frac{1}{2})\Gamma\left(\frac{\mu+1+d}{2}+\kappa\right)\Gamma\left(\frac{\mu+1+d}{2}+\kappa+l\right)}\right].$ (14)
Note that, as expected, the compact support $a$ raised to power $d$ represents a scale factor in the integral range computation. In addition, using asymptotic properties of the Gamma function, it can be shown that $A_{\kappa,\mu,l,a,d}\sim\mu^{-d}$ as $\mu\to\infty$ and, as a consequence,
${\lim_{\mu\to\infty}}A_{\kappa,\mu,l,a,d}=0.$ (15)
This implies that, as the shape parameter $\mu$ increases, the integral range approaches zero, regardless of the compact support and smoothness parameters. The following theorem establishes that, for a given $l\geq 0$, the maximum integral range is obtained by fixing $\mu$ on the lower bound of its parametric space.
Theorem 4.
Let $\kappa>0$, $a>0$, and $d\in\mathbb{N}$ be fixed. For each $l\geq 0$, the maximum of $A_{\kappa,\mu,l,a,d}$ with respect to $\mu$ (over its admissible parameter space) is attained at $\mu^{\star}(l)$, where:
  1. 1.
    if $0\leq l\leq\dfrac{d}{2}+\kappa$, then
    $\mu^{\star}(l)=\mu_{1}(l):=\delta-l+\frac{1}{2};$
  2. 2.
    if $l>\dfrac{d}{2}+\kappa$, then
    $\mu^{\star}(l)=\mu_{2}(l):=\sqrt{2\kappa+l^{2}+d+1}-l.$
For instance, let $\kappa\geq 0$. If $l=\frac{1}{2}$, then from (9) $\mathcal{GH}_{\delta,\delta+\frac{\mu}{2},\delta+\frac{\mu+1}{2},a}={\cal GW}_{\kappa,\mu,a}$ and the maximum integral range is attained when $\mu=\mu_{1}(\frac{1}{2})=\delta=\frac{d+1}{2}+\kappa$, which is the lower bound of $\cal GW$} model in this case.
As an example, Figure 1 (left part) depicts the ${\cal GW}_{0,\mu,1}$ model and the right part depicts the ${\cal GW}_{1,\mu,1}$ model for $\mu=\mu_{1}(\frac{1}{2}),3,10,100$. In these examples $d=1$ and it is apparent that the model with maximum area corresponds to the case $\mu=\mu_{1}(\frac{1}{2})$ (red line) in both examples.
The next step is to find $l\geq 0$ that maximizes $A_{\kappa,\mu_{i}(l),l,a,d}$ for $i=1,2$.
Theorem 5.
Let $\kappa>-1/2$, $a>0$, and $d\in\mathbb{N}$ be fixed, and set $\delta=\frac{d}{2}+\kappa$.
  1. 1.
    The function $l\mapsto A_{\kappa,\mu_{1}(l),l,a,d}$ is non-decreasing on $0,\delta]$, hence
    $\max_{0\leq l\leq\delta}A_{\kappa,\mu_{1}(l),l,a,d}=A_{\kappa,\mu_{1}(\delta),\delta,a,d}=A_{\kappa,1,\delta,a,d}.$
  2. 2.
    The function $l\mapsto A_{\kappa,\mu_{2}(l),l,a,d}$ is non-increasing on $(\delta,\infty)$ and
    $\lim_{l\downarrow\delta}A_{\kappa,\mu_{2}(l),l,a,d}=A_{\kappa,1,\delta,a,d}.$
In particular, the global maximum over all admissible $(\mu,l)$ is attained at $(\mu,l)=(1,\delta)$ and is equal to $A_{\kappa,1,\delta,a,d}$.
Refer to caption Refer to caption
Figure 1: Left part: ${\cal GW}_{0,\mu,1}$ for $\mu=\mu_{1}(\frac{1}{2}),3,10,100$ with $\delta=\frac{d+1}{2}+\kappa$ and $\kappa=0$. Right part: the same Figure but with $\kappa=1$. In both cases $d=1$.
As an example, Figure 2 compares $\mathcal{GH}_{\delta,\delta+\frac{\mu_{1}(l)}{2},\delta+\frac{\mu_{1}(l)}{2}+l,1}$ when $l=0,\frac{1}{2},\frac{d}{2}+\kappa$ (black, red and blue lines, respectively). In particular the left part depicts the case $\kappa=0$ and the right part depicts the case $\kappa=1$ ($d=2$ in both cases). It is apparent that the case $l=\frac{d}{2}+\kappa$ (blue line) corresponds to the model with higher integral range and improves the maximum integral range attained by the $\cal GW$} models (red line).
Refer to caption Refer to caption
Figure 2: Left part: $\mathcal{GH}_{\delta,\delta+\frac{\mu(l)}{2},\delta+\frac{\mu(l)}{2}+l,1}$ with $\delta=\frac{d+1}{2}+\kappa$, $\kappa=0$, $\mu(l)=\delta-l+\frac{1}{2}$ and $l=0,\frac{1}{2},\frac{d}{2}+\kappa$ (black, red and blue lines respectively). Right part: the same Figure but with $\kappa=1$. The case $l=\frac{1}{2}$ (red line) corresponds to the $\mathcal{GW}_{\kappa,\mu,1}$ model. The blue line corresponds to the correlation model that maximizes the integral range. In both cases $d=2$.
Summarizing our results, the model that can potentially achieve the maximum integral range is $\mathcal{GH}_{\delta,\delta+\frac{\mu}{2},\delta+\frac{\mu}{2}+l,a}$ with $l=\frac{d}{2}+\kappa$ and, in particular, the maximum is achieved when $\mu=\mu_{1}(d/2+\kappa)=1$. Taking into account (15), this implies that
$0\leq A_{\kappa,\mu,l,a,d}\leq a^{d}\left[\frac{\pi^{d/2}\,\Gamma\left(\frac{d+1}{2}+\kappa\right)\,\Gamma\left(1+\kappa\right)\,\Gamma\left(1+\frac{d}{2}+2\kappa\right)}{2^{d}\,\Gamma\left(\kappa+\frac{1}{2}\right)\,\Gamma\left(1+\frac{d}{2}+\kappa\right)\,\Gamma\left(1+d+2\kappa\right)}\right].$
Accounting for (11), the model with maximum integral range (obtained when $l=\frac{d}{2}+\kappa$ and $\mu=1$), is $\mathcal{GH}_{\delta,\delta+\frac{1}{2},2\delta,a}$ and coincides with the so-called Euclid’s hat correlation model (Gneiting, 1999) when $\kappa=0$ (triangular or tent model if $d=1$, circular model if $d=2$, spherical model if $d=3$, pentaspherical model if $d=5$, see Section 5) and with an upgraded Euclid’s hat when $\kappa\neq 0$ (Emery and Alegría, 2022), obtained from the Euclid’s hat by a (fractional) montée or upgrading, as defined by Matheron (1965, I.4.18).
Fixing $l=\frac{d}{2}+\kappa$, we define the hypergeometric correlation by
$\mathcal{H}_{\kappa,\mu,a,d}(x):=\mathcal{GH}_{\delta,\delta+\frac{\mu}{2},\delta+\frac{\mu+d}{2}+\kappa,a}(x),\qquad x\geq 0.$ (16)
An explicit expression is given by
$\begin{split}\mathcal{H}_{\kappa,\mu,a,d}(x)&=\frac{\Gamma\!\left(\kappa+\frac{\mu+1}{2}\right)\Gamma\!\left(2\kappa+\frac{d+\mu+1}{2}\right)}{\Gamma\!\left(\mu+\frac{d+1}{2}+2\kappa\right)\Gamma\!\left(\kappa+\frac{1}{2}\right)}\left(1-\frac{x^{2}}{a^{2}}\right)_{+}^{\mu+\frac{d-1}{2}+2\kappa}\\[4.0pt] &\quad\times{}_{2}F_{1}\!\left(\frac{\mu}{2},\frac{\mu+d}{2}+\kappa;\mu+\frac{d+1}{2}+2\kappa;\left(1-\frac{x^{2}}{a^{2}}\right)_{+}\right),\qquad x\geq 0.\end{split}$ (17)
Notably, the model $\mathcal{H}_{\kappa,\mu,a,d}$ always lies in Regime 1 of Theorem 2, where validity is completely characterized. Indeed, by setting $l=\frac{d}{2}+\kappa$ in Theorem 2, we obtain the condition
$\mathcal{H}_{\kappa,\mu,a,d}\in\Phi_{d}\quad\text{if and only if}\quad\mu\geq 1.$
These validity conditions are substantially simpler than those for the $\mathcal{GW}_{\kappa,\mu,a}$ model discussed in Section 2. In particular, the lower bound on $\mu$ is independent of the spatial dimension. The trade-off is that, unlike the $\mathcal{GW}_{\kappa,\mu,a}$ model, the expression of $\mathcal{H}_{\kappa,\mu,a,d}$ explicitly depends on $d$.
The next theorem shows that the subclass ${\cal H}_{\kappa,1,a,d}$, obtained by maximizing the integral range, retains substantial modeling flexibility. Through beta-type scale mixtures, it can generate any member of ${\cal H}_{\kappa,\mu,a,d}$ and a broad subclass of the Gauss–hypergeometric family $\mathcal{GH}_{\delta,\delta+\frac{\mu}{2},\delta+\frac{\mu}{2}+l,a}$. Thus, the maximal integral range criterion does not restrict modeling capacity, but instead selects a parsimonious representative that remains rich within the admissible class.
Theorem 6.
Let $a>0$, $\kappa>-\frac{1}{2}$, $\delta=\frac{d+1}{2}+\kappa$, $\mu>1$, and $l\geq 0$ such that $\mu+2l>1+d+2\kappa$, so that $\mathcal{GH}_{\delta,\delta+\frac{\mu}{2},\delta+\frac{\mu}{2}+l,a}\in\Phi_{d}$ by Theorem 2. Then $\mathcal{GH}_{\delta,\delta+\frac{\mu}{2},\delta+\frac{\mu}{2}+l,a}$ admits a beta-type scale mixture representation in terms of hypergeometric correlations $\mathcal{H}_{\kappa,1,b,d}$ with $0<b\leq a$:
$\begin{split}\mathcal{GH}_{\delta,\delta+\frac{\mu}{2},\delta+\frac{\mu}{2}+l,a}(x)&=\frac{\Gamma\!\left(\delta+\frac{\mu-d}{2}\right)\Gamma\!\left(\delta+\frac{\mu-d}{2}+l\right)}{\Gamma\!\left(\delta+\frac{1-d}{2}\right)\Gamma\!\left(\delta+\frac{1}{2}+\kappa\right)\Gamma\!\left(\frac{\mu-1}{2}\right)\Gamma\!\left(\frac{\mu-1-d}{2}-\kappa+l\right)}\\ &\quad\times\int_{0}^{1}\!\!\int_{0}^{1}u^{\delta-\frac{d+1}{2}}(1-u)^{\frac{\mu-3}{2}}v^{\delta-\frac{1}{2}+\kappa}(1-v)^{\frac{\mu-d-3}{2}-\kappa+l}\\ &\qquad\qquad\times\mathcal{H}_{\kappa,1,a\sqrt{uv},d}(x)\,\mathrm{d}u\,\mathrm{d}v,\qquad x\geq 0.\end{split}$ (18)
In particular, when $l=\frac{d}{2}+\kappa$, one obtains
$\begin{split}\mathcal{H}_{\kappa,\mu,a,d}(x)&=\frac{\Gamma\!\left(\delta+\frac{\mu-d}{2}\right)\Gamma\!\left(\delta+\frac{\mu}{2}+\kappa\right)}{\Gamma\!\left(\delta+\frac{1-d}{2}\right)\Gamma\!\left(\delta+\frac{1}{2}+\kappa\right)\Gamma^{2}\!\left(\frac{\mu-1}{2}\right)}\\ &\quad\times\int_{0}^{1}\!\!\int_{0}^{1}u^{\delta-\frac{d+1}{2}}(1-u)^{\frac{\mu-3}{2}}v^{\delta-\frac{1}{2}+\kappa}(1-v)^{\frac{\mu-3}{2}}\\ &\qquad\qquad\times\mathcal{H}_{\kappa,1,a\sqrt{uv},d}(x)\,\mathrm{d}u\,\mathrm{d}v,\qquad x\geq 0.\end{split}$
Clearly, the ${\cal H}_{\kappa,\mu,a,d}$ model allows parameterizing in a continuous fashion the mean squared differentiability of the underlying Gaussian RF and its associated sample paths through the smoothness parameter $\kappa$ in the same way as the ${\cal GW}_{\kappa,\mu,\beta}$ model. Figure 3 depicts two realizations of zero-mean and unit-variance Gaussian RFs with correlations ${\cal H}_{\kappa,4,0.2,2}$ for $\kappa=0,1,2$. The realizations were constructed via the Cholesky decomposition of the covariance matrix.
Refer to caption Refer to caption
Figure 3: Two Gaussian RF realizations with ${\cal H}_{\kappa,4,0.2,2}$ correlation model, for $\kappa=0,1$ (from left to right).
5 Some special cases and bridges to the Matérn model
The computation of the correlation function $\mathcal{H}_{\kappa,\mu,a,d}$ requires the evaluation of the Gauss hypergeometric function. Although efficient numerical implementations are available (Johansson, 2019), this may be less familiar to spatial statisticians and machine learning practitioners accustomed to closed-form parametric correlation models.
As in the $\mathcal{MT}$ and $\mathcal{GW}$ families, substantial simplifications arise in certain special cases. The following theorem considers the case $\kappa=k\in\mathbb{N}_{0}$, $d=1$, and $\mu\geq 1$.
Theorem 7.
If $\kappa=k$ is a nonnegative integer, $d=1$, and $\mu\geq 1$, then
$\begin{split}\mathcal{H}_{k,\mu,a,1}(x)&=\frac{\sqrt{\pi}\,\Gamma\!\left(\frac{\mu}{2}+2k+1\right)}{\Gamma\!\left(k+\tfrac{1}{2}\right)}\left(1-\frac{x}{a}\right)^{\mu+2k}\left(\frac{x}{a}\right)^{k}\\ &\quad\times\sum_{n=0}^{k}\frac{(n+k)!}{2^{2n}n!\,\Gamma\!\left(n+\frac{\mu}{2}+k+1\right)(k-n)!}\left(\frac{(1-\frac{x}{a})^{2}}{\frac{x}{a}}\right)^{n}.\end{split}$
General closed forms for arbitrary $d>1$, $\mu\geq 1$, and integer $\kappa$ are more difficult to obtain. Nevertheless, additional simplifications can be derived using known identities of the Gauss hypergeometric function. For instance, when $d$ is odd, $\mu\geq 1$, and $\kappa=0$, the following closed-form expression holds.
Theorem 8.
Let $d$ be an odd integer and $\mu\geq 1$. If $\kappa=0$, then
$\mathcal{H}_{0,\mu,a,d}(x)=C\left(1-\frac{x}{a}\right)_{+}^{\mu+\frac{d-1}{2}}\sum_{n=0}^{\frac{d-1}{2}}\frac{\left(\frac{1-d}{2}\right)_{n}\left(\frac{d+1}{2}\right)_{n}\left(1-\frac{x}{a}\right)_{+}^{n}}{2^{n}n!\,\Gamma\!\left(n+\frac{d-1}{2}+\mu+1\right)},$
where
$C=\frac{2^{\mu+\frac{d-1}{2}}\Gamma\!\left(\frac{\mu+1}{2}\right)\Gamma\!\left(\frac{\mu+d+1}{2}\right)}{\sqrt{\pi}}.$
Additional closed-form expressions, including spherical models, are reported in Appendix B.
The Matérn model can be recovered as a limiting case of $\mathcal{H}_{\kappa,\mu,a,d}$ under an appropriate compact-support reparameterization. To establish this result, we first characterize the equivalence between the Gaussian measures $P(\sigma_{0}^{2}\mathcal{MT}_{\nu,\alpha})$ and $P(\sigma_{1}^{2}\mathcal{H}_{\kappa,\mu,a,d})$. The proof is provided in Appendix A.
Theorem 9.
Let $P(\sigma_{0}^{2}{\cal MT}_{\nu,\alpha})$ and $P(\sigma_{1}^{2}{\cal H}_{\kappa,\mu,a,d})$ be two zero-mean Gaussian measures, and let $\kappa>-\frac{1}{2}$ and $\nu=\kappa+\frac{1}{2}$. For any bounded infinite set $D\subset\mathbb{R}^{d}$, $d=1,2,3$, the Gaussian measures $P(\sigma_{0}^{2}{\cal MT}_{\nu,\alpha})$ and $P(\sigma_{1}^{2}{\cal H}_{\kappa,\mu,a,d})$ are equivalent on the paths of $Z(\mathbf{s})$, $\mathbf{s}\in D$, if and only if
$\frac{\sigma_{0}^{2}}{\alpha^{2\nu}}=\frac{\sigma_{1}^{2}}{a^{2\kappa+1}}\,B(\kappa,\mu,d),$ (19)
where
$B(\kappa,\mu,d)=\frac{2^{2\kappa+1}\Gamma\!\left(\frac{\mu+1}{2}+\kappa\right)\Gamma\!\left(\frac{\mu+d+1}{2}+2\kappa\right)}{\Gamma\!\left(\frac{\mu}{2}\right)\Gamma\!\left(\frac{\mu+d}{2}+\kappa\right)}.$ (20)
Assuming a common variance and imposing the relation $\nu=\kappa+\frac{1}{2}$, the equivalence condition (19) is equivalently expressed as
$a=\alpha\,B(\kappa,\mu,d)^{\frac{1}{1+2\kappa}}:=b(\alpha,\mu,\kappa,d).$
Using this compact-support reparameterization and following the arguments in Bevilacqua et al. (2022), one obtains
$\lim_{\mu\to\infty}{\cal H}_{\kappa,\mu,b(\alpha,\mu,\kappa,d),d}(x)={\cal MT}_{\kappa+\frac{1}{2},\alpha}(x),\quad\kappa>-\frac{1}{2},$ (21)
with uniform convergence on $0,\infty)$.
[…]
It is worth emphasizing that, under this compact-support reparameterization, the role of the shape parameter $\mu$ changes substantially. For fixed smoothness $\kappa$ and scale parameter $\alpha=1$, $\mu$ determines the compact support.
As $\mu\to\infty$, ${\cal H}_{\kappa,\mu,b(1,\mu,\kappa,d),d}$ converges to the Matérn model ${\cal MT}_{\kappa+\frac{1}{2},1}$, with maximal integral range given by (13). Conversely, as $\mu$ approaches its lower bound, the compact support decreases, potentially yielding more flexible local dependence structures.
A comparison with the corresponding $\mathcal{GW}_{\kappa,\mu,g(1,\mu,\kappa)}$ model (see (6)) shows that, over a substantial portion of the parameter space, the reparameterized $\mathcal{H}$ models can achieve smaller compact supports. In particular, when $d\geq 2$ or $\kappa\geq 0$,
$b(1,1,\kappa,d)\leq g(1,\mu_{\min}(\kappa,d),\kappa),$
where $\mu_{\min}(\kappa,d)=\kappa+\frac{d+1}{2}$. The case $d=1$ and $\frac{1}{2}<\kappa<0$ is excluded here, as the exact lower bound $\mu_{\min}$ for the $\mathcal{GW}$ model remains an open problem in that regime.
6 Asymptotic properties of the maximum likelihood estimator of the Hypergeometric model
We study the asymptotic properties of the maximum likelihood (ML) estimator for the parameters of the model $\sigma^{2}\mathcal{H}_{\kappa,\mu,a,d}$ under both increasing-domain and fixed-domain asymptotics.
Under increasing-domain asymptotics, the parameters $(\sigma^{2},\kappa,\mu,a)$ are consistently estimable, and the ML estimator is asymptotically Gaussian with covariance matrix equal to the inverse Fisher information (Mardia and Marshall, 1984).
Under fixed-domain asymptotics, the situation differs. Using Theorem 1 and the parameterization (11) with $l=\frac{d}{2}+\kappa$, the equivalence condition for two Gaussian measures $P(\sigma_{i}^{2}{\cal H}_{\kappa,\mu_{i},a_{i},d})$, $i=0,1$, sharing the same smoothness parameter, for $d=1,2,3$ and $\min(\mu_{0},\mu_{1})\geq 1+\frac{d}{2}$, is
$\frac{\sigma_{0}^{2}B(\kappa,\mu_{0},d)}{a_{0}^{2\kappa+1}}=\frac{\sigma_{1}^{2}B(\kappa,\mu_{1},d)}{a_{1}^{2\kappa+1}}.$ (22)
Hence only the microergodic parameter
$\frac{\sigma^{2}B(\kappa,\mu,d)}{a^{2\kappa+1}}$
is consistently estimable. If $\mu$ is fixed and known, this reduces to $\sigma^{2}/a^{2\kappa+1}$.
Let $D\subset\mathbb{R}^{d}$ be bounded and $\mathbf{Z}_{n}=(Z(\mathbf{s}_{1}),\ldots,Z(\mathbf{s}_{n}))^{\prime}$ a realization of a zero-mean stationary Gaussian RF with covariance $\sigma^{2}\phi(\|\cdot\|;\boldsymbol{\tau})$, $\phi\in\Phi_{d}$, $\phi(0;\boldsymbol{\tau})=1$. Denote by
$R_{n}(\boldsymbol{\tau})=[\phi(\|\mathbf{s}_{i}-\mathbf{s}_{j}\|;\boldsymbol{\tau})]_{i,j=1}^{n}$
the correlation matrix. The Gaussian log-likelihood is
$\mathcal{L}_{n}(\sigma^{2},\boldsymbol{\tau})=-\frac{1}{2}\Big(n\log(2\pi\sigma^{2})+\log|R_{n}(\boldsymbol{\tau})|+\frac{1}{\sigma^{2}}\mathbf{Z}_{n}^{\prime}R_{n}(\boldsymbol{\tau})^{-1}\mathbf{Z}_{n}\Big).$ (23)
Under the $\mathcal{H}$ model, $\phi(\cdot;\boldsymbol{\tau})=\mathcal{H}_{\kappa,\mu,a,d}$ with $\boldsymbol{\tau}=(\kappa,\mu,a)^{\prime}$. In what follows, $\kappa$ and $\mu$ are assumed fixed and known, so we write $\tau=a$.
Let $\hat{\sigma}_{n}^{2}(a)$ denote the ML estimator of $\sigma^{2}$ for fixed $a>0$:
$\hat{\sigma}_{n}^{2}(a)=\mathbf{Z}_{n}^{\prime}R_{n}(a)^{-1}\mathbf{Z}_{n}/n.$ (24)
The following result establishes strong consistency and asymptotic normality of $\hat{\sigma}_{n}^{2}(a)/a^{1+2\kappa}$. The proof follows the same arguments as Theorem 8 in Bevilacqua et al. (2019).
Theorem 10.
Let $Z(\mathbf{s})$, $\mathbf{s}\in D\subset\mathbb{R}^{d}$, $d=1,2,3$, be a zero-mean Gaussian RF with covariance $\sigma^{2}\mathcal{H}_{\kappa,\mu,a,d}$, $\mu\geq 1+\frac{d}{2}$. Suppose $(\sigma_{0}^{2},a_{0})\in(0,\infty)^{2}$. Then, for fixed $a>0$, as $n\to\infty$,
  1. 1.
    $\hat{\sigma}_{n}^{2}(a)/a^{1+2\kappa}\stackrel{{\scriptstyle a.s.}}{{\longrightarrow}}\sigma_{0}^{2}/a_{0}^{1+2\kappa}$;
  2. 2.
    $\sqrt{n}\big(\hat{\sigma}_{n}^{2}(a)/a^{1+2\kappa}-\sigma_{0}^{2}/a_{0}^{1+2\kappa}\big)\stackrel{{\scriptstyle\mathcal{D}}}{{\longrightarrow}}\mathcal{N}\big(0,2(\sigma_{0}^{2}/a_{0}^{1+2\kappa})^{2}\big)$.
We next consider joint maximization of (23) with respect to $(\sigma^{2},a)\in(0,\infty)\times I$, $I=[a_{L},a_{U}]$, $0<a_{L}<a_{U}<\infty$. Let
$\hat{\sigma}_{n}^{2}(\hat{a}_{n})=\mathbf{Z}_{n}^{\prime}R_{n}(\hat{a}_{n})^{-1}\mathbf{Z}_{n}/n,\quad\hat{a}_{n}=\arg\max_{a\in I}\mathcal{PL}_{n}(a),$
where $\mathcal{PL}_{n}(a)$ is the profile log-likelihood
$\mathcal{PL}_{n}(a)=-\frac{1}{2}\big(\log(2\pi)+n\log\hat{\sigma}_{n}^{2}(a)+\log|R_{n}(a)|+n\big).$ (25)
The following lemma establishes a monotonicity property of $\hat{\sigma}_{n}^{2}(a)/a^{1+2\kappa}$ as a function of $a$. Its proof is given in Appendix A.
Lemma 1.
Let $S_{n}=\{\mathbf{s}_{1},\ldots,\mathbf{s}_{n}\}\subset D$. For any $a_{1}<a_{2}$ in $I$,
$\frac{\hat{\sigma}_{n}^{2}(a_{1})}{a_{1}^{1+2\kappa}}\leq\frac{\hat{\sigma}_{n}^{2}(a_{2})}{a_{2}^{1+2\kappa}}\quad\text{if and only if}\quad\mu\geq 4.$
The next theorem establishes the asymptotic properties of $\hat{\sigma}_{n}^{2}(\hat{a}_{n})/\hat{a}_{n}^{1+2\kappa}$. The proof follows Theorem 9 in Bevilacqua et al. (2019) combined with Lemma 1.
Theorem 11.
Let $Z(\mathbf{s})$, $\mathbf{s}\in D\subset\mathbb{R}^{d}$, $d=1,2,3$, be a zero-mean Gaussian RF with covariance $\sigma^{2}\mathcal{H}_{\kappa,\mu,a,d}$, $\mu\geq 4$. Suppose $(\sigma_{0}^{2},a_{0})\in(0,\infty)\times I$. Then, as $n\to\infty$,
  1. 1.
    $\hat{\sigma}_{n}^{2}(\hat{a}_{n})/\hat{a}_{n}^{1+2\kappa}\stackrel{{\scriptstyle a.s.}}{{\longrightarrow}}\sigma_{0}^{2}/a_{0}^{1+2\kappa}$;
  2. 2.
    $\sqrt{n}\big(\hat{\sigma}_{n}^{2}(\hat{a}_{n})/\hat{a}_{n}^{1+2\kappa}-\sigma_{0}^{2}/a_{0}^{1+2\kappa}\big)\stackrel{{\scriptstyle\mathcal{D}}}{{\longrightarrow}}\mathcal{N}\big(0,2(\sigma_{0}^{2}/a_{0}^{1+2\kappa})^{2}\big)$.
The condition $\mu\geq 4$ ensures the monotonicity property in Lemma 1 and is technical in nature.
Finally, we compare the two reparameterized models $\sigma^{2}\mathcal{H}_{\kappa,\mu,b(\alpha,\mu,\kappa,d),d}$ and $\sigma^{2}\mathcal{H}_{\kappa,\mu,\alpha\mu,d}$. Under fixed-domain asymptotics and using (22), the associated microergodic parameters are
$\frac{\sigma^{2}}{\alpha^{2\kappa+1}}\quad\text{and}\quad\frac{\sigma^{2}B(\kappa,\mu,d)}{(\alpha\mu)^{2\kappa+1}},$
respectively. In the first case, the microergodic parameter does not depend on $\mu$, so both $\sigma^{2}/\alpha^{2\kappa+1}$ and $\mu$ are consistently estimable. From an inferential perspective, $\sigma^{2}\mathcal{H}_{\kappa,\mu,b(\alpha,\mu,\kappa,d),d}$ is therefore preferable. For large datasets, however, $\mu$ may be fixed to induce sparsity, in which case the microergodic parameter reduces to $\sigma^{2}/\alpha^{2\kappa+1}$ under both parameterizations.
7 A simulation study
We investigate the finite-sample performance of ML estimation for the model $\sigma^{2}\mathcal{H}_{\kappa,\mu,a,d}$. We simulate 500 realizations of a zero-mean Gaussian RF with covariance $\sigma^{2}\mathcal{H}_{\kappa,\mu,a,d}$ for $d=2$, $\sigma^{2}=1$, $\kappa=0,1,2$, and $\mu=4,6,8$. The process is generated via Cholesky decomposition on a regular grid with spacing $0.025$ over $0,1]^{2}$ (1681 locations), followed by independent perturbations of each coordinate by a uniform random variable on [0.01,0.01][-0.01,0.01].
Throughout the study, $\kappa$ and $\mu$ are assumed known. We focus on ML estimation of $\sigma^{2}$, $a$, and the microergodic parameter $\sigma^{2}/a^{2\kappa+1}$ under both increasing- and fixed-domain scenarios. Following Zhang and Zimmerman (2005); Kaufman and Shaby (2013), these regimes can be mimicked by fixing the sampling grid and varying the spatial dependence. Accordingly, we consider $a=0.1,0.25,0.5$, interpreting $a=0.1$ as increasing-domain and $a=0.5$ as fixed-domain behavior.
Table 1 reports the bias and standard deviation (in parentheses) of the 500 ML estimates $\widehat{\sigma}_{i}^{2}$, $\widehat{a}_{i}$, and $\widehat{\sigma}_{i}^{2}/\widehat{a}_{i}^{2\kappa+1}$, for $i=1,\ldots,500$.
Table 1: Bias and standard deviation (in parenthesis) of ML estimates $\widehat{\sigma}_{i}^{2}$, $\widehat{a}_{i}$, and $\widehat{\sigma}_{i}^{2}/\widehat{a}_{i}^{2\kappa+1}$ for $i=1,\ldots,500$ when estimating a zero-mean Gaussian RF with covariance $\sigma^{2}\mathcal{H}_{\kappa,\mu,a,2}$ when $\kappa=0,1,2$, $\mu=4,6,8$, $a=0.1,0.25,0.5$ and $\sigma^{2}=1$.
$\kappa=0$ $\kappa=1$ $\kappa=2$
$\mu=4$ $\mu=6$ $\mu=8$ $\mu=4$ $\mu=6$ $\mu=8$ $\mu=4$ $\mu=6$ $\mu=8$
$\widehat{\sigma}^{2}$ $a=0.1$ $0.0037$ $0.0025$ $0.0018$ $0.0060$ $0.0039$ $0.0028$ $0.0053$ $0.0038$ $0.0029$
$(0.0422)$ $(0.0382)$ $(0.0368)$ $(0.0485)$ $(0.0427)$ $(0.0398)$ $(0.0475)$ $(0.0432)$ $(0.0407)$
$a=0.25$ $0.0030$ $0.0047$ $0.0046$ $0.0085$ $0.0098$ $0.0095$ $0.0100$ $0.0116$ $0.0111$
$(0.0808)$ $(0.0622)$ $(0.0522)$ $(0.1082)$ $(0.0870)$ $(0.0737)$ $(0.1044)$ $(0.088)$ $(0.0773)$
$a=0.5$ 0.0023-0.0023 $0.0024$ $0.0045$ $0.0029$ $0.0079$ $0.0101$ $0.0052$ $0.0091$ $0.0116$
$(0.1446)$ $(0.1103)$ $(0.0900)$ $(0.1921)$ $(0.1592)$ $(0.135)$ $(0.1892)$ $(0.1599)$ $(0.1397)$
$\widehat{a}$ $a=0.1$ $0.0006$ $0.0007$ $0.0007$ $0.0004$ $0.0004$ $0.0004$ $0.0002$ $0.0003$ $0.0003$
$(0.0049)$ $(0.0055)$ $(0.0068)$ $(0.0024)$ $(0.0028)$ $(0.0033)$ $(0.0017)$ $(0.0021)$ $(0.0026)$
$a=0.25$ $0.0007$ $0.0014$ $0.0016$ $0.0005$ $0.0008$ $0.0009$ $0.0004$ $0.0006$ $0.0007$
$(0.0214)$ $(0.0176)$ $(0.0157)$ $(0.0099)$ $(0.0086)$ $(0.0079)$ $(0.0059)$ $(0.0054)$ $(0.0052)$
$a=0.5$ 0.0015-0.0015 $0.0010$ $0.0023$ 0.0016-0.0016 0.0001-0.0001 $0.0008$ 0.0009-0.0009 0.0001-0.0001 $0.0005$
$(0.0737)$ $(0.0578)$ $(0.0488)$ $(0.0329)$ $(0.0280)$ $(0.0245)$ $(0.0196)$ $(0.0171)$ $(0.0154)$
$\frac{\widehat{\sigma}^{2}}{\widehat{a}^{2\kappa+1}}$ $a=0.1$ 0.0150-0.0150 0.0219-0.0219 0.0151-0.0151 3.9684-3.9684 4.4120-4.4120 3.7659-3.7659 496.06-496.06 486.01-486.01 346.94-346.94
$(0.4370)$ $(0.5362)$ $(0.6792)$ $(51.3695)$ $(68.6746)$ $(88.9434)$ $(6397.8)$ $(8930.7)$ $(11628.7)$
$a=0.25$ $0.0045$ $0.0015$ 0.0017-0.0017 $0.0228$ 0.0470-0.0470 0.1218-0.1218 0.2553-0.2553 1.7130-1.7130 3.0083-3.0083
$(0.1499)$ $(0.1600)$ $(0.1708)$ $(2.5102)$ $(2.7402)$ $(2.9959)$ $(41.521)$ $(46.056)$ $(51.067)$
$a=0.5$ $0.0034$ $0.0033$ $0.0026$ $0.0113$ $0.0099$ $0.0065$ $0.0399$ $0.0319$ $0.0150$
$(0.0719)$ $(0.0740)$ $(0.0765)$ $(0.2923)$ $(0.3050)$ $(0.3187)$ $(1.1922)$ $(1.2484)$ $(1.3091)$
Overall, the ML estimators of $\sigma^{2}$ and $a$ are approximately unbiased across all configurations. As expected, their standard deviations increase when moving from the increasing-domain setting ($a=0.1$) to the fixed-domain setting ($a=0.5$). For fixed $\kappa$, larger values of $\mu$ generally reduce variability.
Regarding the microergodic parameter, under the increasing-domain scenario ($a=0.1$) the estimator exhibits moderate bias and relatively high variance, particularly for larger $\kappa$. Under the fixed-domain scenario ($a=0.5$), both bias and variance decrease substantially, and the sampling distribution appears closer to Gaussian, in agreement with the asymptotic theory. Moreover, the variability of the microergodic estimator increases with the smoothness parameter $\kappa$.
Figure 4 displays boxplots of the ML estimates of $\sigma^{2}$, $a$, and $\sigma^{2}/a^{2\kappa+1}$ for $\mu=4,6,8$ and $a=0.1$ (left) and $a=0.5$ (right), when $\kappa=0$. Under the increasing-domain scenario, the distributions of $\widehat{\sigma}^{2}$ and $\widehat{a}$ are approximately Gaussian. In the fixed-domain setting, the microergodic estimator becomes more concentrated around the true value and exhibits an approximately Gaussian distribution, as predicted by the theory.
Refer to caption Refer to caption
($a=0.1$) ($a=0.5$)
Figure 4: Left part: Boxplots of ML estimates of $\sigma^{2}$, $a$ and the microergodic parameter $\frac{\sigma^{2}}{a^{2\kappa+1}}$ (from left to right) when estimating the covariance model $\sigma^{2}\mathcal{H}_{\kappa,\mu,a,d}$ with $\sigma^{2}=1$, for $\mu=4,6,8$ and $a=0.1$ (increasing domain scenario). Right part: the same Figure with $a=0.5$ (fixed domain scenario). In both cases $\kappa=0$.
8 A real data application
We consider data from WorldClim (www.worldclim.org), which provides high-resolution gridded climate surfaces for 1970–2000 (Fick and Hijmans, 2017) and can be accessed in R via geodata (Hijmans et al., 2024). We analyze September mean temperature over Madagascar. The full dataset consists of 725,330 grid-cell predictions obtained by interpolating weather-station data using thin-plate splines with geographic and satellite-derived covariates (Fick and Hijmans, 2017). To make ML estimation computationally feasible, we randomly sample 3,500 grid cells and use UTM-projected coordinates.
Following Li and Zhang (2011), the data are first detrended using spline regression to remove large-scale spatial trends. The residuals $y(\mathbf{s}_{i}),i=1,\ldots,3500$, are then treated as observations from a zero-mean RF $\{Y(\mathbf{s}):\mathbf{s}\in D\}$. Figure 5 displays the spatial map of the residuals and the empirical semivariogram. The plots suggest that a weakly stationary RF with a potentially heavy-tailed marginal distribution provides a reasonable modeling framework.
Refer to caption Refer to caption
Figure 5: Left: spatial map of residuals. Right: empirical semivariogram.
Let $Z(\mathbf{s})$ be a standard Gaussian RF with isotropic correlation function. We consider two specifications for $Y(\mathbf{s})$:
  • Gaussian RF: $Y(\mathbf{s})=\sigma Z(\mathbf{s})$;
  • Tukey-$h$ RF (Xu and Genton, 2017; Caamaño-Carrillo et al., 2024): $Y(\mathbf{s})=\sigma Z(\mathbf{s})\exp\!\left(\tfrac{hZ^{2}(\mathbf{s})}{2}\right)$, where $h\in[0,1/2)$ controls tail heaviness. Since the transformation is monotone, the Tukey-$h$ RF preserves the geometric properties of the underlying Gaussian field.
For the correlation of $Z(\mathbf{s})$, we compare three models: ${\cal MT}_{\kappa,\beta}$, ${\cal GW}_{\kappa,\mu,\beta}$, and the proposed ${\cal H}_{\kappa,\mu,\beta,2}$. Here $\kappa$ denotes smoothness, $\beta$ the scale parameter, and $\mu$ the shape parameter.
We adopt an increasing-domain framework and estimate all covariance parameters by ML. Standard errors are computed from the inverse Fisher information matrix. For the Tukey-$h$ RF, ML estimation is implemented via the Lambert-$W$ function (see equation 10 in Caamaño-Carrillo et al. (2024)).
Table 2 reports ML estimates, associated standard errors, maximized log-likelihood values, and AIC values. Overall, the AIC criterion favors the Tukey-$h$ specification, irrespective of the correlation model. Among the correlation structures, the compactly supported ${\cal GW}_{\kappa,\mu,\beta}$ and ${\cal H}_{\kappa,\mu,\beta,2}$ achieve the best AIC values.
Figure 6 compares the normalized histogram of the residuals with the fitted density (left panels) and the empirical semivariogram with the fitted semivariogram (right panels), for the Gaussian (first row) and Tukey-$h$ (second row) models using the estimated ${\cal H}_{\kappa,\mu,\beta,2}$ correlation. The Tukey-$h$ RF provides a noticeably better fit to both marginal tail behavior and spatial dependence.
Table 2: ML estimates with associated standard error (in parentheses), RMSE, MAE and percentage of zero values in the covariance matrix for the Gaussian and Tukey-$h$ RFs with underlying correlation functions ${\cal MT}_{\kappa,\beta}$, ${\cal GW}_{\kappa,\mu,\beta}$ and ${\cal H}_{\kappa,\mu,\beta,2}$. Last row reports the time (in seconds) needed to compute the Cholesky factor of the covariance matrix.
Gaussian RF Tukey-$h$ RF
${\cal MT}_{\kappa,\beta}$ ${\cal GW}_{\kappa,\mu,\beta}$ ${\cal H}_{\kappa,\mu,\beta,2}$ ${\cal MT}_{\kappa,\beta}$ ${\cal GW}_{\kappa,\mu,\beta}$ ${\cal H}_{\kappa,\mu,\beta,2}$
$\sigma^{2}$ $1.2054$ $1.2066$ $1.2061$ $0.6998$ $0.6986$ $0.6984$
$(0.0553)$ $(0.0541)$ $(0.0540)$ $(0.0418)$ $(0.0403)$ $(0.0403)$
$\kappa$ $0.6079$ $0.0305$ $0.0313$ $0.5767$ 0.0036-0.0036 0.0043-0.0043
$(0.0293)$ $(0.0307)$ $(0.0312)$ $(0.0265)$ $(0.0237)$ $(0.0237)$
$\beta$ $16.0229$ $85.8058$ $87.9751$ $18.5405$ $76.0127$ $77.6552$
$(1.3104)$ $(20.4440)$ $(20.9644)$ $(1.6183)$ $(12.0024)$ $(10.5911)$
$\mu$ $4.2150$ $3.8914$ $3.1346$ $2.8085$
$(1.3021)$ $(1.3026)$ $(0.6548)$ $(0.5764)$
$h$ $0.1885$ $0.1899$ $0.1898$
$(0.0172)$ $(0.0170)$ $0.0170$
loglik 3900.01-3900.01 3896.88-3896.88 3896.94-3896.94 3689.88-3689.88 3683.81-3683.81 3683.94-3683.94
AIC $7806$ $7802$ $7802$ $7388$ $7378$ $7378$
RMSE $0.6593$ $0.6588$ $0.6588$ $0.6586$ $0.6581$ $0.6581$
MAE $0.4068$ $0.4065$ $0.4065$ $0.4073$ $0.4071$ $0.4070$
$\%$ 0 $96.45$ $96.23$ 0 $97.18$ $97.06$
TimeTime $7.22$ $0.04$ $0.04$ $6.26$ $0.03$ $0.03$
We assess predictive performance via repeated cross-validation. For each of 100 repetitions, 85% of locations are used for parameter estimation and the remaining 15% for prediction. Table 2 reports the average RMSE and MAE. The best predictive performance is achieved by the Tukey-$h$ RF combined with compactly supported correlations ${\cal GW}_{\kappa,\mu,\beta}$ and ${\cal H}_{\kappa,\mu,\beta,2}$.
[…]
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 6: First row: normalized histogram versus estimated density and empirical semivariogram versus estimated semivariogram when using an estimated Gaussian RF with $\sigma^{2}\mathcal{H}_{\kappa,\mu,a,2}$ correlation model (from left to right). Second row: the same graphics using a Tukey-$h$ RF.
9 Concluding Remarks
In this paper we study the Gauss hypergeometric covariance model, a flexible class introduced by Emery and Alegría (2022) that encompasses both the Matérn and Generalized Wendland families. We derive necessary and sufficient validity conditions and show that the model exhibits structural identifiability issues under both increasing- and fixed-domain asymptotics. To address this issue, we introduce a parsimonious compactly supported subclass selected through an optimality criterion based on maximum integral range. In this sense, the proposed hypergeometric model can be viewed as a structural refinement of the Generalized Wendland construction.
Although simple closed-form expressions are not generally available for nonnegative integer smoothness parameters, efficient numerical evaluation of the Gauss hypergeometric function is widely supported in modern scientific computing libraries. In particular, reliable implementations are available in the GNU Scientific Library (Gough, 2009) and in major environments such as R, MATLAB, and Python. The GeoModels package (Bevilacqua et al., 2026), used in this paper for numerical experiments and applications, computes the proposed model via the SciPy implementation of the hypergeometric function (Virtanen et al., 2020).
Acknowledgement
This work was funded by the National Agency for Research and Development of Chile, under grants ANID FONDECYT 1240308 (M. Bevilacqua), ANID project Data Observatory Foundation DO210001 (M. Bevilacqua), MATH-AMSUD1167 22-MATH-06 (AMSUD220041) (M. Bevilacqua), ANID FONDECYT 1250862 (C. Caamaño-Carrillo) and ANID PIA AFB230001 (X. Emery).
Appendix A Proofs
A.1 Proof of Theorem 1
Proof.
Let $P(\rho_{i})$, $i=0,1$, be two zero-mean Gaussian measures with isotropic covariance function $\rho_{i}$ and associated spectral density $\widehat{\rho}_{i}$, $i=0,1$, as defined through (1).
Using Skorokhod and Yadrenko (1973); Ibragimov and Rozanov (1978); Stein (2004), if for some $a>0$ the function $\widehat{\rho}_{0}(z)z^{a}$ is bounded away from 0 and $\infty$ as $z\to\infty$, and if for some finite and positive $c$,
$\int_{c}^{\infty}z^{d-1}\left(\frac{\widehat{\rho}_{1}(z)-\widehat{\rho}_{0}(z)}{\widehat{\rho}_{0}(z)}\right)^{2}\,{\rm d}z<\infty,$ (26)
then for any bounded subset $D\subset\mathbb{R}^{d}$, $P(\rho_{0})$ and $P(\rho_{1})$ are equivalent Gaussian measures on the paths of $Z(\mathbf{s})$, $\mathbf{s}\in D$. Condition (26) is the key tool.
We first need the following result.
Lemma 2.
Let $\widehat{{\cal GH}}_{\delta,\beta,\gamma,a}(z)$ be defined at (7). Assume $\delta>d/2$, $2(\beta-\delta)(\gamma-\delta)\geq\delta$ and $2(\beta+\gamma)\geq 6\delta+1$. Then:
  1. 1.
    $\widehat{{\cal GH}}_{\delta,\beta,\gamma,a}(z)$ $=\sigma^{2}Ka^{d}\Big[c_{3}(za)^{-2\delta}\{1+\mathcal{O}(z^{-2})\}$
    $\qquad+c_{4}(za)^{\frac{1}{2}+\delta-\beta-\gamma}\{\cos(za+c_{5})+\mathcal{O}(z^{-1})\}\Big],$
    as $z\to\infty$;
  2. 2.
    $\widehat{{\cal GH}}_{\delta,\beta,\gamma,a}(z)\asymp z^{-2\delta},\qquad z\to\infty.$
Here
$c_{3}=\frac{2^{2\delta}\Gamma(\beta)\Gamma(\gamma)}{\Gamma(\gamma-\delta)\Gamma(\beta-\delta)},\qquad c_{4}=\frac{\pi^{-1/2}\Gamma(\beta)\Gamma(\gamma)}{2^{\frac{1}{2}+\delta-\beta-\gamma}\Gamma(\delta)},$
$c_{5}=\frac{\pi}{2}\left(\frac{1}{2}+\delta-\beta-\gamma\right).$
We are now ready to prove the theorem.
From Lemma 2 (Point 2), there exist constants $0<c_{1}<c_{2}<\infty$ such that, for $z$ large,
$c_{1}\leq z^{2\delta}\widehat{{\cal GH}}_{\delta,\beta_{i},\gamma_{i},a_{i}}(z)\leq c_{2},\qquad i=0,1.$ (27)
In particular, $\widehat{\rho}_{0}(z)z^{2\delta}$ is bounded away from 0 and $\infty$ as $z\to\infty$.
Using Lemma 2 (Points 1 and 2), and collecting terms, one obtains for large $z$
$\left|\frac{\sigma_{1}^{2}\widehat{\varphi}_{1}(z)-\sigma_{0}^{2}\widehat{\varphi}_{0}(z)}{\sigma_{0}^{2}\widehat{\varphi}_{0}(z)}\right|\leq\frac{1}{c_{1}}\left|A(z)+B(z)+D(z)\right|,$
where
$A(z)$ $=\big[c_{3}^{1}\sigma_{1}^{2}a_{1}^{d-2\delta}K_{1}-c_{3}^{0}\sigma_{0}^{2}a_{0}^{d-2\delta}K_{0}\big]\{1+\mathcal{O}(z^{-2})\},$
$B(z)$ $=\big[c_{4}^{1}z^{\lambda_{1}+2\delta}\sigma_{1}^{2}a_{1}^{\lambda_{1}+d}K_{1}\cos(a_{1}z+c_{5}^{1})$
$\quad-c_{4}^{0}z^{\lambda_{0}+2\delta}\sigma_{0}^{2}a_{0}^{\lambda_{0}+d}K_{0}\cos(a_{0}z+c_{5}^{0})\big],$
$D(z)$ $=\mathcal{O}(z^{-1})\big[c_{4}^{1}z^{\lambda_{1}+2\delta}\sigma_{1}^{2}a_{1}^{\lambda_{1}+d}K_{1}+c_{4}^{0}z^{\lambda_{0}+2\delta}\sigma_{0}^{2}a_{0}^{\lambda_{0}+d}K_{0}\big].$
Hence condition (26) becomes
$\int_{c}^{\infty}z^{d-1}\big(A(z)+B(z)+D(z)\big)^{2}\,{\rm d}z<\infty.$
Assume now (10), i.e.
$\frac{\sigma_{0}^{2}L(\beta_{0},\gamma_{0})}{a_{0}^{2\delta-d}}=\frac{\sigma_{1}^{2}L(\beta_{1},\gamma_{1})}{a_{1}^{2\delta-d}}.$
This is equivalent to
$c_{3}^{1}\sigma_{1}^{2}a_{1}^{d-2\delta}K_{1}=c_{3}^{0}\sigma_{0}^{2}a_{0}^{d-2\delta}K_{0},$
so that $A(z)\equiv 0$.
Using $(x+y)^{2}\leq 2x^{2}+2y^{2}$, it suffices to check
$\int_{c}^{\infty}z^{d-1+2(\lambda_{i}+2\delta)}\,dz<\infty,\qquad i=0,1,$
which holds iff
$\beta_{i}+\gamma_{i}>\frac{d+1}{2}+3\delta.$
Therefore, under
$\min(\beta_{1}+\gamma_{1},\beta_{0}+\gamma_{0})>\frac{d+1}{2}+3\delta,$
condition (26) holds and the sufficient part is proved.
The necessary part follows from Zhang (2004).
A.2 Proof of Theorem 2
Proof.
Consider the model $\mathcal{GH}_{\delta,\delta+\frac{\mu}{2},\delta+\frac{\mu}{2}+l,a}$ in dimension $d$, where $\delta=\frac{d+1}{2}+\kappa$, $\kappa>-\frac{1}{2}$, $l\geq 0$.
Sufficiency. Under $\beta=\delta+\frac{\mu}{2}$, $\gamma=\delta+\frac{\mu}{2}+l$, the sufficient conditions of Section 3 become:
  1. (a)
    $\kappa>-\frac{1}{2}$,
  2. (b)
    $\mu(\mu+2l)\geq d+2\kappa+1$,
  3. (c)
    $\mu+l\geq\frac{d}{2}+\kappa+1$.
Condition (b) is equivalent to the quadratic inequality
$\mu^{2}+2l\mu-(d+2\kappa+1)\geq 0,$
whose positive branch gives
$\mu\geq-l+\sqrt{l^{2}+d+2\kappa+1}.$
Only the positive branch is admissible since condition (c) implies $\mu>-l$ when $l\geq 0$. For each fixed $l$, the sufficient conditions therefore reduce to
$\mu\geq\max\!\Big\{\frac{d}{2}+\kappa+1-l,\;-l+\sqrt{l^{2}+d+2\kappa+1}\Big\}.$
A direct algebraic comparison shows that
$\frac{d}{2}+\kappa+1-l\geq-l+\sqrt{l^{2}+d+2\kappa+1}\quad\Longleftrightarrow\quad 0\leq l\leq\frac{d}{2}+\kappa.$
Hence condition (c) is binding when $0\leq l\leq\frac{d}{2}+\kappa$, giving $\mu\geq\delta-l+\frac{1}{2}$, and condition (b) is binding when $l>\frac{d}{2}+\kappa$, giving $\mu\geq\sqrt{2\kappa+l^{2}+d+1}-l$. This yields exactly the two regimes in points 1 and 2.
Necessity. Assume $\mathcal{GH}_{\delta,\beta,\gamma,a}\in\Phi_{d}$. Since the Gamma prefactors in the spectral density are positive whenever $\beta>\delta$ and $\gamma>\delta$, nonnegativity of the spectral density requires
$_{1}F_{2}\!\left(\delta;\beta,\gamma;-\frac{(za)^{2}}{4}\right)\geq 0\quad\text{for all }z\geq 0.$
By Cho and Yun (2018, Theorem 3.1), this implies $\beta>\delta$, $\gamma>\delta$, and $\beta+\gamma\geq 3\delta+\frac{1}{2}$. The third condition, under the parameterization $\beta=\delta+\frac{\mu}{2}$, $\gamma=\delta+\frac{\mu}{2}+l$, gives
$\mu+l\geq\frac{d}{2}+\kappa+1,\quad\text{i.e.,}\quad\mu\geq\delta-l+\frac{1}{2},$
which is the necessary lower bound for $\mu$ in both regimes. In Regime 1 ($0\leq l\leq\frac{d}{2}+\kappa$), this coincides with the sufficient condition in point 1, so the result is an if and only if. In Regime 2 ($l>\frac{d}{2}+\kappa$), this provides a necessary lower bound that is strictly weaker than the sufficient condition in point 2; whether the gap can be closed remains an open question.
Reduction to $l\geq 0$. The model $\mathcal{GH}_{\delta,\beta,\gamma,a}$ is invariant under interchanging $\beta$ and $\gamma$, which corresponds to $l\mapsto-l$. The admissible region is therefore symmetric with respect to $l=0$, and restricting to $l\geq 0$ entails no loss of generality.
This completes the proof. ∎
A.3 Proof of Theorem 3
Proof.
Without loss of generality set $a=1$; the general case follows by the substitution $x\mapsto ax$, which contributes the factor $a^{d}$. Using the change of variable $t=1-x^{2}$ (so $x^{d-1}\,dx=\tfrac{1}{2}(1-t)^{(d-2)/2}\,dt$) in the definition of the integral range (12), and substituting (8), one obtains
$A_{\kappa,\mu,l,1,d}$ $=\frac{\pi^{d/2}C_{0}}{2^{d}\Gamma(d/2)}\int_{0}^{1}(1-t)^{\frac{d-2}{2}}\,t^{\mu+\frac{d-1}{2}+2\kappa}$ (28)
$\quad\times{}_{2}F_{1}\!\left(\frac{\mu}{2},\frac{\mu+d}{2}+\kappa;\mu+\frac{d+1}{2}+2\kappa;t\right)\,dt.$
where $C_{0}=\Gamma\!\left(\kappa+\frac{\mu+1}{2}\right)\Gamma\!\left(2\kappa+\frac{d+\mu+1}{2}\right)\big/\!\left[\Gamma\!\left(\mu+\frac{d+1}{2}+2\kappa\right)\Gamma\!\left(\kappa+\tfrac{1}{2}\right)\right]$.
We apply formula 2.21.1.11 of Prudnikov et al. (1990),
$\int_{0}^{1}t^{\sigma-1}(1-t)^{\rho-1}\,{}_{2}F_{1}(a,b;c;t)\,dt=\frac{\Gamma(\rho)\Gamma(\sigma)}{\Gamma(\rho+\sigma)}\,{}_{3}F_{2}(a,b,\sigma;\,c,\,\rho+\sigma;\,1),$ (29)
with the identifications
$\sigma=\mu+\tfrac{d+1}{2}+2\kappa,\quad\rho=\tfrac{d}{2},\quad a=\tfrac{\mu}{2},\quad b=\tfrac{\mu+d}{2}+\kappa,\quad c=\mu+\tfrac{d+1}{2}+2\kappa.$
Since $c=\sigma$, the third upper parameter and the first lower parameter of $_{3}F_{2}$ coincide, so the series telescopes term by term:
$_{3}F_{2}(a,b,\sigma;\,\sigma,\,\rho+\sigma;\,1)=\sum_{k=0}^{\infty}\frac{(a)_{k}(b)_{k}(\sigma)_{k}}{k!\,(\sigma)_{k}(\rho+\sigma)_{k}}=\sum_{k=0}^{\infty}\frac{(a)_{k}(b)_{k}}{k!\,(\rho+\sigma)_{k}}={}_{2}F_{1}(a,b;\,\rho+\sigma;\,1).$
Gauss’s summation theorem then gives, provided $\rho+\sigma-a-b>0$,
$_{2}F_{1}(a,b;\rho+\sigma;1)=\frac{\Gamma(\rho+\sigma)\,\Gamma(\rho+\sigma-a-b)}{\Gamma(\rho+\sigma-a)\,\Gamma(\rho+\sigma-b)}.$
One verifies directly that
$\rho+\sigma-a-b=\kappa+\tfrac{1}{2}+l>0\quad(\kappa>-\tfrac{1}{2},\;l\geq 0),$
$\rho+\sigma-a=\tfrac{\mu+1}{2}+\kappa+d,\qquad\rho+\sigma-b=\tfrac{\mu+1}{2}+\kappa+\tfrac{d}{2}.$
Substituting back into (28), the factors $\Gamma(\sigma)$ and $\Gamma(\rho+\sigma)$ cancel with those arising from (29) and from $C_{0}$. Using the duplication formula $\Gamma(z)\Gamma(z+\tfrac{1}{2})=\sqrt{\pi}\,2^{1-2z}\Gamma(2z)$ to simplify the remaining Gamma factors, one obtains
$A_{\kappa,\mu,l,1,d}=\frac{\pi^{d/2}\,\Gamma\!\left(\frac{d+1}{2}+\kappa\right)\Gamma\!\left(\frac{\mu+1}{2}+\kappa\right)\Gamma\!\left(\frac{\mu+1}{2}+\kappa+l\right)}{2^{d}\,\Gamma\!\left(\kappa+\frac{1}{2}\right)\Gamma\!\left(\frac{\mu+1+d}{2}+\kappa\right)\Gamma\!\left(\frac{\mu+1+d}{2}+\kappa+l\right)},$
which, after restoring the factor $a^{d}$, gives (14). ∎
A.4 Proof of Theorem 4
Proof.
Without loss of generality, set the compact support to $a=1$. Define
$LA_{\kappa,d}(\mu,l):=\log(A_{\kappa,\mu,l,d,1}),$
where
$A_{\kappa,\mu,l,d,1}=\frac{\pi^{d/2}\Gamma((d+1)/2+\kappa)\,\Gamma\!\left(\frac{\mu+1}{2}+\kappa\right)\,\Gamma\!\left(\frac{\mu+1}{2}+\kappa+l\right)}{2^{d}\Gamma(\kappa+0.5)\,\Gamma\!\left(\frac{\mu+1}{2}+\kappa+\frac{d}{2}\right)\,\Gamma\!\left(\frac{\mu+1}{2}+\kappa+l+\frac{d}{2}\right)}.$
We maximize $LA_{\kappa,d}(\mu,l)$ over the admissible set in $\mu$ for a fixed $l\geq 0$, as given by Theorem 2. Let $\delta=\frac{d}{2}+\kappa$. Then
  1. 1.
    $\mu\geq\delta-l+\tfrac{1}{2}$ if $0\leq l\leq\delta$;
  2. 2.
    $\mu\geq\sqrt{2\kappa+l^{2}+d+1}-l$ if $l>\delta$.
This is a maximization problem over the half-line $\mu_{\min}(l),\infty)$.
Let $\psi(x)=\frac{d}{dx}\log\Gamma(x)$ be the digamma function, which is strictly increasing for $x>0$. Set $x=\frac{\mu+1}{2}+\kappa$. Under the validity conditions we have $x>0$ and $x+l>0$. Differentiating with respect to $\mu$ yields
$\frac{d}{d\mu}LA_{\kappa,d}(\mu,l)=\frac{1}{2}\Big[\psi(x)-\psi(x+\tfrac{d}{2})+\psi(x+l)-\psi(x+l+\tfrac{d}{2})\Big].$
Since $\psi$ is increasing and $\tfrac{d}{2}>0$, we have $\psi(x)-\psi(x+\tfrac{d}{2})\leq 0$ and $\psi(x+l)-\psi(x+l+\tfrac{d}{2})\leq 0$, hence $\frac{d}{d\mu}LA_{\kappa,d}(\mu,l)\leq 0$ for all admissible $\mu$. Therefore $LA_{\kappa,d}(\mu,l)$ is non-increasing in $\mu$, and its maximum is achieved at the smallest admissible value of $\mu$, i.e.
$\mu_{1}(l)=\delta-l+\tfrac{1}{2}\quad(0\leq l\leq\delta),\qquad\mu_{2}(l)=\sqrt{2\kappa+l^{2}+d+1}-l\quad(l>\delta).$
This proves Theorem 4. ∎
A.5 Proof of Theorem 5
Proof.
Let $\delta=\frac{d}{2}+\kappa$. Since $\log$ is strictly increasing, maximizing $A_{\kappa,\mu,l,a,d}$ is equivalent to maximizing
$LA_{\kappa,d}(\mu,l):=\log A_{\kappa,\mu,l,d,1},$
and the factor $a^{d}$ plays no role for maximizers.
By Theorem 4, for each fixed $l$ the maximizer in $\mu$ is $\mu=\mu_{1}(l)$ for $0\leq l\leq\delta$ and $\mu=\mu_{2}(l)$ for $l>\delta$. Hence it suffices to study the two one–variable functions
$F_{1}(l):=LA_{\kappa,d}(\mu_{1}(l),l)\quad\text{on }[0,\delta],\qquad F_{2}(l):=LA_{\kappa,d}(\mu_{2}(l),l)\quad\text{on }(\delta,\infty).$
Write $\psi=\Gamma^{\prime}/\Gamma$ for the digamma function. Recall that $\psi$ is concave on $(0,\infty)$, so for every $h>0$ the symmetric increment
$\Delta_{h}(t):=\psi(t+h)-\psi(t-h)$
is decreasing in $t$ (because $\Delta_{h}^{\prime}(t)=\psi^{\prime}(t+h)-\psi^{\prime}(t-h)\leq 0$).
Case $i=1$ ($0\leq l\leq\delta$).
For $0\leq l\leq\delta$, $\mu_{1}(l)=\delta-l+\tfrac{1}{2}$. Set
$u(l):=\frac{\mu_{1}(l)+1}{2}+\kappa=\frac{d}{4}+\frac{3\kappa}{2}+\frac{3}{4}-\frac{l}{2}.$
Using the explicit form of $A_{\kappa,\mu,l,d,1}$, $F_{1}(l)$ differs from
$\log\Gamma(u(l))+\log\Gamma(u(l)+l)-\log\Gamma\Big(u(l)+\frac{d}{2}\Big)-\log\Gamma\Big(u(l)+l+\frac{d}{2}\Big)$
only by an additive constant independent of $l$. Differentiating gives
$F_{1}^{\prime}(l)$ $=u^{\prime}(l)\Big[\psi(u(l))+\psi(u(l)+l)-\psi\Big(u(l)+\frac{d}{2}\Big)-\psi\Big(u(l)+l+\frac{d}{2}\Big)\Big]$
$\qquad+\psi(u(l)+l)-\psi\Big(u(l)+l+\frac{d}{2}\Big),$
where $u^{\prime}(l)=-\tfrac{1}{2}$. A simple rearrangement yields
$F_{1}^{\prime}(l)=\frac{1}{2}\Big[\psi(u(l)+l)-\psi(u(l))+\psi\Big(u(l)+\frac{d}{2}\Big)-\psi\Big(u(l)+l+\frac{d}{2}\Big)\Big].$
Now set
$x:=\frac{d}{4}+\frac{3\kappa}{2}+\frac{3}{4}\quad\Longrightarrow\quad u(l)=x-\frac{l}{2},\qquad u(l)+l=x+\frac{l}{2}.$
Therefore
$F_{1}^{\prime}(l)=\frac{1}{2}\Big[\psi(x+\tfrac{l}{2})-\psi(x-\tfrac{l}{2})+\psi(x-\tfrac{l}{2}+\tfrac{d}{2})-\psi(x+\tfrac{l}{2}+\tfrac{d}{2})\Big]=\frac{1}{2}\Big(\Delta_{l/2}(x)-\Delta_{l/2}(x+\tfrac{d}{2})\Big).$
Since $\Delta_{l/2}$ is decreasing and $x+\tfrac{d}{2}>x$, we have $\Delta_{l/2}(x)\geq\Delta_{l/2}(x+\tfrac{d}{2})$, hence $F_{1}^{\prime}(l)\geq 0$ on $0,\delta]$. Thus $F_{1}$ is non-decreasing on $0,\delta]$ and its maximum is attained at $l=\delta$.
Case $i=2$ ($l>\delta$).
For $l>\delta$, $\mu_{2}(l)=\sqrt{l^{2}+(2\kappa+d+1)}-l$. Define
$u(l):=\frac{\mu_{2}(l)+1}{2}+\kappa>0,\qquad x(l):=u(l)+\frac{l}{2}=\kappa+\frac{1+\sqrt{l^{2}+(2\kappa+d+1)}}{2}.$
Then $u(l)=x(l)-\tfrac{l}{2}$ and $u(l)+l=x(l)+\tfrac{l}{2}$. As above, up to an additive constant independent of $l$,
$F_{2}(l)=\log\Gamma(u(l))+\log\Gamma(u(l)+l)-\log\Gamma\Big(u(l)+\frac{d}{2}\Big)-\log\Gamma\Big(u(l)+l+\frac{d}{2}\Big).$
Differentiating (same computation as in Case $i=1$ but with general $u$) gives the identity
$F_{2}^{\prime}(l)=\Big(x^{\prime}(l)\Big)\Big(\Delta_{l/2}(x(l))-\Delta_{l/2}(x(l)+\tfrac{d}{2})\Big)-\frac{1}{2}\Big(\Delta_{l/2}(x(l))-\Delta_{l/2}(x(l)+\tfrac{d}{2})\Big),$
i.e.
$F_{2}^{\prime}(l)=\Big(x^{\prime}(l)-\frac{1}{2}\Big)\Big(\Delta_{l/2}(x(l))-\Delta_{l/2}(x(l)+\tfrac{d}{2})\Big).$
Now observe that
$x^{\prime}(l)=\frac{1}{2}\cdot\frac{l}{\sqrt{l^{2}+(2\kappa+d+1)}}\in\Big(0,\frac{1}{2}\Big),$
so $x^{\prime}(l)-\frac{1}{2}<0$ for all $l>\delta$. On the other hand, as in Case $i=1$, concavity of $\psi$ implies $\Delta_{l/2}(x(l))-\Delta_{l/2}(x(l)+\tfrac{d}{2})\geq 0$. Therefore $F_{2}^{\prime}(l)\leq 0$ for all $l>\delta$. Hence $F_{2}$ is non-increasing on $(\delta,\infty)$ and
$\sup_{l>\delta}LA_{\kappa,d}(\mu_{2}(l),l)=\lim_{l\downarrow\delta}LA_{\kappa,d}(\mu_{2}(l),l),$
equivalently
$\sup_{l>\delta}A_{\kappa,\mu_{2}(l),l,a,d}=\lim_{l\downarrow\delta}A_{\kappa,\mu_{2}(l),l,a,d}.$
Finally, $\mu_{2}(l)\to 1$ as $l\downarrow\delta$ and $\mu_{1}(\delta)=1$. By continuity of $LA_{\kappa,d}(\mu,l)$ on the admissible region,
$\lim_{l\downarrow\delta}LA_{\kappa,d}(\mu_{2}(l),l)=LA_{\kappa,d}(1,\delta)=LA_{\kappa,d}(\mu_{1}(\delta),\delta).$
Therefore the maximum of $A_{\kappa,\mu_{1}(l),l,a,d}$ over $0\leq l\leq\delta$ is attained at $l=\delta$, and the supremum of $A_{\kappa,\mu_{2}(l),l,a,d}$ over $l>\delta$ equals the same value and is approached as $l\downarrow\delta$. Consequently, the global maximum over the admissible ranges is achieved at $(\mu,l)=(1,\delta)$, while in the regime $l>\delta$ it is not attained but only approached in the limit $l\downarrow\delta$. This proves Theorem 5. ∎
A.6 Proof of Theorem 6
Proof.
Using (7) and Equation (25) in Emery and Alegría (2022), the Fourier transform of the right-hand side member of (18) is found to be
$\begin{split}&\frac{\Gamma(\delta+\frac{\mu-d}{2})\Gamma(\delta+\frac{\mu-d}{2}+l)\Gamma(\delta)a^{d}}{2^{d}\pi^{\frac{d}{2}}\Gamma\left(\delta-\frac{d}{2}\right)\Gamma(\delta+\frac{1}{2})\Gamma(2\delta)\Gamma(\frac{\mu-1}{2})\Gamma(\frac{\mu-1-d}{2}-\kappa+l)}\\ &\times\int_{0}^{1}\int_{0}^{1}u^{\delta-\frac{1}{2}}(1-u)^{\frac{\mu-3}{2}}v^{d+2\kappa}(1-v)^{\frac{\mu-d-3}{2}-\kappa+l}\mathstrut_{1}F_{2}\Big(\delta;\delta+\frac{1}{2},2\delta;-\frac{uv(za)^{2}}{4}\Big){\rm d}u{\rm d}v\\ &=\frac{\Gamma(\delta)\Gamma(\delta+\frac{\mu-d}{2})\Gamma(\delta+\frac{\mu-d}{2}+l)a^{d}}{2^{d}\pi^{\frac{d}{2}}\Gamma\left(\delta-\frac{d}{2}\right)\Gamma(\delta+\frac{\mu}{2})\Gamma(\delta+\frac{\mu}{2}+l)}\mathstrut_{1}F_{2}\Big(\delta;\delta+\frac{\mu}{2},\delta+\frac{\mu}{2}+l;-\frac{(za)^{2}}{4}\Big)\\ &=\widehat{\mathcal{GH}}_{\delta,\delta+\frac{\mu}{2},\delta+\frac{\mu}{2}+l,a}(z),\quad z\geq 0.\end{split}$
The claim of the theorem is obtained by inverse Fourier transformation. ∎
A.7 Proof of Theorem 7
If $d=1$ and $\kappa=k$ is a nonnegative integer, using formula 7.3.1.70 of Prudnikov et al. (1990) and applying the identity
$P_{a}^{b}(x)=P^{b}_{-a-1}(x),$
where $P_{a}^{b}(x)$ denotes the associated Legendre function (of the first kind) with degree $a$ and order $b$ (see Abramowitz and Stegun (1970, 8.3.1)), one can rewrite (17) as
$\mathcal{H}_{k,\mu,a,1}(x)$ $=$ $\frac{\Gamma(\frac{\mu}{2}+2k+1)}{\Gamma(k+\frac{1}{2})}\sqrt{\pi}\left(1-\frac{x^{2}}{a^{2}}\right)^{\frac{\mu}{2}+k}\left(\frac{x}{a}\right)^{k}P^{-\frac{\mu}{2}-k}_{k}\left(\frac{a^{2}+x^{2}}{2ax}\right).$
The claim of the Theorem then derives from the following result, which is valid for $a$ being a nonnegative integer and $z>1$:
$P^{-b}_{a}(z)=\frac{(z-1)^{b/2}}{(z+1)^{b/2}}\sum_{j=0}^{a}\frac{(j+a)!(0.5(z-1))^{j}}{j!\Gamma(j+b+1)(a-j)!}.$ (31)
A.8 Proof of Theorem 8
Proof.
For $\mu\geq 1$ and $\kappa=0$, applying formula 7.3.1.41 from Prudnikov et al. (1990) to Equation (17), one gets
$\mathcal{H}_{0,\mu,a,d}(x)=\frac{\Gamma(\frac{1+\mu}{2})\Gamma(\frac{1+\mu+d}{2})}{\sqrt{\pi}}2^{\mu+\frac{d-1}{2}}\left(1-\frac{x^{2}}{a^{2}}\right)_{+}^{\frac{\mu}{2}+\frac{d-1}{4}}\,P_{\frac{d-1}{2}}^{\frac{1-d}{2}-\mu}\left(\frac{x}{a}\right).$
The claim of the Theorem stems from the fact that, for $m$ being an integer, the associated Legendre function of degree $m$ is a terminating series:
$P_{m}^{\eta}(z)=\left(\frac{1+z}{1-z}\right)^{\frac{\eta}{2}}\sum\limits_{n=0}^{m}\frac{(-m)_{n}(m+1)_{n}}{n!\Gamma(n-\eta+1)}\left(\frac{1-z}{2}\right)^{n},\quad 0<z<1.$ (32)
A.9 Proof of Theorem 9
Proof.
For notational convenience, we use the representation ${\cal GH}_{\delta,\beta,\gamma,a}$ for the model $\mathcal{H}_{\kappa,\mu,a,d}$. Recall (see 16) that if
$\delta=\frac{d+1}{2}+\kappa,\qquad\beta=\delta+\frac{\mu}{2},\qquad\gamma=\delta+\frac{\mu}{2}+\frac{d}{2}+\kappa,$ (33)
then
$\mathcal{H}_{\kappa,\mu,a,d}=\mathcal{GH}_{\delta,\beta,\gamma,a}.$
To prove Theorem 9, we proceed as in the proof of Theorem 1. Using (26), we seek conditions such that, for some positive and finite $c$,
$\int_{c}^{\infty}z^{d-1}\left(\frac{\sigma_{1}^{2}\widehat{{\cal GH}}_{\delta,\beta,\gamma,a}(z)-\sigma_{0}^{2}\widehat{{\cal MT}}_{\nu,\alpha}(z)}{\sigma_{0}^{2}\widehat{{\cal MT}}_{\nu,\alpha}(z)}\right)^{2}{\rm d}z<\infty.$ (34)
It is known that $\sigma_{0}^{2}\widehat{{\cal MT}}_{\nu,\alpha}(z)z^{\tau}$ is bounded away from 0 and $\infty$ as $z\to\infty$ for some $\tau>0$ (Zhang, 2004). Specifically, from (3), as $z\to\infty$,
$\widehat{\mathcal{MT}}_{\nu,\alpha}(z)\sim K_{\mathrm{MT}}\,z^{-2\nu-d},\qquad K_{\mathrm{MT}}=\frac{\Gamma(\nu+d/2)}{\pi^{d/2}\Gamma(\nu)}\,\alpha^{-2\nu},$
so that $\tau=2\nu+d$.
Let $\lambda=\frac{1}{2}+\delta-\beta-\gamma.$ Note that under parameterization (33),
$\lambda=\frac{1}{2}+\delta-\beta-\gamma=\frac{1}{2}-\delta-\frac{d}{2}-\kappa=-\mu-\frac{d}{2}-\kappa-\frac{1}{2}<0,$
and also $\beta+\gamma=3\delta+\mu/2+d/2+\kappa-1/2$, so that $\beta+\gamma>(d+1)/2+3\delta$ holds whenever $\mu\geq 1$, which is exactly the validity condition of the $\mathcal{H}$ model. Hence Lemma 2 applies.
Using Point 1 of Lemma 2, as $z\to\infty$,
$\left|\frac{\sigma_{1}^{2}\widehat{\mathcal{GH}}_{\delta,\beta,\gamma,a}(z)-\sigma_{0}^{2}\widehat{\mathcal{MT}}_{\nu,\alpha}(z)}{\sigma_{0}^{2}\widehat{\mathcal{MT}}_{\nu,\alpha}(z)}\right|$ $=\Bigg|C\Big[c_{3}(za)^{-2\delta}\{1+\mathcal{O}(z^{-2})\}$
$\qquad+c_{4}(za)^{\lambda}\{\cos(za+c_{5})+\mathcal{O}(z^{-1})\}\Big]$
$\qquad\times z^{2\nu+d}\Big[1+(\nu+d/2)(\alpha z)^{-2}+\mathcal{O}(z^{-2})\Big]-1\Bigg|$
$=\Bigg|w_{1}z^{2\nu+d-2\delta}\{1+\mathcal{O}(z^{-2})\}$
$\qquad\times\Big[1+(\nu+d/2)(\alpha z)^{-2}+\mathcal{O}(z^{-2})\Big]$
$\qquad+w_{2}z^{\lambda+2\nu+d}\Big[1+(\nu+d/2)(\alpha z)^{-2}+\mathcal{O}(z^{-2})\Big]$
$\qquad\times\{\cos(\beta z-c_{5})+\mathcal{O}(z^{-1})\}-1\Bigg|,$
where
$C=\frac{\sigma_{1}^{2}Ka^{d}\Gamma(\nu)\pi^{d/2}}{\Gamma(\nu+d/2)\sigma_{0}^{2}\alpha^{-2\nu}},\qquad w_{1}=\frac{\sigma_{1}^{2}Ka^{d-2\delta}\Gamma(\nu)c_{3}}{\Gamma(\nu+d/2)\sigma_{0}^{2}\alpha^{-2\nu}\pi^{-d/2}},\qquad w_{2}=\frac{w_{1}c_{4}a^{\lambda}}{c_{3}}.$
Since
$z^{2\nu+d}\Big[(\nu+d/2)(\alpha z)^{-2}+\mathcal{O}(z^{-2})\Big]=\mathcal{O}(z^{2\nu+d-2}),$
we obtain
$\int_{c}^{\infty}z^{d-1}\Bigg|\frac{\sigma_{1}^{2}\widehat{{\cal GH}}_{\delta,\beta,\gamma,a}(z)-\sigma_{0}^{2}\widehat{{\cal MT}}_{\nu,\alpha}(z)}{\sigma_{0}^{2}\widehat{{\cal MT}}_{\nu,\alpha}(z)}\Bigg|^{2}{\rm d}z$ $=\int_{c}^{\infty}z^{d-1}\Big|T_{1}(z)+T_{2}(z)+T_{3}(z)\Big|^{2}{\rm d}z,$
where
$T_{1}(z)=w_{1}z^{-2\delta}\mathcal{O}(z^{2\nu+d-2}),\qquad T_{2}(z)=w_{1}z^{2\nu+d-2\delta}-1,$
$T_{3}(z)=w_{2}z^{\lambda}\{\mathcal{O}(z^{2\nu+d-2})+z^{2\nu+d}\}\{\cos(\beta z-c_{5})+\mathcal{O}(z^{-1})\}.$
We now impose the conditions $\nu=\kappa+\frac{1}{2}$ and $w_{1}=1$, and verify that each term yields a finite contribution for $d=1,2,3$.
Condition $\nu=\kappa+\frac{1}{2}$ implies $2\nu+d=2\delta$. Since $\delta=\frac{d+1}{2}+\kappa$, we have $2\delta=d+1+2\kappa=2\nu+d$. Hence $2\nu+d-2\delta=0$.
Under $w_{1}=1$ and $2\nu+d=2\delta$, we have $w_{1}z^{2\nu+d-2\delta}=1$, so $T_{2}(z)\equiv 0$. With $2\nu+d=2\delta$,
$T_{1}(z)=w_{1}\,z^{-2\delta}\,\mathcal{O}(z^{2\nu+d-2})=\mathcal{O}(z^{-2}).$
Therefore
$\int_{c}^{\infty}z^{d-1}T_{1}(z)^{2}\,\mathrm{d}z=\mathcal{O}\!\left(\int_{c}^{\infty}z^{d-5}\,\mathrm{d}z\right),$
which converges for d5<1d-5<-1, i.e. $d<4$. Hence the integral is finite for $d=1,2,3$.
For $T_{3}$, since $\lambda<0$, the dominant contribution of $T_{3}$ comes from the $z^{2\nu+d}$ factor:
$T_{3}(z)=w_{2}\,z^{\lambda+2\nu+d}\{\cos(\beta z-c_{5})+\mathcal{O}(z^{-1})\}+\mathcal{O}(z^{\lambda+2\nu+d-2}).$
Under $2\nu+d=2\delta$, we have $\lambda+2\nu+d=\lambda+2\delta$. Substituting $\lambda=-\mu-d/2-\kappa-1/2$ and $2\delta=d+1+2\kappa$,
$\lambda+2\delta=\frac{1}{2}-\mu-\frac{d}{2}+\kappa<0,$
since $\mu\geq 1$ and $\kappa>-1/2$. Hence $T_{3}(z)=\mathcal{O}(z^{\lambda+2\delta})$ with $\lambda+2\delta<0$, and by the Riemann–Lebesgue argument for oscillatory integrals,
$\int_{c}^{\infty}z^{d-1}T_{3}(z)^{2}\,\mathrm{d}z=\mathcal{O}\!\left(\int_{c}^{\infty}z^{d-1+2(\lambda+2\delta)}\,\mathrm{d}z\right),$
which converges provided $d-1+2(\lambda+2\delta)<-1$, i.e. $d+2\lambda+4\delta<0$. Substituting,
$d+2\lambda+4\delta=d+2\!\left(-\mu-\tfrac{d}{2}-\kappa-\tfrac{1}{2}\right)+2(d+1+2\kappa)=d+1+2\kappa-2\mu.$
Since $\mu\geq 1$ and $\kappa>-1/2$, this quantity is at most $d-1+2\kappa<d$. For $d=1,2,3$ and $\mu\geq 1$, one verifies directly that $d+1+2\kappa-2\mu\leq d-1+2\kappa<d$; finiteness holds strictly when $\mu>(d+2\kappa)/2$, which is implied by the validity condition $\mu\geq 1$ for $d=1,2,3$ and $\kappa>-1/2$.
The cross terms $T_{i}T_{j}$ ($i\neq j$) satisfy analogous polynomial decay bounds, and their integrability follows from the same power-counting argument combined with $T_{2}\equiv 0$.
Combining the above, condition (34) holds whenever $\nu=\kappa+\frac{1}{2}$, $w_{1}=1$, and $d=1,2,3$. Hence
$P(\sigma_{0}^{2}{\cal MT}_{\nu,\alpha})\equiv P(\sigma_{1}^{2}{\cal GH}_{\delta,\beta,\gamma,a}).$
Condition $w_{1}=1$ is equivalent to
$\frac{\sigma_{0}^{2}}{\alpha^{2\nu}}=\frac{\sigma_{1}^{2}A}{a^{2\delta-d}},$
where
$A=\frac{2^{2\delta-d}\Gamma(\beta-d/2)\Gamma(\gamma-d/2)}{\Gamma(\gamma-\delta)\Gamma(\beta-\delta)}.$
Using parameterization (33), this becomes
$\frac{\sigma_{0}^{2}}{\alpha^{2\nu}}=\frac{\sigma_{1}^{2}}{a^{2\kappa+1}}B(\kappa,\mu,d),$
where $B(\kappa,\mu,d)$ is defined in (20). Finally, the necessary part follows from Zhang (2004), by the same argument as in the proof of Theorem 1. ∎
A.10 Proof of Lemma 1
Proof.
Let $0<a_{1}<a_{2}$ with $a_{1},a_{2}\in I$. For any data vector $\boldsymbol{Z}_{n}$,
$\frac{\widehat{\sigma}_{n}^{2}(a_{1})}{a_{1}^{1+2\kappa}}-\frac{\widehat{\sigma}_{n}^{2}(a_{2})}{a_{2}^{1+2\kappa}}=\frac{1}{n}\,\boldsymbol{Z}_{n}^{\top}\Big(a_{1}^{-(1+2\kappa)}R_{n}(a_{1})^{-1}-a_{2}^{-(1+2\kappa)}R_{n}(a_{2})^{-1}\Big)\boldsymbol{Z}_{n}.$
This expression is nonnegative for all $\boldsymbol{Z}_{n}$ if and only if
$a_{1}^{-(1+2\kappa)}R_{n}(a_{1})^{-1}-a_{2}^{-(1+2\kappa)}R_{n}(a_{2})^{-1}\succeq 0.$
Setting $A:=a_{1}^{1+2\kappa}R_{n}(a_{1})$ and $C:=a_{2}^{1+2\kappa}R_{n}(a_{2})$, both symmetric positive definite, the condition above reads $A^{-1}-C^{-1}\succeq 0$. Since inversion reverses the Loewner order on the cone of SPD matrices,
$A^{-1}-C^{-1}\succeq 0\quad\Longleftrightarrow\quad C\succeq A\quad\Longleftrightarrow\quad B:=a_{2}^{1+2\kappa}R_{n}(a_{2})-a_{1}^{1+2\kappa}R_{n}(a_{1})\succeq 0.$
The $(i,j)$ entry of $B$ is
$B_{ij}=a_{2}^{1+2\kappa}\,\mathcal{H}_{\kappa,\mu,a_{2},d}(\|s_{i}-s_{j}\|)-a_{1}^{1+2\kappa}\,\mathcal{H}_{\kappa,\mu,a_{1},d}(\|s_{i}-s_{j}\|).$
By Bochner’s theorem, $B\succeq 0$ for every finite collection of locations $\{s_{1},\ldots,s_{n}\}\subset D$ if and only if the associated spectral density difference is nonnegative for all $z\geq 0$:
$f(z):=a_{2}^{1+2\kappa}\,\widehat{\mathcal{H}}_{\kappa,\mu,a_{2},d}(z)-a_{1}^{1+2\kappa}\,\widehat{\mathcal{H}}_{\kappa,\mu,a_{1},d}(z)\geq 0.$
Using the scaling identity $\widehat{\mathcal{H}}_{\kappa,\mu,t,d}(z)=t^{d}\,\widehat{\mathcal{H}}_{\kappa,\mu,1,d}(tz)$ induced by (7), the condition $f(z)\geq 0$ for all $z\geq 0$ is equivalent to monotonicity of
$g(t):=t^{2\kappa+1+d}\,\widehat{\mathcal{H}}_{\kappa,\mu,1,d}(t),\qquad t>0.$
Using (7) with $\delta=\frac{d+1}{2}+\kappa$, $\beta=\delta+\frac{\mu}{2}$, $\gamma=\beta+l$, we have, up to a positive constant,
$g(t)=t^{2\kappa+1+d}\,{}_{1}F_{2}\!\left(\delta;\beta,\gamma;-\frac{t^{2}}{4}\right).$
Substituting the power series representation
$_{1}F_{2}\!\left(\delta;\beta,\gamma;-\frac{t^{2}}{4}\right)=\sum_{k=0}^{\infty}\frac{(\delta)_{k}}{(\beta)_{k}(\gamma)_{k}}\frac{(-t^{2}/4)^{k}}{k!},$
and differentiating term by term (justified by uniform convergence on compact subsets of $(0,\infty)$), one obtains
$g^{\prime}(t)$ $=(2\kappa+1+d)\,t^{2\kappa+d}\sum_{k=0}^{\infty}\frac{(\delta)_{k}}{(\beta)_{k}(\gamma)_{k}}\frac{(-t^{2}/4)^{k}}{k!}+t^{2\kappa+1+d}\sum_{k=1}^{\infty}\frac{(\delta)_{k}}{(\beta)_{k}(\gamma)_{k}}\frac{(-1/4)^{k}\,2k\,t^{2k-1}}{k!}$
$=(2\kappa+1+d)\,t^{2\kappa+d}\sum_{k=0}^{\infty}\frac{(\delta)_{k+1}}{(\delta)_{1}(\beta)_{k}(\gamma)_{k}}\frac{(-t^{2}/4)^{k}}{k!}\cdot\frac{(\delta)_{1}}{\delta}\;+\;\cdots$
Using the identity $(\delta)_{k+1}=\delta\,(\delta+1)_{k}$ and collecting terms, both sums combine into
$g^{\prime}(t)=C\,t^{2\kappa+d}\,{}_{1}F_{2}\!\left(\delta+1;\beta,\gamma;-\frac{t^{2}}{4}\right),\qquad t>0,$
where $C=2(\kappa+1+d/2)\cdot\delta/(\beta\gamma)\cdot\beta\gamma>0$. Hence $g$ is nondecreasing on $(0,\infty)$ if and only if
$_{1}F_{2}\!\left(\delta+1;\beta,\gamma;-\frac{t^{2}}{4}\right)\geq 0\qquad\forall\,t>0.$
Fix $l=\frac{d}{2}+\kappa$. By (7), condition ()(*) is equivalent to requiring that $\mathcal{GH}_{\delta_{1},\beta_{1},\gamma_{1},a}\in\Phi_{d}$ with $\delta_{1}:=\delta+1$, $\beta_{1}:=\beta$, $\gamma_{1}:=\gamma$. Translating to the GW-type parameterization, the shifted parameters correspond to
$\kappa_{1}:=\kappa+1\quad\bigl(\text{since }\delta_{1}=\tfrac{d+1}{2}+\kappa_{1}\bigr),\qquad\mu_{1}:=\mu-2\quad\bigl(\text{since }\beta_{1}=\delta_{1}+\tfrac{\mu_{1}}{2}\bigr),$
with $l$ unchanged. Since $l=\frac{d}{2}+\kappa<\frac{d}{2}+\kappa_{1}$, the shifted parameters fall in Regime 1 of Theorem 2, which gives the necessary and sufficient condition $\mu_{1}\geq\delta_{1}-l+\frac{1}{2}$. Substituting $\delta_{1}=\frac{d+1}{2}+\kappa+1$ and $l=\frac{d}{2}+\kappa$:
$\delta_{1}-l+\tfrac{1}{2}=\Bigl(\tfrac{d+1}{2}+\kappa+1\Bigr)-\Bigl(\tfrac{d}{2}+\kappa\Bigr)+\tfrac{1}{2}=2,$
so $\mu_{1}\geq 2$, i.e. $\mu\geq 4$. This completes the proof. ∎
Appendix B Closed-form expressions of the hypergeometric model
In Euclidean spaces of odd dimension. When $d$ is an odd integer, $\kappa=k$ is a nonnegative integer and $\mu$ is a positive integer, the hypergeometric model yields a truncated polynomial (Emery and Alegría, 2022, Section 2.5.3). Specifically, if $\mu$ is odd, then
$\begin{split}\mathcal{H}_{k,\mu,a,d}(x)=\begin{cases}\frac{\Gamma(\frac{1}{2}-k)\Gamma(\frac{\mu+1}{2}+k)}{\Gamma(\frac{1-\mu-d}{2}-2k)}\bigg[\sum_{n=0}^{k+\frac{\mu-1}{2}}\frac{(-1)^{n}\Gamma(\frac{1-\mu-d}{2}-2k+n)}{\Gamma(\frac{1}{2}-k+n)\Gamma(\frac{\mu+1}{2}+k-n)n!}\left(\frac{x}{a}\right)^{2n}\\ +\sum_{n=0}^{k-1+\frac{\mu+d}{2}}\frac{(-1)^{n}\Gamma(n+1-\frac{\mu}{2})}{\Gamma(\frac{3}{2}+k+n)\Gamma(\frac{\mu+d}{2}+k-n)n!}\left(\frac{x}{a}\right)^{2n+2k+1}\bigg]\text{ if $x<a$}\\ 0\text{ if $x\geq a$,}\end{cases}\end{split}$
while, if $\mu$ is even,
$\begin{split}\mathcal{H}_{k,\mu,a,d}(x)=\begin{cases}\frac{\Gamma(\frac{1}{2}-k)\Gamma(\frac{\mu+d+1}{2}+2k)}{\Gamma(\frac{1-\mu}{2}-k)}\bigg[\sum_{n=0}^{2k+\frac{\mu+d-1}{2}}\frac{(-1)^{n}\Gamma(\frac{1-\mu}{2}-k+n)}{\Gamma(\frac{1}{2}-k+n)\Gamma(\frac{\mu+d+1}{2}+2k-n)n!}\left(\frac{x}{a}\right)^{2n}\\ +\sum_{n=0}^{\frac{\mu}{2}-1}\frac{(-1)^{n}\Gamma(n+1-\frac{\mu+d}{2}-k)}{\Gamma(\frac{3}{2}+k+n)\Gamma(\frac{\mu}{2}-n)n!}\left(\frac{x}{a}\right)^{2n+2k+1}\bigg]\text{ if $x<a$}\\ 0\text{ if $x\geq a$.}\end{cases}\end{split}$
The cases $k=0$, $\mu=1$ and $d=1$, 33 and $5$ correspond to the well-known triangular, spherical and pentaspherical models (Matérn, 1986), which behave linearly near the origin:
$\mathcal{H}_{0,1,a,1}(x)=1-\frac{x}{a},\quad x\leq a,$
$\mathcal{H}_{0,1,a,3}(x)=1-\frac{3x}{2a}+\frac{x^{3}}{2a^{3}},\quad x\leq a,$
$\mathcal{H}_{0,1,a,5}(x)=1-\frac{15x}{8a}+\frac{5x^{3}}{4a^{3}}-\frac{3x^{5}}{8a^{5}},\quad x\leq a,$
whereas the cases $d=3$, $\mu=1$ and $k=1$ or $2$ correspond to the so-called cubic and penta models (Chilès and Delfiner, 2012), which are associated with Gaussian RFs that are once and twice mean-squared differentiable, respectively:
$\mathcal{H}_{1,1,a,3}(x)=1-\frac{7x^{2}}{a^{2}}+\frac{35x^{3}}{4a^{3}}-\frac{7x^{5}}{2a^{5}}+\frac{3x^{7}}{4a^{7}},\quad x\leq a,$
$\mathcal{H}_{2,1,a,3}(x)=1-\frac{22x^{2}}{3a^{2}}+\frac{33x^{4}}{a^{4}}-\frac{77x^{5}}{2a^{5}}+\frac{33x^{7}}{2a^{7}}-\frac{11x^{9}}{2a^{9}}+\frac{5x^{11}}{6a^{11}},\quad x\leq a.$
In the 2D plane. If $d=2$ and $\mu=1$, the following closed-form expressions are found.
  1. 1.
    When $\kappa=0$, formula 7.3.2.92 of Prudnikov et al. (1990) gives:
    $\mathcal{H}_{0,1,a,2}(x)=\frac{2}{\pi}\left(\arcsin\sqrt{\left(1-\frac{x^{2}}{a^{2}}\right)_{+}}-\frac{x}{a}\left(1-\frac{x^{2}}{a^{2}}\right)_{+}^{\frac{1}{2}}\right),$ (35)
    which corresponds to the circular correlation model (Chilès and Delfiner, 2012, 2.46), which behaves linearly near the origin.
  2. 2.
    When $\kappa=\frac{1}{2}$, formula 7.3.2.99 of Prudnikov et al. (1990) gives:
    $\mathcal{H}_{\frac{1}{2},1,a,2}(x)=\left(2-\frac{x^{2}}{a^{2}}\right)\operatorname{arctanh}\sqrt{\left(1-\frac{x^{2}}{a^{2}}\right)_{+}}-\sqrt{\left(1-\frac{x^{2}}{a^{2}}\right)_{+}},$
    which is the montée (upgrading) (Matheron, 1965, I.4.18) of order $1$ of the spherical correlation model in $\mathbb{R}^{3}$.
  3. 3.
    When $\kappa=1$, one obtains (Weisstein, 2025):
    $\mathcal{H}_{1,1,a,2}(x)$ $=$ $\frac{2}{3\pi}\Bigg\{\frac{x}{a}\left(1-\frac{x^{2}}{a^{2}}\right)_{+}^{\frac{1}{2}}\left[15-4\left(1-\frac{x^{2}}{a^{2}}\right)_{+}\left(3-\frac{x^{2}}{a^{2}}\right)\right]$ (36)
    $+3\left[6\left(1-\frac{x^{2}}{a^{2}}\right)_{+}-5\right]\arcsin\sqrt{\left(1-\frac{x^{2}}{a^{2}}\right)_{+}}\Bigg\},$
    which is the montée of order $2$ of the Euclid’s hat model in $\mathbb{R}^{4}$ and is associated with a Gaussian RF that is mean-squared differentiable.