### v3

# Parsimonious Compactly Supported Covariance Models in the Gauss Hypergeometric Class: Identifiability, Reparameterizations, and Asymptotic Properties

Abstract
We study isotropic covariance functions in the Gauss hypergeometric ($\mathcal{GH}$) class, a flexible family that encompasses the celebrated Matérn ($\mathcal{MT}$) and Generalized Wendland ($\mathcal{GW}$) models. We refine validity conditions for positive definiteness in the $\mathcal{GH}$ family and clarify two distinct sources of non-identifiability: a structural one, due to an intrinsic symmetry in the shape parameters, and the classical fixed-domain one, where different parameter values may induce equivalent Gaussian measures and only microergodic combinations are consistently estimable.
To remove the structural ambiguity, we first restrict the parameter space to obtain a unique representation of the correlation function. We then introduce a parsimonious compactly supported subclass selected by a maximum integral range criterion, leading to a new hypergeometric model, denoted $\mathcal{H}$, with simple validity conditions and sparse covariance matrices. The proposed model can be viewed as a canonical refinement of the $\mathcal{GW}$ family and admits compact-support reparameterizations that recover the Matérn model as a limit case.
Under fixed-domain asymptotics, we establish strong consistency and asymptotic normality for likelihood-based estimation of the associated microergodic parameter, including cases where nuisance covariance parameters are estimated over compact sets. Simulation experiments and a climate-data application illustrate finite-sample behavior and show that $\mathcal{H}$ can improve upon $\mathcal{MT}$ while achieving performance comparable to $\mathcal{GW}$ in terms of fit, prediction, and computational efficiency for both Gaussian and Tukey-hh random fields.
Keywords: Fixed domain asymptotics, Gaussian random fields, Generalized Wendland model, Integral range, Sparse matrices.
1 Introduction
Gaussian random fields (RFs) are widely used in spatial statistics, machine learning, numerical analysis, and computer experiments (Cressie, 1993; Stein, 1999; Banerjee et al., 2004; Rasmussen and Williams, 2006; Wendland, 2004; Santner et al., 2003). Under Gaussianity, the dependence structure is fully characterized by the covariance function, whose specification is central for modeling, prediction, and uncertainty quantification.
For RFs on Euclidean spaces, isotropic covariance (kernel) functions— depending only on Euclidean distance—provide a tractable and widely used modeling framework. Among parametric isotropic models, the Matérn ($\mathcal{MT}$) class is arguably the most prominent (Porcu et al., 2024), owing to its smoothness parameter, which continuously controls the mean-square differentiability of the sample paths.
A compactly supported alternative is the generalized Wendland ($\mathcal{GW}$) family (Gneiting, 2002; Bevilacqua et al., 2019), which also allows for continuous control of smoothness while yielding sparse covariance matrices and substantial computational gains (Duff et al., 1989; Davis, 2006) when working with covariance matrices. Connections between $\mathcal{MT}$ and $\mathcal{GW}$ models have been extensively studied (Bevilacqua et al., 2019, 2022).
In this paper, we focus on the Gauss hypergeometric ($\mathcal{GH}$) covariance function introduced by Emery and Alegría (2022). This highly flexible class includes the Matérn ($\mathcal{MT}$) and generalized Wendland ($\mathcal{GW}$) models as special cases and has been recognized as a prominent extension of the Matérn family (Porcu et al., 2024). Although Emery and Alegría (2022) derived convenient sufficient conditions for positive definiteness, a sharper description of the admissible parameter space and a systematic inferential analysis are still lacking. The aim of this paper is to address these issues.
This paper provides several theoretical and methodological contributions. First, we sharpen the validity conditions of the $\mathcal{GH}$ class by deriving explicit constraints that characterize positive definiteness over a substantial portion of the parameter space, thereby improving the results of Emery and Alegría (2022). Second, we clarify that the $\mathcal{GH}$ family is not identifiable in the classical parametric sense: the parameter-to-correlation mapping is not injective due to an intrinsic symmetry in the shape parameters. We distinguish this structural non-identifiability from the well-known inferential limitations under fixed-domain asymptotics, where different parameter values may induce equivalent Gaussian measures and only microergodic parameter combinations are consistently estimable.
To address the structural non-identifiability, we first impose a restriction on the parameter space of the $\mathcal{GH}$ model so that each admissible parameter set corresponds to a unique correlation function.
Once this structural issue is resolved, we pursue a separate goal: identifying a parsimonious and canonical compactly supported subclass within $\mathcal{GH}$. To this end, for fixed smoothness and compact support we adopt an integral-range criterion, selecting the subclass that can attain the largest isotropic integral range, that is, the total mass of the correlation over $\mathbb{R}^{d}$. Since the integral range is proportional to the spectral density at zero frequency, this criterion favors models with the largest low-frequency spectral power, namely the strongest large-scale dependence compatible with the imposed constraints. The integral range thus provides a global and threshold-free measure of dependence. (Vanmarcke, 1983; Lantuéjoul, 2002; Hristopulos and Zukovic, 2011); in dimension $d=1$, it reduces to the area under the correlation function.
The resulting hypergeometric ($\mathcal{H}$) model constitutes the main contribution of this paper. It provides a canonical and parsimonious representative of the $\mathcal{GH}$ class, preserving compact support, continuous smoothness parameterization, and modeling flexibility and can be viewed as a refinement of the ($\mathcal{GW}$) family. Notably, the $\mathcal{H}$ model satisfies simple and explicit necessary and sufficient validity conditions, reducing to a single lower bound on the shape parameter that is independent of the spatial dimension. In addition, it admits a beta-type scale-mixture representation and may be interpreted as a generalization of the Euclid’s hat model (Gneiting, 1999), encompassing the classical spherical covariance functions widely used in geostatistics (Chilès and Delfiner, 2012).
We further show that a compact-support reparameterization of the $\mathcal{H}$ model yields the Matérn model as a special limit case, thereby extending to the hypergeometric framework the Wendland-to-Matérn connection proved in Bevilacqua et al. (2022).
Finally, for a fixed and known smoothness parameter, we establish strong consistency and asymptotic normality for likelihood-based estimation of the microergodic parameter under the proposed $\mathcal{H}$ model, including settings where scale and shape covariance parameters are estimated over compact sets. These results extend, in the hypergeometric framework, the fixed-domain asymptotic theory developed for compactly supported models in Bevilacqua et al. (2019).
A simulation study investigates finite-sample behavior under increasing- and fixed-domain scenarios, and a real-data analysis of climate data illustrates that the proposed $\mathcal{H}$ model can outperform the $\mathcal{MT}$ model and achieve performance comparable to the $\mathcal{GW}$ model, in terms of model fit, prediction accuracy, and computational efficiency, for both Gaussian and Tukey-$h$ RFs (Xu and Genton, 2017).
The $\mathcal{H}$ model is implemented in the GeoModels package (Bevilacqua et al., 2026). The remainder of the paper is organized as follows. Section 2 reviews the $\mathcal{MT}$ and $\mathcal{GW}$ models. Section 3 introduces the $\mathcal{GH}$ model and discusses identifiability. Section 4 provides validity conditions and derives the $\mathcal{H}$ subclass. Section 5 presents closed-form expressions and compact-support reparameterizations. Section 6 establishes asymptotic properties of the microergodic estimator. Sections 7 and 8 report simulation and real-data results. All proofs are collected in Appendix A.
2 Background: the Matérn and Generalized Wendland correlation models
We denote by $\{Z(\boldsymbol{s}):\boldsymbol{s}\in D\}$ a zero-mean Gaussian RF on a bounded set $D\subset\mathbb{R}^{d}$ with stationary isotropic covariance function
${\rm Cov}\!\left(Z(\boldsymbol{s}),Z(\boldsymbol{s}^{\prime})\right)=\sigma^{2}\,C(\boldsymbol{h})=\sigma^{2}\,\phi(\|\boldsymbol{h}\|),\qquad\boldsymbol{h}=\boldsymbol{s}^{\prime}-\boldsymbol{s},$
where $\sigma^{2}>0$ is the variance parameter and $\phi:[0,\infty)\to\mathbb{R}$ is a correlation function satisfying $\phi(0)=1$. We consider the class $\Phi_{d}$ of continuous mappings $\phi$ such that $\phi(\|\cdot\|)$ is positive semidefinite on $\mathbb{R}^{d}$ (equivalently, $\phi\in\Phi_{d}$). In particular,
${\rm corr}\!\left(Z(\boldsymbol{s}),Z(\boldsymbol{s}^{\prime})\right)=C(\boldsymbol{h})=\phi(\|\boldsymbol{h}\|),\qquad\boldsymbol{s},\boldsymbol{s}^{\prime}\in D,$
with $\|\cdot\|$ denoting the Euclidean norm. Gaussian RFs with such covariance functions are called weakly stationary and isotropic.
Schoenberg (1938) characterized the class $\Phi_{d}$ as scale mixtures of the characteristic functions of random vectors uniformly distributed on the spherical shell of $\mathbb{R}^{d}$, with any finite and nonnegative measure, $F$:
$\phi(x)=\int_{0}^{\infty}\Omega_{d}(xr)F(\text{d}r),\qquad x\geq 0,$
with $\Omega_{d}(x)=\Gamma(d/2)(2/x)^{d/2-1}J_{d/2-1}(x)$ and $J_{\nu}$ the Bessel function of the first kind of order $\nu$. The class $\Phi_{d}$ is nested, with the inclusion relation $\Phi_{1}\supset\Phi_{2}\supset\ldots\supset\Phi_{\infty}$ being strict, and where $\Phi_{\infty}:=\bigcap_{d\geq 1}\Phi_{d}$ is the class of continuous mappings $\phi$, the radial version of which is positive semidefinite on any $d$-dimensional Euclidean space.
Fourier transforms of radial versions of members of $\Phi_{d}$, for a given $d$, have a simple expression, as reported in Yaglom (1987) or Stein (1999). For a member $\phi$ of the class $\Phi_{d}$ such that $\phi(\|\cdot\|)$ is absolutely integrable in $\mathbb{R}^{d}$, we define its isotropic spectral density as
$\widehat{\phi}(z)=\frac{z^{1-d/2}}{(2\pi)^{\frac{d}{2}}}\int_{0}^{\infty}u^{d/2}J_{d/2-1}(uz)\phi(u){\rm d}u,\qquad z\geq 0.$ (1)
In the following we review two important examples:
  1. 1.
    The Matérn ($\cal MT$}) family
    The Matérn model is a two-parameter globally supported correlation function defined as (Matérn, 1986):
    ${\cal MT}_{\nu,a}(x)=\frac{2^{1-\nu}}{\Gamma(\nu)}\left(\frac{x}{a}\right)^{\nu}{\cal K}_{\nu}\left(\frac{x}{a}\right),\qquad x>0,$ (2)
    where ${\cal K}_{\nu}$ is the modified Bessel function of the second kind of order $\nu$ (Abramowitz and Stegun, 1970, 9.6.22). The associated spectral density, $\widehat{{\cal MT}}_{a,\nu}$, is given by
    $\widehat{{\cal MT}}_{\nu,a}(z)=\frac{\Gamma(\nu+d/2)}{\pi^{d/2}\Gamma(\nu)}\frac{a^{d}}{(1+a^{2}z^{2})^{\nu+d/2}},\qquad z\geq 0.$ (3)
    Arguments in Stein (1999) show that ${\cal MT}_{\nu,a}$ belongs to $\Phi_{\infty}$ when both the smoothness parameter $\nu$ and the scale parameter $a$ are strictly positive.
    A key feature of the Matérn model is the smoothness parameter $\nu$, which determines the mean-square differentiability of the sample paths of the associated Gaussian RF. In particular, for any integer $m=0,1,\ldots$, the sample paths are $m$-times mean-square differentiable (in any direction) if and only if $\nu>m$. In the special case $\nu=m+\tfrac{1}{2}$, the correlation function reduces to an exponential term multiplied by a polynomial of degree $m$.
  2. 2.
    The Generalized Wendland ($\cal GW$}) family
    The $\cal GW$} model is a three-parameters compactly supported correlation function defined as (Gneiting, 2002; Bevilacqua et al., 2019):
    $\begin{split}{\cal GW}_{\kappa,\mu,a}(x)&=\frac{\Gamma(\kappa)\Gamma(2\kappa+\mu+1)}{\Gamma(2\kappa)\Gamma(\kappa+\mu+1)2^{\mu+1}}\left(1-\frac{x^{2}}{a^{2}}\right)^{\kappa+\mu}_{+}\\ &\times{}_{2}F_{1}\left(\frac{\mu}{2},\frac{\mu+1}{2};\kappa+\mu+1;\left(1-\frac{x^{2}}{a^{2}}\right)_{+}\right),\quad x\geq 0,\end{split}$ (4)
    with $a>0$ a compact support parameter, $\kappa>-\frac{1}{2}$ a smoothness parameter and $\mu$ a shape parameter, and where $(\cdot)_{+}$ stands for the positive part function and $_{2}F_{1}$ for the Gauss hypergeometric function (Abramowitz and Stegun, 1970, 15.1.1). The associated spectral density is given by:
    $\widehat{{\cal GW}}_{\kappa,\mu,a}(z)=La^{d}\mathstrut_{1}F_{2}\Big(\lambda;\lambda+\frac{\mu}{2},\lambda+\frac{\mu}{2}+\frac{1}{2};-\frac{(za)^{2}}{4}\Big),\quad z>0,$ (5)
    where $_{1}F_{2}$ is a generalized hypergeometric function, $\lambda=(d+1)/2+\kappa$ and $L=\frac{2^{-d}\pi^{-\frac{d}{2}}\Gamma(\mu+2\kappa+1)\Gamma(2\kappa+d)\Gamma(\kappa)}{\Gamma\left(\kappa+{d}/{2}\right)\Gamma(\mu+2\kappa+d+1)\Gamma(2\kappa)}$ is a normalization constant.
    For a given $\kappa>-\frac{1}{2}$, the validity conditions for $\mathcal{GW}_{\kappa,\mu,a}\in\Phi_{d}$ follow as a special case of Theorem 2 by setting $l=\frac{1}{2}$:
    • (A)
      $\mathcal{GW}_{\kappa,\mu,a}\in\Phi_{d}$ if and only if $\mu\geq\kappa+\frac{d+1}{2}$, whenever $\kappa\geq 0$ or $d\geq 2$;
    • (B)
      $\mathcal{GW}_{\kappa,\mu,a}\in\Phi_{1}$ if $\mu\geq\frac{\sqrt{8\kappa+9}-1}{2}$, and only if $\mu\geq\kappa+1$, for $\frac{1}{2}<\kappa<0$.
    Since $\kappa+1<\frac{\sqrt{8\kappa+9}-1}{2}$ for all $\frac{1}{2}<\kappa<0$, a gap exists between the necessary and sufficient conditions. This partially corrects Bevilacqua et al. (2024), which claimed in case (B) that $\mu\geq\frac{\sqrt{8\kappa+9}-1}{2}$ is both necessary and sufficient for $d=1$. Whether this lower bound is also necessary remains an open problem.
    The ${\cal GW}_{\kappa,\mu,a}$ model, as in the ${\cal MT}_{\nu,a}$ case, allows parameterizing in a continuous fashion the mean squared differentiability of the underlying Gaussian RF and its associated sample paths through the smoothness parameter $\kappa$. Specifically, for any integer $m=1,2\ldots$, the sample paths of the ${\cal GW}_{\kappa,\mu,a}$ model are $m$ times differentiable, in any direction, if and only if $\kappa>m-0.5$ and for $0\leq\kappa\leq 0.5$ they are not differentiable.
    When $\kappa=k\in\mathbb{N}$, the Generalized Wendland correlation simplifies into the product of an Askey (truncated power) correlation with a polynomial of degree $k$.
    Recently, Bevilacqua et al. (2022) considered a compact support reparameterized version of the $\cal GW$} model, and proved that
    $\lim_{\mu\to\infty}{\cal GW}_{\kappa,\mu,g(a,\mu,\kappa)}(x)={\cal MT}_{\kappa+1/2,a}(x),\quad\kappa>-0.5,$ (6)
    where $g(a,\mu,\kappa)=a\left(\frac{\Gamma(\mu+2\kappa+1)}{\Gamma(\mu)}\right)^{\frac{1}{1+2\kappa}}$, with uniform convergence on $0,\infty)$.
    As a consequence, the reparameterized $\cal GW$} model can be viewed as a generalization of the $\cal MT$} model with an additional parameter that, for given smoothness and spatial dependence parameters, allows switching from the world of flexible compactly supported covariance functions to the world of flexible globally supported covariance functions.
3 The Gauss hypergeometric model and the identifiability problem
The $\mathcal{GH}$ class in Emery and Alegría (2022) has been defined starting from the following spectral density:
$\widehat{{\cal GH}}_{\delta,\beta,\gamma,a}(z)=\frac{\Gamma(\delta)\Gamma(\beta-\frac{d}{2})\Gamma(\gamma-\frac{d}{2})a^{d}}{2^{d}\pi^{\frac{d}{2}}\Gamma\left(\delta-\frac{d}{2}\right)\Gamma(\beta)\Gamma(\gamma)}\mathstrut_{1}F_{2}\Big(\delta;\beta,\gamma;-\frac{(za)^{2}}{4}\Big),\quad z\geq 0.$ (7)
Note that (7) has basically the same form of the spectral density of the $\cal GW$} model in (5). However, in (7) there is no dependence between the parameters of the $\mathstrut{}_{1}F_{2}$ function as in the $\cal GW$} case.
Using some results in Cho and Yun (2018), Emery and Alegría (2022) show that, for a given $a>0$, the following conditions:
$(a)\ \delta>\frac{d}{2},\qquad(b)\ 2(\beta-\delta)(\gamma-\delta)\geq\delta,\qquad(c)\ \beta+\gamma\geq 3\delta+\frac{1}{2}\,$,
are sufficient for the positivity of the spectral density on $0,\infty)$. The associated four-parameter isotropic correlation model is given by:
$\begin{split}\mathcal{GH}_{\delta,\beta,\gamma,a}(x)&=\frac{\Gamma(\beta-\frac{d}{2})\Gamma(\gamma-\frac{d}{2})}{\Gamma(\beta-\delta+\gamma-\frac{d}{2})\Gamma(\delta-\frac{d}{2})}\left(1-\frac{x^{2}}{a^{2}}\right)_{+}^{\beta-\delta+\gamma-\frac{d}{2}-1}\\ &\times{}_{2}F_{1}\left(\beta-\delta;\gamma-\delta;\beta-\delta+\gamma-\frac{d}{2};\left(1-\frac{x^{2}}{a^{2}}\right)_{+}\right),\quad x\geq 0.\end{split}$ (8)
Here $a>0$ is a compact support parameter, $\delta>\frac{d}{2}$ is a smoothness parameter, and $\beta$ and $\gamma$ are two additional shape parameters.
As in the $\cal MT$} and $\cal GW$} cases, the $\cal G$H} model allows parameterizing in a continuous fashion the mean squared differentiability of the underlying Gaussian RF and its associated sample paths through the smoothness parameter $\delta$.
Furthermore, it is not difficult to see that the $\cal G$H} model includes the $\cal GW$} and the $\cal MT$} models as special cases (Emery and Alegría, 2022). For instance:
$\mathcal{GH}_{\delta,\delta+\frac{\mu}{2},\delta+\frac{\mu}{2}+\frac{1}{2},a}(x)={\cal GW}_{\kappa,\mu,a}(x),$ (9)
where $\delta=\frac{d+1}{2}+\kappa$.
Although the $\cal G$H} family is highly flexible, it is not identifiable in the usual parametric sense. Indeed, even under increasing-domain asymptotics—where one would in principle expect all parameters to be estimable (Mardia and Marshall, 1984)—the parameter-to-correlation map is not injective. In particular,
$\mathcal{GH}_{\delta,\beta,\gamma,a}(x)=\mathcal{GH}_{\delta,\gamma,\beta,a}(x),$
because the Gauss hypergeometric function is symmetric in its first two arguments:
$_{2}F_{1}(a,b;\cdot\,;x)={}_{2}F_{1}(b,a;\cdot\,;x).$
Hence distinct parameter values (obtained by swapping $\beta$ and $\gamma$) induce exactly the same correlation function. This implies that additional constraints are required.
A related inferential limitation arises under fixed-domain (infill) asymptotics, where distinct parameter values may induce equivalent Gaussian measures. The following theorem establishes the equivalence condition for two distinct Gaussian measures with $\mathcal{GH}$ correlation models that share a common smoothness parameter. Hereafter, we denote by $P(\rho_{i})$, $i=0,1$, a zero-mean Gaussian measure with an isotropic covariance function $\rho_{i}$.
Theorem 1.
Let $P(\sigma_{i}^{2}{\cal GH}_{\delta,\beta_{i},\gamma_{i},a_{i}})$, $i=0,1$, be two zero-mean Gaussian measures and let $\delta>\frac{d}{2}$. If $\frac{d+1}{2}+3\delta<\min(\beta_{1}+\gamma_{1},\beta_{0}+\gamma_{0}),$ and $2(\beta_{i}-\delta)(\gamma_{i}-\delta)\geq\delta.$ for any bounded infinite set $D\subset\mathbb{R}^{d}$, $d=1,2,3$, the Gaussian measures $P(\sigma_{1}^{2}{\cal GH}_{\delta,\beta_{1},\gamma_{1},a_{1}})$ and $P(\sigma_{0}^{2}{\cal GH}_{\delta,\beta_{0},\gamma_{0},a_{0}})$ are equivalent on the paths of $\{Z(\boldsymbol{s}):\boldsymbol{s}\in D\}$, if and only if
$\frac{\sigma_{0}^{2}L(\beta_{0},\gamma_{0})}{a_{0}^{2\delta-d}}=\frac{\sigma_{1}^{2}L(\beta_{1},\gamma_{1})}{a_{1}^{2\delta-d}},$ (10)
where $L(\beta_{i},\gamma_{i})=\dfrac{2^{2\delta-d}\Gamma(\beta_{i}-d/2)\Gamma(\gamma_{i}-d/2)}{\Gamma(\gamma_{i}-\delta)\Gamma(\beta_{i}-\delta)}$, $i=0,1$.
The restriction to $d=1,2,3$ is typical when studying equivalence of Gaussian measures for isotropic models on $\mathbb{R}^{d}$, and is not specific to the $\cal G$H} class. Indeed, the same dimensional restriction appears in the equivalence results for the $\cal MT$} model (Zhang, 2004), and stems from the integrability conditions required by the criteria of Stein (2004).
An immediate consequence of Theorem 1 is that, for fixed $\delta$, the microergodic parameter is given by
$\frac{\sigma^{2}L(\beta,\gamma)}{a^{2\delta-d}}.$
Because $L(\beta,\gamma)=L(\gamma,\beta)$, this quantity does not distinguish $(\beta,\gamma)$ from $(\gamma,\beta)$. Thus, while the microergodic parameter is the relevant identifiable feature under fixed-domain asymptotics, the individual shape parameters $\beta$ and $\gamma$ are not identifiable without an additional constraint.
Summarizing, the $\mathcal{GH}$ correlation model is highly flexible and includes the $\mathcal{GW}$ and $\mathcal{MT}$ models as special cases. However, it is not identifiable in the usual parametric sense because swapping $(\beta,\gamma)$ leaves the correlation unchanged. This issue can be observed under both increasing and fixed domain scenario.
4 Fixing the identifiability problem: the Hypergeometric correlation model
In this section, we first refine the validity conditions of the $\mathcal{GH}$ model, clarifying the relationship between necessary and sufficient constraints. We then address the non-uniqueness discussed in the previous section by imposing an appropriate restriction on the parameter space, which restores a unique representation of the correlation function. Finally, we introduce a parsimonious compactly supported subclass selected through an explicit optimality criterion.
To this end, we adopt a $\cal GW$}-type reparameterization of the $\mathcal{GH}_{\delta,\beta,\gamma,a}$ model and set
$\delta=\frac{d+1}{2}+\kappa,\qquad\beta=\delta+\frac{\mu}{2},\qquad\gamma=\beta+l,\qquad l\geq 0.$ (11)
As noted in (9), this parametrization includes the $\cal GW$} model as the special case $l=\frac{1}{2}$.
The restriction $l\geq 0$ in (11) is equivalent to imposing the ordering constraint $\gamma\geq\beta$ on the original $\mathcal{GH}_{\delta,\beta,\gamma,a}$ family. This removes the intrinsic label-switching ambiguity induced by the symmetry of $_{2}F_{1}$ in its upper parameters, namely $\mathcal{GH}_{\delta,\beta,\gamma,a}=\mathcal{GH}_{\delta,\gamma,\beta,a}$, and thus yields a unique representation up to this exchange.
The following theorem improves Theorem 1 in Emery and Alegría (2022) by providing sharp validity conditions for the reparameterized model $\mathcal{GH}_{\delta,\delta+\frac{\mu}{2},\delta+\frac{\mu}{2}+l,a}$.
Theorem 2.
Let $\kappa>-\frac{1}{2}$, $\delta=\frac{d+1}{2}+\kappa$, $a>0$. Then:
  1. 1.
    If $0\leq l\leq\frac{d}{2}+\kappa$, then
    $\mathcal{GH}_{\delta,\delta+\frac{\mu}{2},\delta+\frac{\mu}{2}+l,a}\in\Phi_{d}\quad\text{if and only if}\quad\mu\geq\delta-l+\frac{1}{2}.$
  2. 2.
    If $l>\frac{d}{2}+\kappa$, then
    $\mathcal{GH}_{\delta,\delta+\frac{\mu}{2},\delta+\frac{\mu}{2}+l,a}\in\Phi_{d}\quad\text{if}\quad\mu\geq\sqrt{2\kappa+l^{2}+d+1}-l,$
    and only if $\mu\geq\delta-l+\frac{1}{2}.$
Remark 1.
Tracing back the conditions of Theorem 2 to the original parametrization $(\delta,\beta,\gamma,a)$, and assuming $\delta>d/2$ and $\gamma\geq\beta$ we obtain:
  1. 1.
    If $0\leq\gamma-\beta\leq\delta-\frac{1}{2}$, then
    $\mathcal{GH}_{\delta,\beta,\gamma,a}\in\Phi_{d}\quad\text{if and only if}\quad\beta+\gamma\geq 3\delta+\frac{1}{2}.$
  2. 2.
    If $\gamma-\beta>\delta-\frac{1}{2}$, then
    $\mathcal{GH}_{\delta,\beta,\gamma,a}\in\Phi_{d}\quad\text{if}\quad\beta+\gamma\geq 2\delta+\sqrt{(\gamma-\beta)^{2}+2\delta},$
    and only if $\beta+\gamma\geq 3\delta+\frac{1}{2}.$
In particular, on the broad regime $\gamma-\beta\leq\delta-\tfrac{1}{2}$ the sufficient conditions (a)–(c) in Section 3 reduce to the single inequality $\beta+\gamma\geq 3\delta+\tfrac{1}{2}$ and become sharp, since this inequality is also necessary. This sharpens the earlier validity region reported in Emery and Alegría (2022). Outside this regime, the sufficient bound above is explicit while a residual gap with the necessary condition remains.
For $l=\frac{1}{2}$, the necessary and sufficient validity conditions for $\mathcal{GW}_{\kappa,\mu,a}\in\Phi_{d}$ discussed in Section 2 are recovered. Indeed, if $l=\frac{1}{2}\leq\frac{d}{2}+\kappa$ (which holds for all $d\geq 2$ and $\kappa>-\frac{1}{2}$, and for $d=1$ with $\kappa\geq 0$), then Regime 1 of Theorem 2 yields $\mu\geq\delta=\frac{d+1}{2}+\kappa$. For $d=1$ and $\frac{1}{2}<\kappa<0$, instead $l=\frac{1}{2}>\frac{d}{2}+\kappa$, so Regime 2 applies, yielding the sufficient condition $\mu\geq\frac{1}{2}(\sqrt{8\kappa+9}-1)$ and the necessary condition $\mu\geq\kappa+1$, consistently with the discussion in Section 2.
Theorem 2 characterizes validity for each admissible value of $l$. However, even after restricting to $l\geq 0$ the reparameterized model still involves two shape parameters, $\mu$ and $l$. In the spirit of the $\cal GW$} model, we therefore seek a parsimonious specification by treating $l$ as fixed (or by linking it to the remaining parameters and/or to the spatial dimension), so that inference targets a single interpretable shape parameter $\mu$. This naturally raises the question: which choice of $l\geq 0$ is optimal? We address this question by introducing an explicit optimality criterion that selects a parsimonious subclass once smoothness and compact support are fixed. Specifically, we choose $l$ so that the resulting $\mathcal{GH}$ subclass can attain the largest isotropic integral range.
For an isotropic correlation $\phi\in\Phi_{d}$, the integral range is
$A=\int_{\mathbb{R}^{d}}\phi(\|\boldsymbol{h}\|)\,\mathrm{d}\boldsymbol{h}=\frac{2\pi^{\frac{d}{2}}}{\Gamma(\frac{d}{2})}\int_{0}^{\infty}x^{d-1}\phi(x)\,\mathrm{d}x.$ (12)
This quantity, originally introduced by Lantuéjoul (1991), provides a global and threshold-free measure of dependence (Yaglom, 1987; Matheron, 1989; Hristopulos and Zukovic, 2011).
If $\phi(\|\cdot\|)$ is absolutely integrable in $\mathbb{R}^{d}$ (in particular, if it is compactly supported), then
$A=(2\pi)^{d}\,\widehat{\phi}(0),$
so the integral range is proportional to the spectral density at zero frequency.
For compactly supported correlations, maximizing $A$ therefore targets models with maximal low-frequency spectral power, that is, the strongest large-scale dependence compatible with the imposed constraints. While alternative optimality principles could be considered, they would typically depend on the sampling design or on a specific loss function. In contrast, the integral range criterion is intrinsic to the model class and yields a reproducible, one-to-one specification (through $l$) within the family.
For a parametric model $\phi_{\boldsymbol{\theta}}\in\Phi_{d}$, we denote the associated integral range by $A_{\boldsymbol{\theta},d}$. For example, for the Matérn correlation one has
$A_{\nu,a,d}=2^{d}\pi^{d/2}a^{d}\,\frac{\Gamma\!\left(\nu+\frac{d}{2}\right)}{\Gamma(\nu)}.$ (13)
The following theorem provides the integral range of $\mathcal{GH}_{\delta,\delta+\frac{\mu}{2},\delta+\frac{\mu}{2}+l,a}$.
Theorem 3.
The integral range of $\mathcal{GH}_{\delta,\delta+\frac{\mu}{2},\delta+\frac{\mu}{2}+l,a}$ is given by
$A_{\kappa,\mu,l,a,d}=a^{d}\left[\frac{\pi^{d/2}\Gamma(\frac{d+1}{2}+\kappa)\,\Gamma\left(\frac{\mu+1}{2}+\kappa\right)\,\Gamma\left(\frac{\mu+1}{2}+\kappa+l\right)}{\Gamma(\kappa+\frac{1}{2})\,\Gamma\left(\frac{\mu+1+d}{2}+\kappa\right)\,\Gamma\left(\frac{\mu+1+d}{2}+\kappa+l\right)}\right].$ (14)
Note that the compact support $a$ raised to power $d$ represents a scale factor in the integral range computation. In addition, using asymptotic properties of the Gamma function, it can be shown that $A_{\kappa,\mu,l,a,d}\sim\mu^{-d}$ as $\mu\to\infty$ and, as a consequence,
$\lim_{\mu\to\infty}A_{\kappa,\mu,l,a,d}=0.$ (15)
This implies that, as the shape parameter $\mu$ increases, the integral range approaches zero, regardless of the compact support and smoothness parameters.
The following theorem establishes that, for a given $l\geq 0$, the integral range is maximized by taking $\mu$ at the lower bound of its admissible domain. In what follows, we call admissible the parameter region ensured by Theorem 2. In Regime 2 of Theorem 2, validity is guaranteed only through the explicit sufficient bound. Throughout Sections 45, whenever we optimize over $(\mu,l)$ we therefore interpret the admissible region as the parameter set ensured by the sufficient conditions in Theorem 2.
Theorem 4.
Let $\kappa>-\frac{1}{2}$, $a>0$, and $d\in\mathbb{N}$ be fixed, and set $\delta=\frac{d+1}{2}+\kappa$. For each $l\geq 0$, the map $\mu\mapsto A_{\kappa,\mu,l,a,d}$ is non-increasing on its admissible domain. Hence its maximum is attained at the smallest admissible value of $\mu$. In particular,
  1. 1.
    if $0\leq l\leq d/2+\kappa$, then the maximum is attained at μ^(l)=μ_1(l):=δ-l+12;
  2. 2.
    if $l>d/2+\kappa$, then the maximum is attained at μ^(l)=μ_2(l):=2κ+l^2+d+1-l.
For instance, if $l=\frac{1}{2}$ and $d\geq 2$ (or $d=1$ and $\kappa\geq 0$), then by (9)
$\mathcal{GH}_{\delta,\delta+\frac{\mu}{2},\delta+\frac{\mu+1}{2},a}={\cal GW}_{\kappa,\mu,a},$
and the maximum integral range (over the admissible $\mu$) is attained at $\mu=\mu_{1}(\frac{1}{2})=\delta=\frac{d+1}{2}+\kappa$, which coincides with the lower bound of the $\cal GW$} model in this case.
As an example, consider the case $d=1$ and $\kappa=0,1$. Figure 1 (left) depicts ${\cal GW}_{0,\mu,1}$ and Figure 1 (right) depicts ${\cal GW}_{1,\mu,1}$ for $\mu=\mu_{1}(\frac{1}{2}),3,10,100$. It is clear that, in both panels, the correlation with the largest area under the curve is obtained for $\mu=\mu_{1}\!\left(\tfrac{1}{2}\right)$ (red line).
The next step is to determine which $l\geq 0$ yields the largest attainable integral range, i.e., which $l$ maximizes $A_{\kappa,\mu^{\star}(l),l,a,d}$.
Theorem 5.
Let $\kappa>-1/2$, $a>0$, and $d\in\mathbb{N}$ be fixed.
  1. 1.
    The function $l\mapsto A_{\kappa,\mu_{1}(l),l,a,d}$ is non-decreasing on $0,d/2+\kappa]$, hence
    $\max_{0\leq l\leq d/2+\kappa}A_{\kappa,\mu_{1}(l),l,a,d}=A_{\kappa,1,d/2+\kappa,a,d}.$
  2. 2.
    The function $l\mapsto A_{\kappa,\mu_{2}(l),l,a,d}$ is non-increasing on $(d/2+\kappa,\infty)$ and
    $\lim_{l\downarrow d/2+\kappa}A_{\kappa,\mu_{2}(l),l,a,d}=A_{\kappa,1,d/2+\kappa,a,d}.$
In particular, the largest attainable integral range over all admissible $(\mu,l)$ is achieved at $(\mu,l)=(1,d/2+\kappa)$ and equals $A_{\kappa,1,d/2+\kappa,a,d}$.
Refer to caption Refer to caption
Figure 1: Left part: ${\cal GW}_{0,\mu,1}$ for $\mu=\mu_{1}(\frac{1}{2}),3,10,100$ with $\mu_{1}(\frac{1}{2})=\delta=\frac{d+1}{2}+\kappa$ and $\kappa=0$. Right part: the same Figure but with $\kappa=1$. In both cases $d=1$.
As an example, Figure 2 compares $\mathcal{GH}_{\delta,\delta+\frac{\mu_{1}(l)}{2},\delta+\frac{\mu_{1}(l)}{2}+l,1}$ when $l=0,\frac{1}{2},d/2+\kappa$ (black, red and blue lines, respectively). In particular the left part depicts the case $\kappa=0$ and the right part depicts the case $\kappa=1$ ($d=2$ in both cases). It is apparent that the case $l=d/2+\kappa$ (blue line) corresponds to the model with higher integral range and improves the maximum integral range attained by the $\cal GW$} models (red line).
Refer to caption Refer to caption
Figure 2: Left part: $\mathcal{GH}_{\delta,\delta+\frac{\mu(l)}{2},\delta+\frac{\mu(l)}{2}+l,1}$ with $\delta=\frac{d+1}{2}+\kappa$, $\kappa=0$, $\mu(l)=\delta-l+\frac{1}{2}$ and $l=0,\frac{1}{2},d/2+\kappa$ (black, red and blue lines respectively). Right part: the same Figure but with $\kappa=1$. The case $l=\frac{1}{2}$ (red line) corresponds to the $\mathcal{GW}_{\kappa,\mu,1}$ model. The blue line corresponds to the correlation model that maximizes the integral range. In both cases $d=2$.
Summarizing, Theorem 5 shows that the largest attainable integral range over all admissible $(\mu,l)$ is achieved at $(\mu,l)=(1,d/2+\kappa).$ In particular, fixing $l=d/2+\kappa$ identifies the unique $\mathcal{GH}$ specification that can attain this maximum (with the maximum attained at $\mu=1$).
Accounting for (11), the extremal member attaining the maximal integral range (i.e., $l=d/2+\kappa$ and $\mu=1$) is $\mathcal{GH}_{\delta,\delta+\frac{1}{2},2\delta,a}$. This coincides with the Euclid’s hat correlation model (Gneiting, 1999) when $\kappa=0$ (triangular/tent for $d=1$, circular for $d=2$, spherical for $d=3$, pentaspherical for $d=5$, see Section 5), and with an upgraded Euclid’s hat when $\kappa\neq 0$ (Emery and Alegría, 2022), obtained via a (fractional) montée (Matheron, 1965, I.4.18).
Motivated by this, we fix $l=d/2+\kappa$ and define the hypergeometric correlation by
$\mathcal{H}_{\kappa,\mu,a,d}(x):=\mathcal{GH}_{\delta,\delta+\frac{\mu}{2},\delta+\frac{\mu+d}{2}+\kappa,a}(x),\qquad x\geq 0.$ (16)
An explicit expression is given by
$\begin{split}\mathcal{H}_{\kappa,\mu,a,d}(x)&=\frac{\Gamma\!\left(\kappa+\frac{\mu+1}{2}\right)\Gamma\!\left(2\kappa+\frac{d+\mu+1}{2}\right)}{\Gamma\!\left(\mu+\frac{d+1}{2}+2\kappa\right)\Gamma\!\left(\kappa+\frac{1}{2}\right)}\left(1-\frac{x^{2}}{a^{2}}\right)_{+}^{\mu+\frac{d-1}{2}+2\kappa}\\[4.0pt] &\quad\times{}_{2}F_{1}\!\left(\frac{\mu}{2},\frac{\mu+d}{2}+\kappa;\mu+\frac{d+1}{2}+2\kappa;\left(1-\frac{x^{2}}{a^{2}}\right)_{+}\right),\qquad x\geq 0.\end{split}$ (17)
Notably, the model $\mathcal{H}_{\kappa,\mu,a,d}$ always lies in Regime 1 of Theorem 2, where validity is completely characterized. Indeed, by setting $l=d/2+\kappa$ in Theorem 2, we obtain
$\mathcal{H}_{\kappa,\mu,a,d}\in\Phi_{d}\quad\text{if and only if}\quad\mu\geq 1.$
Refer to caption Refer to caption
Figure 3: Two Gaussian RF realizations with ${\cal H}_{\kappa,4,0.2,2}$ correlation model, for $\kappa=0,1$ (from left to right).
These validity conditions are substantially simpler than those for the $\mathcal{GW}_{\kappa,\mu,a}$ model discussed in Section 2. In particular, the lower bound on $\mu$ is independent of the spatial dimension. The trade-off is that, unlike the $\mathcal{GW}_{\kappa,\mu,a}$ model, the expression of $\mathcal{H}_{\kappa,\mu,a,d}$ explicitly depends on $d$.
The next theorem shows that the extremal subclass ${\cal H}_{\kappa,1,a,d}$, which attains the maximal integral range, retains substantial modeling flexibility. Through beta-type scale mixtures, it can generate any member of ${\cal H}_{\kappa,\mu,a,d}$ and a broad subclass of the Gauss–hypergeometric family $\mathcal{GH}_{\delta,\delta+\frac{\mu}{2},\delta+\frac{\mu}{2}+l,a}$. Thus, the maximal integral range criterion does not restrict modeling capacity, but instead selects a parsimonious representative that remains rich within the admissible class.
Theorem 6.
Let $a>0$, $\kappa>-\frac{1}{2}$, $\delta=\frac{d+1}{2}+\kappa$, $\mu>1$ and $l\geq 0$. Assume that the parameters $(\mu,l)$ belong to the admissible region of Theorem 2, so that $\mathcal{GH}_{\delta,\delta+\frac{\mu}{2},\delta+\frac{\mu}{2}+l,a}\in\Phi_{d}$. In addition, assume $\mu+2l>1+d+2\kappa$. Then $\mathcal{GH}_{\delta,\delta+\frac{\mu}{2},\delta+\frac{\mu}{2}+l,a}$ admits a beta-type scale mixture representation in terms of $\mathcal{H}_{\kappa,1,b,d}$ with $b=a\sqrt{uv}$:
$\begin{split}\mathcal{GH}_{\delta,\delta+\frac{\mu}{2},\delta+\frac{\mu}{2}+l,a}(x)&=\frac{\Gamma\!\left(\delta+\frac{\mu-d}{2}\right)\Gamma\!\left(\delta+\frac{\mu-d}{2}+l\right)}{\Gamma\!\left(\delta+\frac{1-d}{2}\right)\Gamma\!\left(\delta+\frac{1}{2}+\kappa\right)\Gamma\!\left(\frac{\mu-1}{2}\right)\Gamma\!\left(\frac{\mu-1-d}{2}-\kappa+l\right)}\\ &\quad\times\int_{0}^{1}\!\!\int_{0}^{1}u^{\delta-\frac{d+1}{2}}(1-u)^{\frac{\mu-3}{2}}v^{\delta-\frac{1}{2}+\kappa}(1-v)^{\frac{\mu-d-3}{2}-\kappa+l}\\ &\qquad\qquad\times\mathcal{H}_{\kappa,1,a\sqrt{uv},d}(x)\,\mathrm{d}u\,\mathrm{d}v,\qquad x\geq 0.\end{split}$ (18)
In particular, when $l=d/2+\kappa$, one obtains
$\begin{split}\mathcal{H}_{\kappa,\mu,a,d}(x)&=\frac{\Gamma\!\left(\delta+\frac{\mu-d}{2}\right)\Gamma\!\left(\delta+\frac{\mu}{2}+\kappa\right)}{\Gamma\!\left(\delta+\frac{1-d}{2}\right)\Gamma\!\left(\delta+\frac{1}{2}+\kappa\right)\Gamma^{2}\!\left(\frac{\mu-1}{2}\right)}\\ &\quad\times\int_{0}^{1}\!\!\int_{0}^{1}u^{\delta-\frac{d+1}{2}}(1-u)^{\frac{\mu-3}{2}}v^{\delta-\frac{1}{2}+\kappa}(1-v)^{\frac{\mu-3}{2}}\\ &\qquad\qquad\times\mathcal{H}_{\kappa,1,a\sqrt{uv},d}(x)\,\mathrm{d}u\,\mathrm{d}v,\qquad x\geq 0.\end{split}$
[…]
5 Some special cases and bridges to the Matérn model
The computation of the correlation function $\mathcal{H}_{\kappa,\mu,a,d}$ requires the evaluation of the Gauss hypergeometric function. Although efficient numerical implementations are available (Johansson, 2019), this may be less familiar to spatial statisticians and machine learning practitioners accustomed to closed-form parametric correlation models.
As in the $\mathcal{MT}$ and $\mathcal{GW}$ families, substantial simplifications arise in certain special cases. The following theorem considers the case $\kappa=k\in\mathbb{N}_{0}$, $d=1$, and $\mu\geq 1$.
Theorem 7.
If $\kappa=k$ is a nonnegative integer, $d=1$, and $\mu\geq 1$, then, for $x\geq 0$,
$\mathcal{H}_{k,\mu,a,1}(x)=\begin{cases}\displaystyle\frac{\sqrt{\pi}\,\Gamma\!\left(\frac{\mu}{2}+2k+1\right)}{\Gamma\!\left(k+\tfrac{1}{2}\right)}\left(1-\frac{x}{a}\right)^{\mu+2k}\left(\frac{x}{a}\right)^{k}\\[6.00006pt] \displaystyle\qquad\times\sum_{n=0}^{k}\frac{(n+k)!}{2^{2n}n!\,\Gamma\!\left(n+\frac{\mu}{2}+k+1\right)(k-n)!}\left(\frac{\left(1-\frac{x}{a}\right)^{2}}{\frac{x}{a}}\right)^{n},&0\leq x<a,\\[11.00008pt] 0,&x\geq a.\end{cases}$
General closed forms for arbitrary $d>1$, $\mu\geq 1$, and integer $\kappa$ are more difficult to obtain. Nevertheless, additional simplifications can be derived using known identities of the Gauss hypergeometric function. For instance, when $d$ is odd, $\mu\geq 1$, and $\kappa=0$, the following closed-form expression holds.
Theorem 8.
Let $d$ be an odd integer and $\mu\geq 1$. If $\kappa=0$, then
$\mathcal{H}_{0,\mu,a,d}(x)=C\left(1-\frac{x}{a}\right)_{+}^{\mu+\frac{d-1}{2}}\sum_{n=0}^{\frac{d-1}{2}}\frac{\left(\frac{1-d}{2}\right)_{n}\left(\frac{d+1}{2}\right)_{n}\left(1-\frac{x}{a}\right)_{+}^{n}}{2^{n}n!\,\Gamma\!\left(n+\frac{d-1}{2}+\mu+1\right)},$
where $C=\frac{2^{\mu+\frac{d-1}{2}}\Gamma\!\left(\frac{\mu+1}{2}\right)\Gamma\!\left(\frac{\mu+d+1}{2}\right)}{\sqrt{\pi}}.$
Additional closed-form expressions are reported in Appendix B, including the spherical models (Chilès and Delfiner, 2012) that are widely used in geostatistical practice and that are recovered as special cases of the proposed hypergeometric family.
The Matérn model can be recovered as a limiting case of $\mathcal{H}_{\kappa,\mu,a,d}$ under an appropriate compact-support reparameterization. To establish this result, we first characterize the equivalence between the Gaussian measures $P(\sigma_{0}^{2}\mathcal{MT}_{\nu,\alpha})$ and $P(\sigma_{1}^{2}\mathcal{H}_{\kappa,\mu,a,d})$. The proof is provided in Appendix A.
Theorem 9.
Let $P(\sigma_{0}^{2}{\cal MT}_{\nu,\alpha})$ and $P(\sigma_{1}^{2}{\cal H}_{\kappa,\mu,a,d})$ be two zero-mean Gaussian measures, and let $\kappa>-\frac{1}{2}$ , $\nu=\kappa+\frac{1}{2}$ and $\mu>1+d/2$ . For any bounded infinite set $D\subset\mathbb{R}^{d}$, $d=1,2,3$, the Gaussian measures $P(\sigma_{0}^{2}{\cal MT}_{\nu,\alpha})$ and $P(\sigma_{1}^{2}{\cal H}_{\kappa,\mu,a,d})$ are equivalent on the paths of $\{Z(\boldsymbol{s}):\boldsymbol{s}\in D\}$, if and only if
$\frac{\sigma_{0}^{2}}{\alpha^{2\nu}}=\frac{\sigma_{1}^{2}}{a^{2\kappa+1}}\,F(\kappa,\mu,d),$ (19)
where
$F(\kappa,\mu,d)=\frac{2^{2\kappa+1}\Gamma\!\left(\frac{\mu+1}{2}+\kappa\right)\Gamma\!\left(\frac{\mu+d+1}{2}+2\kappa\right)}{\Gamma\!\left(\frac{\mu}{2}\right)\Gamma\!\left(\frac{\mu+d}{2}+\kappa\right)}.$ (20)
Assuming a common variance and imposing the relation $\nu=\kappa+\frac{1}{2}$, the equivalence condition (19) is equivalently expressed as
$a=\alpha\,F(\kappa,\mu,d)^{\frac{1}{1+2\kappa}}:=b(\alpha,\mu,\kappa,d).$
Using this compact-support reparameterization and following the same arguments in Bevilacqua et al. (2022), one obtains
$\lim_{\mu\to\infty}{\cal H}_{\kappa,\mu,b(\alpha,\mu,\kappa,d),d}(x)={\cal MT}_{\kappa+\frac{1}{2},\alpha}(x),\quad\kappa>-\frac{1}{2},$ (21)
with uniform convergence on $0,\infty)$.
[…]
It is worth emphasizing that, under this compact-support reparameterization, the role of the shape parameter $\mu$ changes substantially. For fixed smoothness $\kappa$ and scale parameter $\alpha$, $\mu$ determines the compact support.
As $\mu\to\infty$, ${\cal H}_{\kappa,\mu,b(\alpha,\mu,\kappa,d),d}$ converges to the Matérn model ${\cal MT}_{\kappa+\frac{1}{2},\alpha}$, with maximal integral range given by (13).
6 Asymptotic properties of the maximum likelihood estimator of the Hypergeometric model
We study the asymptotic properties of likelihood-based estimators under fixed-domain (infill) asymptotics for the model $\sigma^{2}\mathcal{H}_{\kappa,\mu,a,d}$, with $\kappa$ fixed and known.
Using Theorem 1 and the parameterization (11) with $l=d/2+\kappa$, the equivalence condition for two Gaussian measures $P(\sigma_{i}^{2}\mathcal{H}_{\kappa,\mu_{i},a_{i},d})$, $i=0,1$, sharing the same $\kappa$, for $d=1,2,3$ and $\min(\mu_{0},\mu_{1})>1+d/2$, is
$\frac{\sigma_{0}^{2}\,F(\kappa,\mu_{0},d)}{a_{0}^{2\kappa+1}}=\frac{\sigma_{1}^{2}\,F(\kappa,\mu_{1},d)}{a_{1}^{2\kappa+1}}.$ (22)
Hence, under fixed-domain (infill) asymptotics, the individual parameters $\sigma^{2}$, $\mu$ and $a$ are not consistently identifiable; only the microergodic parameter
$\theta(\sigma^{2},\mu,a):=\frac{\sigma^{2}\,F(\kappa,\mu,d)}{a^{2\kappa+1}}$ (23)
is consistently estimable. In our fixed-domain limit theory we assume $\mu>1+d/2$ for $d=1,2,3$, which is the condition ensuring equivalence of Gaussian measures (see Theorem 1). This requirement is stronger than the mere validity condition for the $\mathcal{H}$ model, which is simply $\mu\geq 1$.
Let $D\subset\mathbb{R}^{d}$ be a bounded domain and let $S_{n}=\{\mathbf{s}_{1},\ldots,\mathbf{s}_{n}\}\subset D$ be an infill design, i.e. $S_{n}$ becomes dense in $D$ as $n\to\infty$. Let $\mathbf{Z}_{n}=(Z(\mathbf{s}_{1}),\ldots,Z(\mathbf{s}_{n}))^{\top}$ be a realization of a zero-mean stationary Gaussian RF with covariance function $\sigma^{2}\mathcal{H}_{\kappa,\mu,a,d}(\|\cdot\|)$. Write
$R_{n}(\mu,a):=\bigl[\mathcal{H}_{\kappa,\mu,a,d}(\|\mathbf{s}_{i}-\mathbf{s}_{j}\|)\bigr]_{i,j=1}^{n}$
for the $n\times n$ correlation matrix, and
$\ell_{n}(\sigma^{2},\mu,a):=-\frac{1}{2}\Bigl(n\log(2\pi\sigma^{2})+\log\lvert R_{n}(\mu,a)\rvert+\frac{1}{\sigma^{2}}\mathbf{Z}_{n}^{\top}R_{n}(\mu,a)^{-1}\mathbf{Z}_{n}\Bigr)$ (24)
for the Gaussian log-likelihood. Throughout this section $\kappa$ is fixed and known, while $\mu$ and $a$ may be either fixed or estimated as specified below.
For any $(\mu,a)$, the profile ML estimator of $\sigma^{2}$ is
$\hat{\sigma}_{n}^{2}(\mu,a):=\frac{\mathbf{Z}_{n}^{\top}R_{n}(\mu,a)^{-1}\mathbf{Z}_{n}}{n},$ (25)
and the corresponding plug-in estimator of the microergodic parameter (23) is
$\hat{\theta}_{n}(\mu,a):=\theta\!\bigl(\hat{\sigma}_{n}^{2}(\mu,a),\mu,a\bigr)=\frac{\hat{\sigma}_{n}^{2}(\mu,a)\,F(\kappa,\mu,d)}{a^{2\kappa+1}}.$ (26)
The profile log-likelihood obtained by substituting (25) into (24) is
$\mathcal{PL}_{n}(\mu,a):=-\frac{1}{2}\Bigl(n\log\bigl(2\pi\hat{\sigma}_{n}^{2}(\mu,a)\bigr)+\log\lvert R_{n}(\mu,a)\rvert+n\Bigr).$ (27)
When $\mu$ is fixed and known, the factor $F(\kappa,\mu,d)$ in (23) is a known constant. Thus, the equivalence condition (22) reduces to $\sigma_{0}^{2}/a_{0}^{2\kappa+1}=\sigma_{1}^{2}/a_{1}^{2\kappa+1}$, and the microergodic parameter can equivalently be taken as $\sigma^{2}/a^{2\kappa+1}$ (up to the known multiplicative constant $F(\kappa,\mu,d)$). We keep the notation (23) for uniformity, noting that all statements can be rescaled by $F(\kappa,\mu,d)$ when $\mu$ is fixed.
The following result establishes strong consistency and asymptotic normality of $\hat{\theta}_{n}(\mu,a)$ when both $\mu$ and $a$ are treated as fixed. The proof follows the same arguments as Theorem 8 in Bevilacqua et al. (2019) and is omitted.
Theorem 10.
Let $\{Z(\boldsymbol{s}):\boldsymbol{s}\in D\}$, $D\subset\mathbb{R}^{d}$, $d=1,2,3$, be a zero-mean Gaussian RF with covariance $\sigma_{0}^{2}\,\mathcal{H}_{\kappa,\mu,a_{0},d}$, where $\kappa$ and $\mu>1+d/2$ are fixed and known, and $(\sigma_{0}^{2},a_{0})\in(0,\infty)^{2}$. Then, for any fixed $a>0$, as $n\to\infty$:
  1. (i)
    $\hat{\theta}_{n}(\mu,a)\xrightarrow{a.s.}\theta_{0}:=\sigma_{0}^{2}\,F(\kappa,\mu,d)/a_{0}^{2\kappa+1}$;
  2. (ii)
    $\sqrt{n}\,\bigl(\hat{\theta}_{n}(\mu,a)-\theta_{0}\bigr)\xrightarrow{\mathcal{D}}\mathcal{N}\bigl(0,\,2\theta_{0}^{2}\bigr)$.
We now estimate $a$ by ML over the compact set $I=[a_{L},a_{U}]$:
$\hat{a}_{n}:=\arg\max_{a\in I}\,\mathcal{PL}_{n}(\mu,a),\quad 0<a_{L}<a_{U}<\infty,$ (28)
and study the asymptotic behaviour of the random variable $\hat{\theta}_{n}(\mu,\hat{a}_{n})$. The key technical ingredient is a uniform stochastic control of $\hat{\theta}_{n}(\mu,a)$ over $a\in I$. Lemma 2 (Appendix A) provides this control and implies that replacing a fixed $a$ by the random ML estimator $\hat{a}_{n}$ does not affect the fixed-domain limit distribution of the microergodic estimator.
Theorem 11.
Under the same assumptions as Theorem 10, and $a_{0}\in I$, let $\hat{a}_{n}$ be as in (28). Then, as $n\to\infty$:
  1. (i)
    $\hat{\theta}_{n}(\mu,\hat{a}_{n})\xrightarrow{p}\theta_{0}$;
  2. (ii)
    $\sqrt{n}\bigl(\hat{\theta}_{n}(\mu,\hat{a}_{n})-\theta_{0}\bigr)\xrightarrow{\mathcal{D}}\mathcal{N}(0,2\theta_{0}^{2})$.
We finally allow $\mu$ to be unknown and estimate $(\mu,a)$ jointly by ML over the compact set $J\times I$, where $J=[\mu_{L},\mu_{U}]$ and $\mu_{L}>1+d/2$:
$(\hat{\mu}_{n},\hat{a}_{n}):=\arg\max_{(\mu,a)\in J\times I}\,\mathcal{PL}_{n}(\mu,a).$ (29)
As in Case 2, the key step is to control $\hat{\theta}_{n}(\mu,a)$ uniformly over the parameter region $J\times I$. Lemma 3 (Appendix A) provides a two-parameter analogue of Lemma 2. Together with Theorem 10, this yields the following result.
Theorem 12.
Under the same assumptions as Theorem 10, with $(\sigma_{0}^{2},\mu_{0},a_{0})\in(0,\infty)\times J\times I$ and $\mu_{0}>1+d/2$, let $(\hat{\mu}_{n},\hat{a}_{n})$ be as in (29). Then, as $n\to\infty$:
  1. (i)
    $\hat{\theta}_{n}(\hat{\mu}_{n},\hat{a}_{n})\xrightarrow{p}\theta_{0}$;
  2. (ii)
    $\sqrt{n}\,\bigl(\hat{\theta}_{n}(\hat{\mu}_{n},\hat{a}_{n})-\theta_{0}\bigr)\xrightarrow{\mathcal{D}}\mathcal{N}\bigl(0,\,2\theta_{0}^{2}\bigr)$.
Theorems 1112 show that the fixed-domain limit theory for the microergodic parameter under the hypergeometric model is stable with respect to ML estimation of nuisance covariance parameters over compact sets. Our proofs rely on a different strategy than the classical sandwich arguments based on finite-sample monotonicity (Kaufman and Shaby, 2013; Bevilacqua et al., 2019; Ma and Bhadra, 2022): we establish uniform stochastic control of the plug-in microergodic estimator over the relevant compact parameter sets (see Appendix A). This approach yields more general results than those currently available for related compactly supported families, such as the generalized Wendland model, where the restriction $\mu\geq(d+1)/2+\kappa+3$ is required (Bevilacqua et al., 2019).
7 A simulation study
We study the finite-sample behavior of ML estimation under the proposed model $\sigma^{2}\mathcal{H}_{\kappa,\mu,a,d}$. We generate $500$ independent realizations of a zero-mean Gaussian random field on $0,1]^{2}$ ($d=2$) with $\sigma^{2}=1$, $\kappa\in\{0,1,2\}$, and $\mu\in\{4,8\}$. Each realization is simulated via Cholesky decomposition at $n=2500$ locations randomly sampled uniformly over the unit square.
Throughout, we treat $\kappa$ as known and focus on ML estimation of $\sigma^{2}$, $a$, $\mu$, and the microergodic quantity $\sigma^{2}F(\kappa,\mu,d)/a^{2\kappa+1}$. To create a gradual transition from increasing-domain to fixed-domain behavior while keeping the sampling design fixed, we follow Zhang and Zimmerman (2005); Kaufman and Shaby (2013) and vary the scale parameter $a$. Specifically, we consider $a\in\{0.1,0.3,0.6,0.9\}$. Here $a=0.1$ corresponds to weak dependence relative to the unit-square domain and thus approximates an increasing-domain regime, whereas $a=0.9$ induces strong dependence and is much closer to a fixed-domain setting.
Table 1 reports the empirical bias and standard deviation (in parentheses) of the $500$ ML estimates $\widehat{\sigma}^{2}_{i}$, $\widehat{a}_{i}$, $\widehat{\mu}_{i}$ and $\widehat{\sigma}^{2}_{i}F(\kappa,\widehat{\mu}_{i},d)/\widehat{a}_{i}^{2\kappa+1}$, $i=1,\ldots,500$. The main pattern is clear: the individual covariance parameters $\sigma^{2}$, $a$, and $\mu$ become more variable as $a$ increases, whereas the microergodic parameter behaves in the opposite way.
Table 1: Bias and standard deviation (in parentheses) of ML estimates $\widehat{\sigma}^{2}_{i}$, $\widehat{a}_{i}$$\widehat{\mu}_{i}$, and $\widehat{\sigma}^{2}_{i}F(\kappa,\widehat{\mu}_{i},d)/\widehat{a}_{i}^{2\kappa+1}$ for $i=1,\ldots,500$ when estimating a zero-mean Gaussian RF with covariance $\sigma^{2}\mathcal{H}_{\kappa,\mu,a,2}$ on $0,1]^{2}$ ($n=2500$ locations), with $\sigma^{2}=1$, $\kappa\in\{0,1,2\}$, $\mu\in\{4,8\}$, and $a\in\{0.1,0.3,0.6,0.9\}$.
$\kappa=0$ $\kappa=1$ $\kappa=2$
$\mu=4$ $\mu=8$ $\mu=4$ $\mu=8$ $\mu=4$ $\mu=8$
$\widehat{\sigma}^{2}$ $a=0.1$ 0.00299 0.00107 0.00433 0.00184 0.00408 0.00212
(0.04138) (0.03205) (0.04914) (0.03759) (0.04794) (0.03889)
$a=0.3$ 0.01575 0.00504 0.01504 0.00770 0.01173 0.00726
(0.10407) (0.05976) (0.13271) (0.08844) (0.12980) (0.09514)
$a=0.6$ 0.06435 0.01181 0.04389 0.01385 0.03073 0.01553
(0.26849) (0.11312) (0.25614) (0.16863) (0.23655) (0.18127)
$a=0.9$ 0.12097 0.02437 0.09481 0.01848 0.05189 0.02400
(0.38703) (0.17580) (0.41788) (0.24205) (0.32244) (0.25633)
$\widehat{a}$ $a=0.1$ 0.02090 0.08978 0.00276 0.01635 0.00137 0.00612
(0.10465) (0.20471) (0.01397) (0.06865) (0.00827) (0.02363)
$a=0.3$ 0.81440 0.32342 0.02153 0.05872 0.00511 0.01092
(2.75327) (0.75816) (0.10263) (0.26379) (0.04194) (0.06661)
$a=0.6$ 3.91933 2.29317 0.55575 0.45367 0.04601 0.07026
(8.37251) (4.53301) (2.64211) (1.50576) (0.22104) (0.33102)
$a=0.9$ 7.46177 4.50776 1.66114 1.55866 0.16050 0.22433
(14.27313) (7.73620) (5.73980) (4.14495) (0.77112) (0.80880)
$\widehat{\mu}$ $a=0.1$ 1.00129 8.12410 0.17091 1.81217 0.10804 0.81383
(5.12648) (18.51370) (0.88371) (7.52792) (0.66348) (3.13054)
$a=0.3$ 12.17228 9.21399 0.42205 1.98131 0.12180 0.42181
(41.02159) (21.40586) (2.01811) (8.89166) (0.96416) (2.54729)
$a=0.6$ 29.22591 32.41524 5.42300 7.57716 0.55245 1.34734
(61.35913) (63.46363) (25.61297) (24.98706) (2.54928) (6.27535)
$a=0.9$ 36.69791 42.83158 10.93102 17.38350 1.29228 2.84537
(68.21306) (72.15506) (37.54848) (45.95964) (6.09577) (10.03183)
$\frac{\widehat{\sigma}^{2}F(\kappa,\mu,d)}{\widehat{a}^{2\kappa+1}}$ $a=0.1$ 0.02847 -0.15330 567.68210 7656.58100 $1.40740\times 10^{7}$ $4.32344\times 10^{8}$
(1.79328) (5.08178) (12695.11000) (107860.90000) ($2.18681\times 10^{8}$) ($3.79118\times 10^{9}$)
$a=0.3$ 0.00880 -0.04215 -0.59747 -0.93738 -2994.48500 -48566.15000
(0.45226) (0.95325) (296.09960) (1651.58800) (448665.40000) (4813284.00000)
$a=0.6$ 0.00331 -0.01499 0.66440 -0.59339 165.89200 -325.76500
(0.21584) (0.43426) (32.70379) (167.22660) (11788.42000) (115466.10000)
$a=0.9$ 0.00148 -0.01088 0.05882 -0.61039 -4.95065 -62.59252
(0.14205) (0.28237) (9.47927) (46.79391) (1482.45500) (13819.55000)
This is especially evident for $\widehat{a}$ and $\widehat{\mu}$. For instance, when $\kappa=0$ and $\mu=4$, the empirical standard deviation of $\widehat{a}$ increases from $0.10465$ at $a=0.1$ to $2.75327$ at $a=0.3$, then to $8.37251$ at $a=0.6$, and to $14.27313$ at $a=0.9$; the standard deviation of $\widehat{\mu}$ rises from $5.12648$ to $41.02159$, then to $61.35913$, and finally to $68.21306$. The estimator $\widehat{\sigma}^{2}$ is more stable, but it follows the same qualitative trend.
These results are fully consistent with fixed-domain asymptotics: the individual covariance parameters are not consistently estimable, and stronger dependence effectively reduces the amount of independent information. The effect of $\mu$ is visible but secondary relative to the role of $a$. In several configurations, $\mu=8$ leads to somewhat larger variability than $\mu=4$, especially for $\widehat{\mu}$ and in the stronger-dependence regimes. Likewise, the overall inferential difficulty increases with $\kappa$, particularly for quantities involving the scale parameter.
The microergodic estimator behaves in the opposite way. For
$\widehat{\theta}=\frac{\widehat{\sigma}^{2}F(\kappa,\widehat{\mu},d)}{\widehat{a}^{2\kappa+1}},$
both bias and dispersion decrease markedly as $a$ increases. This is most striking for larger $\kappa$. For instance, when $\kappa=2$ and $\mu=4$, the empirical standard deviation of $\widehat{\theta}$ drops from $2.18681\times 10^{8}$ at $a=0.1$, to $4.48665\times 10^{5}$ at $a=0.3$, to $1.17884\times 10^{4}$ at $a=0.6$, and to $1482.45500$ at $a=0.9$. Even for $\kappa=0$, where $\widehat{\theta}$ is already much more stable, the same monotone improvement is evident.
This contrast is the central empirical message of the simulation study. As the regime moves from increasing-domain-like to fixed-domain-like behavior, the individual covariance parameters deteriorate, while the microergodic parameter improves. The distinction becomes sharper as $\kappa$ increases, since the factor $a^{-(2\kappa+1)}$ amplifies scale estimation error in weak-dependence settings, but defines the identifiable combination under fixed-domain asymptotics.
Figure 4 shows normal QQ-plots for the standardized microergodic estimates in the strong-dependence setting ($a=0.9$) with $\kappa=1$, comparing $\mu=4$ (left) and $\mu=8$ (right). Specifically, we plot
$Z_{i}\;=\;\frac{\sqrt{n}\,\big(\widehat{\theta}_{i}-\theta_{0}\big)}{\sqrt{2}\,\theta_{0}},\qquad\widehat{\theta}_{i}=\widehat{\sigma}^{2}_{i}F(\kappa,\widehat{\mu}_{i},d)/\widehat{a}_{i}^{2\kappa+1},\quad\theta_{0}=\sigma^{2}F(\kappa,\mu,d)/a^{2\kappa+1},$
for $i=1,\dots,500$. In both cases, the QQ-plots align reasonably well with the standard normal reference line, supporting the relevance of the fixed-domain Gaussian approximation at this sample size.
Overall, the simulations strongly support the theoretical results of Section 6: the individual covariance parameters are increasingly unstable as dependence strengthens, whereas the microergodic parameter shows the opposite trend and is the meaningful inferential target under fixed-domain asymptotics.
Refer to caption Refer to caption
($\mu=4$) ($\mu=8$)
Figure 4: Normal $Q$-plots of standardized microergodic estimates in the fixed-domain setting ($a=0.9$) with $\kappa=1$ and $\sigma^{2}=1$, comparing $\mu=4$ (left) and $\mu=8$ (right). The standardized statistics are $Z_{i}=\sqrt{n}\,(\widehat{\theta}_{i}-\theta_{0})/(\sqrt{2}\,\theta_{0})$, where $\widehat{\theta}_{i}=\widehat{\sigma}^{2}_{i}F(\kappa,\widehat{\mu}_{i},d)/\widehat{a}_{i}^{2\kappa+1}$, $\theta_{0}=\sigma^{2}F(\kappa,\mu,d)/a^{2\kappa+1}$, and $n=2500$.
8 A real data application
We consider data from WorldClim (www.worldclim.org), which provides high-resolution gridded climate surfaces for 1970–2000 (Fick and Hijmans, 2017) and can be accessed in R via geodata (Hijmans et al., 2024). We analyze September mean temperature over Madagascar. The full dataset consists of 725,330 grid-cell predictions obtained by interpolating weather-station data using thin-plate splines with geographic and satellite-derived covariates (Fick and Hijmans, 2017). To make ML estimation computationally feasible, we randomly sample 3,500 grid cells and use UTM-projected coordinates.
Following Li and Zhang (2011), the data are first detrended using spline regression to remove large-scale spatial trends. The residuals $y(\mathbf{s}_{i}),i=1,\ldots,3500$, are then treated as observations from a zero-mean RF $\{Y(\mathbf{s}):\mathbf{s}\in D\}$. Figure 5 displays the spatial map of the residuals and the empirical semivariogram. The plots suggest that a weakly stationary RF with a potentially heavy-tailed marginal distribution provides a reasonable modeling framework.
Refer to caption Refer to caption
Figure 5: Left: spatial map of residuals. Right: empirical semivariogram.
Let $Z(\mathbf{s})$ be a standard Gaussian RF with isotropic correlation function. We consider two specifications for $Y(\mathbf{s})$:
  • Gaussian RF: $Y(\mathbf{s})=\sigma Z(\mathbf{s})$;
  • Tukey-$h$ RF (Xu and Genton, 2017; Caamaño-Carrillo et al., 2024): $Y(\mathbf{s})=\sigma Z(\mathbf{s})\exp\!\left(\tfrac{hZ^{2}(\mathbf{s})}{2}\right)$, where $h\in[0,1/2)$ controls tail heaviness. Since the transformation is monotone, the Tukey-$h$ RF preserves the geometric properties of the underlying Gaussian field.
For the correlation of $Z(\mathbf{s})$, we compare three models: ${\cal MT}_{\kappa,a}$, ${\cal GW}_{\kappa,\mu,a}$, and the proposed ${\cal H}_{\kappa,\mu,a,2}$. Here $\kappa$ denotes smoothness, $a$ the scale parameter, and $\mu$ the shape parameter.
We adopt an increasing-domain framework and estimate all covariance parameters by ML. Standard errors are computed from the inverse Fisher information matrix. For the Tukey-$h$ RF, ML estimation is implemented via the Lambert-$W$ function (see equation 10 in Caamaño-Carrillo et al. (2024)).
Table 2 reports ML estimates, associated standard errors, maximized log-likelihood values, and AIC values. Overall, the AIC criterion favors the Tukey-$h$ specification, irrespective of the correlation model. Among the correlation structures, the compactly supported ${\cal GW}_{\kappa,\mu,a}$ and ${\cal H}_{\kappa,\mu,a,2}$ achieve the best AIC values.
Figure 6 compares the normalized histogram of the residuals with the fitted density (left panels) and the empirical semivariogram with the fitted semivariogram (right panels), for the Gaussian (first row) and Tukey-$h$ (second row) models using the estimated ${\cal H}_{\kappa,\mu,a,2}$ correlation. The Tukey-$h$ RF provides a noticeably better fit to both marginal tail behavior and spatial dependence.
Table 2: ML estimates with associated standard error (in parentheses), RMSE, MAE and percentage of zero values in the covariance matrix for the Gaussian and Tukey-$h$ RFs with underlying correlation functions ${\cal MT}_{\kappa,a}$, ${\cal GW}_{\kappa,\mu,a}$ and ${\cal H}_{\kappa,\mu,a,2}$. Last row reports the time (in seconds) needed to compute the Cholesky factor of the covariance matrix.
Gaussian RF Tukey-$h$ RF
${\cal MT}_{\kappa,a}$ ${\cal GW}_{\kappa,\mu,a}$ ${\cal H}_{\kappa,\mu,a,2}$ ${\cal MT}_{\kappa,a}$ ${\cal GW}_{\kappa,\mu,a}$ ${\cal H}_{\kappa,\mu,a,2}$
$\sigma^{2}$ $1.2054$ $1.2066$ $1.2061$ $0.6998$ $0.6986$ $0.6984$
$(0.0553)$ $(0.0541)$ $(0.0540)$ $(0.0418)$ $(0.0403)$ $(0.0403)$
$\kappa$ $0.6079$ $0.0305$ $0.0313$ $0.5767$ 0.0036-0.0036 0.0043-0.0043
$(0.0293)$ $(0.0307)$ $(0.0312)$ $(0.0265)$ $(0.0237)$ $(0.0237)$
$a$ $16.0229$ $85.8058$ $87.9751$ $18.5405$ $76.0127$ $77.6552$
$(1.3104)$ $(20.4440)$ $(20.9644)$ $(1.6183)$ $(12.0024)$ $(10.5911)$
$\mu$ $4.2150$ $3.8914$ $3.1346$ $2.8085$
$(1.3021)$ $(1.3026)$ $(0.6548)$ $(0.5764)$
$h$ $0.1885$ $0.1899$ $0.1898$
$(0.0172)$ $(0.0170)$ $(0.0170)$
loglik 3900.01-3900.01 3896.88-3896.88 3896.94-3896.94 3689.88-3689.88 3683.81-3683.81 3683.94-3683.94
AIC $7806$ $7802$ $7802$ $7388$ $7378$ $7378$
RMSE $0.6593$ $0.6588$ $0.6588$ $0.6586$ $0.6581$ $0.6581$
MAE $0.4068$ $0.4065$ $0.4065$ $0.4073$ $0.4071$ $0.4070$
$\%$ 0 $96.45$ $96.23$ 0 $97.18$ $97.06$
TimeTime $7.22$ $0.04$ $0.04$ $6.26$ $0.03$ $0.03$
We assess predictive performance via repeated cross-validation. For each of 100 repetitions, 85% of locations are used for parameter estimation and the remaining 15% for prediction. Table 2 reports the average RMSE and MAE. The best predictive performance is achieved by the Tukey-$h$ RF combined with compactly supported correlations ${\cal GW}_{\kappa,\mu,a}$ and ${\cal H}_{\kappa,\mu,a,2}$.
[…]
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 6: First row: normalized histogram versus estimated density and empirical semivariogram versus estimated semivariogram when using an estimated Gaussian RF with $\sigma^{2}\mathcal{H}_{\kappa,\mu,a,2}$ correlation model (from left to right). Second row: the same graphics using a Tukey-$h$ RF.
9 Concluding Remarks
We revisited the Gauss hypergeometric covariance class (Emery and Alegría, 2022), a flexible family that includes the celebrated Matérn and generalized Wendland models as special cases. We first refined its validity conditions, deriving sharp constraints over a substantial region of the parameter space and pinpointing where a gap between necessary and sufficient conditions still remains.
We then clarified that the class features two distinct non-identifiability mechanisms: a structural one, stemming from an intrinsic symmetry in the shape parameters, and the classical fixed-domain one, whereby different parameter values can induce equivalent Gaussian measures and only microergodic combinations are consistently estimable. The structural ambiguity is removed by an ordering constraint, yielding a unique representation on the restricted space.
Within this identifiable formulation, we introduced a parsimonious compactly supported subclass, $\mathcal{H}$, selected via a maximum integral range criterion. This choice provides a canonical representative with maximal low-frequency power among admissible models with fixed smoothness and support. The resulting family generalizes the classical spherical model (Chilès and Delfiner, 2012) while retaining continuous smoothness control, compact support, and simple necessary and sufficient validity conditions. We also derived scale-mixture representations and explicit links to the Matérn model through compact-support reparameterizations.
From an inferential standpoint, we developed fixed-domain asymptotics for likelihood-based estimation under $\mathcal{H}$. For fixed smoothness, we proved consistency and asymptotic normality for the microergodic-parameter estimator, and established stability when nuisance parameters are estimated by ML over compact sets. Simulations and a climate-data application indicate that $\mathcal{H}$ can offer a competitive compromise between Matérn and generalized Wendland models, combining accuracy with sparse covariance matrices and notable computational gains, for both Gaussian and Tukey-hh random fields.
Several directions remain open, including closing the remaining validity gap in Theorem 2, extending fixed-domain theory to unknown smoothness, and developing scalable alternatives to full ML. Finally, anisotropic, nonstationary, and multivariate extensions would further broaden the practical scope of the hypergeometric framework.
Although closed forms are available only in special cases, efficient numerical evaluation of Gauss hypergeometric functions is widely supported, including in the GNU Scientific Library (Gough, 2009) and in major environments such as R, MATLAB, and Python. The GeoModels package (Bevilacqua et al., 2026) used in this paper computes the proposed model via the SciPy implementation (Virtanen et al., 2020).
Acknowledgement
This work was funded by the National Agency for Research and Development of Chile, under grants ANID FONDECYT 1240308 (M. Bevilacqua), ANID project Data Observatory Foundation DO210001 (M. Bevilacqua), MATH-AMSUD1167 22-MATH-06 (AMSUD220041) (M. Bevilacqua), ANID FONDECYT 1250862 (C. Caamaño-Carrillo) and ANID PIA AFB230001 (X. Emery).
Appendix A Proofs
A.1 Proof of Theorem 1
A preliminary Lemma is needed for the proof of Theorem 1. The following result can be obtained from the asymptotic expansion of the generalized hypergeometric function $_{1}F_{2}$ for large argument, as given in Luke (1969).
Lemma 1.
Let $\widehat{{\cal GH}}_{\delta,\beta,\gamma,a}(z)$ be defined as in (7). Assume $\delta>\frac{d}{2}$, $2(\beta-\delta)(\gamma-\delta)\geq\delta$, and $2(\beta+\gamma)\geq 6\delta+1$. Then:
  1. 1.
    As $z\to\infty$,
    $\widehat{{\cal GH}}_{\delta,\beta,\gamma,a}(z)=K_{\delta,d}\Big[c_{3}\,(za)^{-2\delta}\{1+\mathcal{O}(z^{-2})\}+c_{4}\,(za)^{\frac{1}{2}+\delta-\beta-\gamma}\{\cos(za+c_{5})+\mathcal{O}(z^{-1})\}\Big],$
    where
    $c_{3}=\frac{2^{2\delta}\Gamma(\beta)\Gamma(\gamma)}{\Gamma(\gamma-\delta)\Gamma(\beta-\delta)},\qquad c_{4}=\frac{\pi^{-1/2}\Gamma(\beta)\Gamma(\gamma)}{2^{\frac{1}{2}+\delta-\beta-\gamma}\Gamma(\delta)},\qquad c_{5}=\frac{\pi}{2}\!\left(\frac{1}{2}+\delta-\beta-\gamma\right).$
  2. 2.
    $\widehat{{\cal GH}}_{\delta,\beta,\gamma,a}(z)\asymp z^{-2\delta}$ as $z\to\infty$.
Proof of Theorem 1.
Let
$\rho_{i}(\boldsymbol{h})=\sigma_{i}^{2}\,\varphi_{i}(\|\boldsymbol{h}\|),\qquad\varphi_{i}(x)={\cal GH}_{\delta,\beta_{i},\gamma_{i},a_{i}}(x),\qquad i=0,1,$
and let $f_{i}$ denote the (isotropic) spectral density of $\rho_{i}$ in the sense of (1). Then $f_{i}(z)=\sigma_{i}^{2}\,\widehat{\varphi}_{i}(z)$, where $\widehat{\varphi}_{i}(z)=\widehat{\cal GH}_{\delta,\beta_{i},\gamma_{i},a_{i}}(z)$ is given by (7). Since we are considering valid covariance models, $\widehat{\varphi}_{i}(z)\geq 0$ for all $z\geq 0$, and the Gamma factors in (7) are finite (in particular, $\beta_{i}>d/2$, $\gamma_{i}>d/2$ and $\beta_{i}>\delta$, $\gamma_{i}>\delta$).
We use the standard equivalence criterion for stationary Gaussian measures on bounded domains (see, e.g., Ibragimov and Rozanov (1978) and Stein (2004)): if for some $\eta>0$ the function $f_{0}(z)\,z^{\eta}$ is bounded away from 0 and $\infty$ as $z\to\infty$, and if for some $c>0$
$\int_{c}^{\infty}z^{d-1}\left(\frac{f_{1}(z)-f_{0}(z)}{f_{0}(z)}\right)^{2}\,dz<\infty,$ (30)
then $P(\rho_{0})$ and $P(\rho_{1})$ are equivalent on the paths over any bounded $D\subset\mathbb{R}^{d}$. (For $d=1,2,3$ this condition is also sharp in the present regularly-varying setting; see the necessity argument below.)
Set, for $i=0,1$,
$\lambda_{i}:=\frac{1}{2}+\delta-\beta_{i}-\gamma_{i}.$
By assumption, for each $i=0,1$ we have $\delta>\frac{d}{2}$ and $2(\beta_{i}-\delta)(\gamma_{i}-\delta)\geq\delta.$ Moreover, the theorem assumes $\beta_{i}+\gamma_{i}>\frac{d+1}{2}+3\delta,$ hence $2(\beta_{i}+\gamma_{i})>(d+1)+6\delta\geq 6\delta+2>6\delta+1,$ so that in particular $2(\beta_{i}+\gamma_{i})\geq 6\delta+1$. Therefore all assumptions of Lemma 1 hold for $\widehat{\varphi}_{i}$. By Lemma 1 (applied to $\widehat{\varphi}_{i}$), as $z\to\infty$,
$\widehat{\varphi}_{i}(z)=K_{\delta,d}\,L(\beta_{i},\gamma_{i})\,a_{i}^{\,d-2\delta}\,z^{-2\delta}\,\{1+\mathcal{O}(z^{-2})\}\;+\;\mathcal{O}\!\big(z^{\lambda_{i}}\big),$ (31)
where $K_{\delta,d}:=\Gamma(\delta)/(2^{2\delta}\pi^{d/2}\Gamma(\delta-d/2))>0$ depends only on $(\delta,d)$ and
$L(\beta_{i},\gamma_{i})=\frac{2^{2\delta-d}\Gamma(\beta_{i}-d/2)\Gamma(\gamma_{i}-d/2)}{\Gamma(\gamma_{i}-\delta)\Gamma(\beta_{i}-\delta)}.$
Multiplying by $\sigma_{i}^{2}$, we obtain for $f_{i}(z)=\sigma_{i}^{2}\widehat{\varphi}_{i}(z)$:
$f_{i}(z)=A_{i}\,z^{-2\delta}\{1+\mathcal{O}(z^{-2})\}+\mathcal{O}\!\big(z^{\lambda_{i}}\big),\qquad A_{i}:=\sigma_{i}^{2}\,K_{\delta,d}\,L(\beta_{i},\gamma_{i})\,a_{i}^{\,d-2\delta}.$ (32)
Since the theorem assumes $\beta_{i}+\gamma_{i}>\frac{d+1}{2}+3\delta,$ we have
$\lambda_{i}+2\delta=\frac{1}{2}+3\delta-(\beta_{i}+\gamma_{i})<-\frac{d}{2},$
hence $z^{\lambda_{i}}=o(z^{-2\delta})$ and (32) yields
$f_{i}(z)\sim A_{i}\,z^{-2\delta}\qquad(z\to\infty),\quad\text{with }A_{i}>0.$ (33)
In particular, $f_{0}(z)\,z^{2\delta}$ is bounded away from 0 and $\infty$ for large $z$, so (30) can be applied with $\eta=2\delta$.
Sufficiency. Assume (10). Since $K_{\delta,d}>0$ is common to both models, (10) is equivalent to $A_{0}=A_{1}$ (because $A_{i}$ differs from $\sigma_{i}^{2}L(\beta_{i},\gamma_{i})/a_{i}^{2\delta-d}$ only by the common factor $K_{\delta,d}$). Hence, by (32),
$f_{1}(z)-f_{0}(z)=A_{0}z^{-2\delta}\mathcal{O}(z^{-2})\;+\;\mathcal{O}\!\big(z^{\lambda_{0}}\big)\;+\;\mathcal{O}\!\big(z^{\lambda_{1}}\big).$
Dividing by $f_{0}(z)\sim A_{0}z^{-2\delta}$ gives
$\frac{f_{1}(z)-f_{0}(z)}{f_{0}(z)}=\mathcal{O}(z^{-2})+\mathcal{O}\!\big(z^{\lambda_{0}+2\delta}\big)+\mathcal{O}\!\big(z^{\lambda_{1}+2\delta}\big).$
Therefore, for large $z$,
$z^{d-1}\left(\frac{f_{1}(z)-f_{0}(z)}{f_{0}(z)}\right)^{2}\;\lesssim\;z^{d-1}z^{-4}\;+\;z^{d-1}z^{2(\lambda_{0}+2\delta)}\;+\;z^{d-1}z^{2(\lambda_{1}+2\delta)}.$
The first term is integrable at $\infty$ iff $d<4$, i.e. $d=1,2,3$. For the other terms, note that
$d-1+2(\lambda_{i}+2\delta)=d-1+2\Big(\tfrac{1}{2}+3\delta-(\beta_{i}+\gamma_{i})\Big)=d+6\delta-2(\beta_{i}+\gamma_{i}),$
which is <1<-1 exactly when $\beta_{i}+\gamma_{i}>(d+1)/2+3\delta$, assumed in the theorem. Thus the integral (30) is finite, and by the criterion $P(\rho_{0})$ and $P(\rho_{1})$ are equivalent on bounded $D$.
Necessity: Assume the theorem’s condition $\beta_{i}+\gamma_{i}>(d+1)/2+3\delta$ holds for $i=0,1$ and suppose (10) fails. Then $A_{1}\neq A_{0}$. Using (33),
$\frac{f_{1}(z)-f_{0}(z)}{f_{0}(z)}=\frac{A_{1}z^{-2\delta}-A_{0}z^{-2\delta}+o(z^{-2\delta})}{A_{0}z^{-2\delta}\{1+o(1)\}}\longrightarrow\frac{A_{1}-A_{0}}{A_{0}}\neq 0\qquad(z\to\infty).$
Hence there exist $c>0$ and $\varepsilon>0$ such that for all $z\geq c$,
$\left|\frac{f_{1}(z)-f_{0}(z)}{f_{0}(z)}\right|\geq\varepsilon,$
and therefore
$\int_{c}^{\infty}z^{d-1}\left(\frac{f_{1}(z)-f_{0}(z)}{f_{0}(z)}\right)^{2}dz\;\geq\;\varepsilon^{2}\int_{c}^{\infty}z^{d-1}\,dz=\infty.$
Thus (30) fails and the measures cannot be equivalent. Consequently, equivalence implies $A_{1}=A_{0}$, which is exactly (10).
Combining the sufficiency and necessity arguments proves the claimed “if and only if” statement. ∎
A.2 Proof of Theorem 2
Proof.
Consider the model $\mathcal{GH}_{\delta,\delta+\frac{\mu}{2},\delta+\frac{\mu}{2}+l,a}$ in dimension $d$, where $\delta=\frac{d+1}{2}+\kappa$, $\kappa>-\frac{1}{2}$.
Note that the model $\mathcal{GH}_{\delta,\beta,\gamma,a}$ is invariant under interchanging $\beta$ and $\gamma$, which corresponds to $l\mapsto-l$. The admissible region is therefore symmetric with respect to $l=0$, and restricting to $l\geq 0$ entails no loss of generality. We henceforth assume $l\geq 0$.
Sufficiency. Under $\beta=\delta+\frac{\mu}{2}$, $\gamma=\delta+\frac{\mu}{2}+l$, the sufficient conditions of Section 3 become:
  1. (a)
    $\kappa>-\frac{1}{2}$,
  2. (b)
    $\mu(\mu+2l)\geq d+2\kappa+1$,
  3. (c)
    $\mu+l\geq\frac{d}{2}+\kappa+1$.
Condition (b) is equivalent to the quadratic inequality
$\mu^{2}+2l\mu-(d+2\kappa+1)\geq 0,$
whose positive branch gives
$\mu\geq-l+\sqrt{l^{2}+d+2\kappa+1}.$
Only the positive branch is admissible since condition (c) implies $\mu>-l$ when $l\geq 0$. For each fixed $l$, the sufficient conditions therefore reduce to
$\mu\geq\max\!\Big\{\frac{d}{2}+\kappa+1-l,\;-l+\sqrt{l^{2}+d+2\kappa+1}\Big\}.$
A direct algebraic comparison shows that
$\frac{d}{2}+\kappa+1-l\geq-l+\sqrt{l^{2}+d+2\kappa+1}\quad\Longleftrightarrow\quad 0\leq l\leq\frac{d}{2}+\kappa.$
Hence condition (c) is binding when $0\leq l\leq\frac{d}{2}+\kappa$, giving $\mu\geq\delta-l+\frac{1}{2}$, and condition (b) is binding when $l>\frac{d}{2}+\kappa$, giving $\mu\geq\sqrt{2\kappa+l^{2}+d+1}-l$. This yields exactly the two regimes in points 1 and 2.
Necessity. Assume $\mathcal{GH}_{\delta,\beta,\gamma,a}\in\Phi_{d}$. Since the Gamma prefactors in the spectral density are positive whenever $\beta>\delta$ and $\gamma>\delta$, nonnegativity of the spectral density requires
$_{1}F_{2}\!\left(\delta;\beta,\gamma;-\frac{(za)^{2}}{4}\right)\geq 0\quad\text{for all }z\geq 0.$
By Cho and Yun (2018, Theorem 3.1), this implies $\beta>\delta$, $\gamma>\delta$, and $\beta+\gamma\geq 3\delta+\frac{1}{2}$. The third condition, under the parameterization $\beta=\delta+\frac{\mu}{2}$, $\gamma=\delta+\frac{\mu}{2}+l$, gives
$\mu+l\geq\frac{d}{2}+\kappa+1,\quad\text{i.e.,}\quad\mu\geq\delta-l+\frac{1}{2},$
which is the necessary lower bound for $\mu$ in both regimes. In Regime 1 ($0\leq l\leq\frac{d}{2}+\kappa$), this coincides with the sufficient condition in point 1, so the result is an if and only if. In Regime 2 ($l>\frac{d}{2}+\kappa$), this provides a necessary lower bound that is strictly weaker than the sufficient condition in point 2; whether the gap can be closed remains an open question.
This completes the proof. ∎
A.3 Proof of Theorem 3
Proof.
Since $\phi(\|\cdot\|)$ is compactly supported it is absolutely integrable and, by the spectral representation associated with (1),
$A=(2\pi)^{d}\,\widehat{\phi}(0).$
For $\phi(x)=\mathcal{GH}_{\delta,\beta,\gamma,a}(x)$, the spectral density is (7). Evaluating at $z=0$ and using $_{1}F_{2}(\cdot;\,0)=1$ gives
$\widehat{{\cal GH}}_{\delta,\beta,\gamma,a}(0)=\frac{\Gamma(\delta)\Gamma(\beta-\frac{d}{2})\Gamma(\gamma-\frac{d}{2})a^{d}}{2^{d}\pi^{\frac{d}{2}}\Gamma\left(\delta-\frac{d}{2}\right)\Gamma(\beta)\Gamma(\gamma)}.$
Hence
$A=(2\pi)^{d}\widehat{{\cal GH}}_{\delta,\beta,\gamma,a}(0)=\pi^{d/2}a^{d}\,\frac{\Gamma(\delta)\Gamma(\beta-\frac{d}{2})\Gamma(\gamma-\frac{d}{2})}{\Gamma\left(\delta-\frac{d}{2}\right)\Gamma(\beta)\Gamma(\gamma)}.$
Substituting $\delta=\frac{d+1}{2}+\kappa$, $\beta=\delta+\frac{\mu}{2}$, $\gamma=\beta+l$ yields (14). ∎
A.4 Proof of Theorem 4
Proof.
Since $A_{\kappa,\mu,l,a,d}=a^{d}\,A_{\kappa,\mu,l,1,d}$, the factor $a^{d}$ is irrelevant for monotonicity and maximizers in $\mu$. Hence we set $a=1$.
Fix $l\geq 0$ and consider
$LA_{\kappa,d}(\mu,l):=\log\!\big(A_{\kappa,\mu,l,1,d}\big),$
where
$A_{\kappa,\mu,l,1,d}=\frac{\pi^{d/2}\Gamma\!\left(\frac{d+1}{2}+\kappa\right)\,\Gamma\!\left(\frac{\mu+1}{2}+\kappa\right)\,\Gamma\!\left(\frac{\mu+1}{2}+\kappa+l\right)}{\Gamma\!\left(\kappa+\frac{1}{2}\right)\,\Gamma\!\left(\frac{\mu+1}{2}+\kappa+\frac{d}{2}\right)\,\Gamma\!\left(\frac{\mu+1}{2}+\kappa+l+\frac{d}{2}\right)}.$
Let $\psi=\Gamma^{\prime}/\Gamma$ be the digamma function and set
$x=\frac{\mu+1}{2}+\kappa\qquad\Big(\text{so }\frac{dx}{d\mu}=\frac{1}{2}\Big).$
On the validity region of the model (hence in particular on the admissible domain in $\mu$), the Gamma factors above are well-defined and positive, so $x>0$ and $x+l>0$. Differentiating with respect to $\mu$ gives
$\frac{d}{d\mu}LA_{\kappa,d}(\mu,l)=\frac{1}{2}\Big[\psi(x)-\psi\!\left(x+\tfrac{d}{2}\right)+\psi(x+l)-\psi\!\left(x+l+\tfrac{d}{2}\right)\Big].$
Since $\psi$ is strictly increasing on $(0,\infty)$ and $d/2>0$, each difference $\psi(x)-\psi(x+d/2)$ and $\psi(x+l)-\psi(x+l+d/2)$ is $\leq 0$. Therefore
$\frac{d}{d\mu}LA_{\kappa,d}(\mu,l)\leq 0\quad\text{for all admissible }\mu,$
hence $\mu\mapsto A_{\kappa,\mu,l,a,d}$ is non-increasing on its admissible domain. Consequently, its maximum is attained at the smallest admissible value of $\mu$.
We now identify this smallest admissible value using Theorem 2. If $0\leq l\leq d/2+\kappa$, Theorem 2 provides the exact validity condition $\mu\geq\mu_{1}(l):=\delta-l+\frac{1}{2}$, hence $\mu^{\star}(l)=\mu_{1}(l)$.
If $l>d/2+\kappa$, Theorem 2 guarantees validity for all $\mu\geq\mu_{2}(l):=\sqrt{2\kappa+l^{2}+d+1}-l$. In this regime, we take as admissible domain the parameter range ensured by Theorem 2 (i.e., $\mu\geq\mu_{2}(l)$), and by monotonicity the maximum over this admissible domain is attained at $\mu^{\star}(l)=\mu_{2}(l)$. ∎
A.5 Proof of Theorem 5
Proof.
Let $\tau=\frac{d}{2}+\kappa$. Since $\log$ is strictly increasing, maximizing $A_{\kappa,\mu,l,a,d}$ is equivalent to maximizing
$LA_{\kappa,d}(\mu,l):=\log A_{\kappa,\mu,l,1,d},$
and the factor $a^{d}$ plays no role for maximizers.
By Theorem 4, for each fixed $l$ the maximizer in $\mu$ is $\mu=\mu_{1}(l)$ for $0\leq l\leq\tau$ and $\mu=\mu_{2}(l)$ for $l>\tau$, where
$\mu_{1}(l)=\tau+1-l\qquad(0\leq l\leq\tau),\qquad\mu_{2}(l)=\sqrt{l^{2}+(2\kappa+d+1)}-l\qquad(l>\tau).$
Hence it suffices to study the two one-variable functions
$F_{1}(l):=LA_{\kappa,d}(\mu_{1}(l),l)\quad\text{on }[0,\tau],\qquad F_{2}(l):=LA_{\kappa,d}(\mu_{2}(l),l)\quad\text{on }(\tau,\infty).$
Write $\psi=\Gamma^{\prime}/\Gamma$ for the digamma function. Recall that $\psi$ is concave on $(0,\infty)$, so for every $h>0$ the symmetric increment
$\Delta_{h}(t):=\psi(t+h)-\psi(t-h)$
is decreasing in $t$ (because $\Delta_{h}^{\prime}(t)=\psi^{\prime}(t+h)-\psi^{\prime}(t-h)\leq 0$).
Case $i=1$ ($0\leq l\leq\tau$).
For $0\leq l\leq\tau$, $\mu_{1}(l)=\tau+1-l$. Set
$u(l):=\frac{\mu_{1}(l)+1}{2}+\kappa=\frac{d}{4}+\frac{3\kappa}{2}+1-\frac{l}{2}.$
Using the explicit form of $A_{\kappa,\mu,l,1,d}$, $F_{1}(l)$ differs from
$\log\Gamma(u(l))+\log\Gamma(u(l)+l)-\log\Gamma\!\left(u(l)+\frac{d}{2}\right)-\log\Gamma\!\left(u(l)+l+\frac{d}{2}\right)$
only by an additive constant independent of $l$. Differentiating gives
$F_{1}^{\prime}(l)$ $=u^{\prime}(l)\left[\psi(u(l))+\psi(u(l)+l)-\psi\!\left(u(l)+\frac{d}{2}\right)-\psi\!\left(u(l)+l+\frac{d}{2}\right)\right]$
$\qquad+\psi(u(l)+l)-\psi\!\left(u(l)+l+\frac{d}{2}\right),$
where $u^{\prime}(l)=-\tfrac{1}{2}$. A simple rearrangement yields
$F_{1}^{\prime}(l)=\frac{1}{2}\left[\psi(u(l)+l)-\psi(u(l))+\psi\!\left(u(l)+\frac{d}{2}\right)-\psi\!\left(u(l)+l+\frac{d}{2}\right)\right].$
Now set
$x:=\frac{d}{4}+\frac{3\kappa}{2}+1\quad\Longrightarrow\quad u(l)=x-\frac{l}{2},\qquad u(l)+l=x+\frac{l}{2}.$
Therefore
$F_{1}^{\prime}(l)=\frac{1}{2}\!\left[\psi\!\left(x+\tfrac{l}{2}\right)-\psi\!\left(x-\tfrac{l}{2}\right)+\psi\!\left(x-\tfrac{l}{2}+\tfrac{d}{2}\right)-\psi\!\left(x+\tfrac{l}{2}+\tfrac{d}{2}\right)\right]=\frac{1}{2}\!\left(\Delta_{l/2}(x)-\Delta_{l/2}\!\left(x+\tfrac{d}{2}\right)\right).$
Since $\Delta_{l/2}$ is decreasing and $x+\tfrac{d}{2}>x$, we have $\Delta_{l/2}(x)\geq\Delta_{l/2}(x+\tfrac{d}{2})$, hence $F_{1}^{\prime}(l)\geq 0$ on $0,\tau]$. Thus $F_{1}$ is non-decreasing on $0,\tau]$ and its maximum is attained at $l=\tau$.
Case $i=2$ ($l>\tau$).
For $l>\tau$, $\mu_{2}(l)=\sqrt{l^{2}+(2\kappa+d+1)}-l$. Define
$u(l):=\frac{\mu_{2}(l)+1}{2}+\kappa>0,\qquad x(l):=u(l)+\frac{l}{2}=\kappa+\frac{1+\sqrt{l^{2}+(2\kappa+d+1)}}{2}.$
Then $u(l)=x(l)-\tfrac{l}{2}$, $u(l)+l=x(l)+\tfrac{l}{2}$, and
$x^{\prime}(l)=\frac{1}{2}\cdot\frac{l}{\sqrt{l^{2}+(2\kappa+d+1)}}\in\left(0,\frac{1}{2}\right),\qquad u^{\prime}(l)=x^{\prime}(l)-\frac{1}{2},\qquad\frac{d}{dl}(u(l)+l)=x^{\prime}(l)+\frac{1}{2}.$
As above, up to an additive constant independent of $l$,
$F_{2}(l)=\log\Gamma(u(l))+\log\Gamma(u(l)+l)-\log\Gamma\!\left(u(l)+\frac{d}{2}\right)-\log\Gamma\!\left(u(l)+l+\frac{d}{2}\right).$
A direct differentiation, using $u^{\prime}(l)=x^{\prime}(l)-\tfrac{1}{2}$ and $\tfrac{d}{dl}(u(l)+l)=x^{\prime}(l)+\tfrac{1}{2}$, gives
$F_{2}^{\prime}(l)=\Bigl(x^{\prime}(l)-\tfrac{1}{2}\Bigr)\bigl[\psi(u)-\psi\!\left(u+\tfrac{d}{2}\right)\bigr]+\Bigl(x^{\prime}(l)+\tfrac{1}{2}\Bigr)\bigl[\psi(u+l)-\psi\!\left(u+l+\tfrac{d}{2}\right)\bigr],$ (34)
where $u=u(l)$. To determine the sign of $F_{2}^{\prime}(l)$, set
$h(t):=\psi(t)-\psi\!\left(t+\frac{d}{2}\right),\qquad t>0.$
Since $\psi$ is concave, $\psi^{\prime}$ is decreasing, hence $h^{\prime}(t)=\psi^{\prime}(t)-\psi^{\prime}(t+\tfrac{d}{2})\geq 0$, so $h$ is non-decreasing. Moreover $h(t)<0$ for all $t>0$ since $\psi$ is increasing. Because $u<u+l$, it follows that
$h(u)\leq h(u+l)<0.$
Since $x^{\prime}(l)-\tfrac{1}{2}<0$, replacing $h(u)$ by the larger value $h(u+l)$ in the first term of (34) gives the upper bound
$F_{2}^{\prime}(l)\leq\Bigl(x^{\prime}(l)-\tfrac{1}{2}\Bigr)h(u+l)+\Bigl(x^{\prime}(l)+\tfrac{1}{2}\Bigr)h(u+l)=2x^{\prime}(l)\,h(u+l)\leq 0.$
Therefore $F_{2}$ is non-increasing on $(\tau,\infty)$ and $\sup_{l>\tau}A_{\kappa,\mu_{2}(l),l,a,d}=\lim_{l\downarrow\tau}A_{\kappa,\mu_{2}(l),l,a,d}.$ Since $\mu_{2}(l)\to 1$ as $l\downarrow\tau$ and $\mu_{1}(\tau)=1$, by continuity of $LA_{\kappa,d}(\mu,l)$ on the admissible region,
$\lim_{l\downarrow\tau}LA_{\kappa,d}(\mu_{2}(l),l)=LA_{\kappa,d}(1,\tau)=LA_{\kappa,d}(\mu_{1}(\tau),\tau).$
Therefore the maximum of $A_{\kappa,\mu_{1}(l),l,a,d}$ over $0\leq l\leq\tau$ is attained at $l=\tau$, and the supremum of $A_{\kappa,\mu_{2}(l),l,a,d}$ over $l>\tau$ equals the same value and is approached as $l\downarrow\tau$. Consequently, the global maximum over the admissible region is achieved at $(\mu,l)=(1,\tau)=(1,d/2+\kappa)$, while in the regime $l>\tau$ it is not attained but only approached in the limit $l\downarrow\tau$. Restoring the factor $a^{d}$ yields the statement. ∎
A.6 Proof of Theorem 6
Proof.
Since all models are compactly supported and hence absolutely integrable, it suffices to verify the identity at the level of spectral densities. Taking the Fourier transform of both sides of (18) and using the scaling identity $\widehat{\mathcal{H}}_{\kappa,1,a\sqrt{uv},d}(z)=(a\sqrt{uv})^{d}\,\widehat{\mathcal{H}}_{\kappa,1,1,d}(za\sqrt{uv})$, the right-hand side of (18) has spectral density
$I(z)=C\,a^{d}\int_{0}^{1}\!\!\int_{0}^{1}u^{A}(1-u)^{B}\,v^{P}(1-v)^{Q}\,(uv)^{d/2}\,\widehat{\mathcal{H}}_{\kappa,1,1,d}(za\sqrt{uv})\,\mathrm{d}u\,\mathrm{d}v,$ (35)
where, from the exponents in (18),
$A=\delta-\tfrac{d+1}{2},\quad B=\tfrac{\mu-3}{2},\quad P=\delta-\tfrac{1}{2}+\kappa,\quad Q=\tfrac{\mu-d-3}{2}-\kappa+l.$
From (7), the spectral density of $\mathcal{H}_{\kappa,1,1,d}$ is
$\widehat{\mathcal{H}}_{\kappa,1,1,d}(w)=K\,{}_{1}F_{2}\!\left(\delta;\,\delta+\tfrac{1}{2},\,2\delta;\,-\tfrac{w^{2}}{4}\right)=K\sum_{k=0}^{\infty}\frac{(\delta)_{k}}{(\delta+\frac{1}{2})_{k}\,(2\delta)_{k}}\frac{(-w^{2}/4)^{k}}{k!},$
where $K>0$ is a constant depending only on $(\kappa,d)$. Substituting $w=za\sqrt{uv}$ and inserting into (35):
$I(z)=C\,K\,a^{d}\sum_{k=0}^{\infty}\frac{(\delta)_{k}}{(\delta+\frac{1}{2})_{k}\,(2\delta)_{k}}\frac{(-z^{2}a^{2}/4)^{k}}{k!}\;\mathcal{B}_{u}(k)\;\mathcal{B}_{v}(k),$
where the two integrals separate as
$\mathcal{B}_{u}(k):=\int_{0}^{1}u^{A+d/2+k}(1-u)^{B}\,\mathrm{d}u,\qquad\mathcal{B}_{v}(k):=\int_{0}^{1}v^{P+d/2+k}(1-v)^{Q}\,\mathrm{d}v.$
By the standard identity $\int_{0}^{1}t^{\alpha+k-1}(1-t)^{\beta-1}\,\mathrm{d}t=B(\alpha,\beta)\,(\alpha)_{k}/(\alpha+\beta)_{k}$, one obtains
$\mathcal{B}_{u}(k)$ $=B\!\left(A+\tfrac{d}{2}+1,\,B+1\right)\cdot\frac{(A+\frac{d}{2}+1)_{k}}{(A+B+\frac{d}{2}+2)_{k}},$ (36)
$\mathcal{B}_{v}(k)$ $=B\!\left(P+\tfrac{d}{2}+1,\,Q+1\right)\cdot\frac{(P+\frac{d}{2}+1)_{k}}{(P+Q+\frac{d}{2}+2)_{k}}.$ (37)
Substituting $A,B,P,Q$ and $\delta=\frac{d+1}{2}+\kappa$:
$A+\tfrac{d}{2}+1=\delta-\tfrac{d+1}{2}+\tfrac{d}{2}+1=\delta+\tfrac{1}{2},\quad A+B+\tfrac{d}{2}+2=\delta-\tfrac{d+1}{2}+\tfrac{\mu-3}{2}+\tfrac{d}{2}+2=\delta+\tfrac{\mu}{2}=\beta,$
$P+\tfrac{d}{2}+1=\delta-\tfrac{1}{2}+\kappa+\tfrac{d}{2}+1=2\delta,\quad P+Q+\tfrac{d}{2}+2=\delta-\tfrac{1}{2}+\kappa+\tfrac{\mu-d-3}{2}-\kappa+l+\tfrac{d}{2}+2=\delta+\tfrac{\mu}{2}+l=\gamma.$
Hence (36) and (37) become
$\mathcal{B}_{u}(k)=B\!\left(\delta+\tfrac{1}{2},\,\tfrac{\mu-1}{2}\right)\frac{(\delta+\frac{1}{2})_{k}}{(\beta)_{k}},\qquad\mathcal{B}_{v}(k)=B\!\left(2\delta,\,\gamma-2\delta\right)\frac{(2\delta)_{k}}{(\gamma)_{k}}.$
Inserting back into the power series, the factors $(\delta+\tfrac{1}{2})_{k}$ and $(2\delta)_{k}$ cancel with those in the denominator of the $\mathcal{H}_{\kappa,1}$ expansion, leaving
$I(z)=C^{\prime}\,a^{d}\sum_{k=0}^{\infty}\frac{(\delta)_{k}}{(\beta)_{k}\,(\gamma)_{k}}\frac{(-z^{2}a^{2}/4)^{k}}{k!}=C^{\prime}\,\widehat{\mathcal{GH}}_{\delta,\beta,\gamma,a}(z),$
where $C^{\prime}=C\cdot K\cdot B(\delta+\frac{1}{2},\frac{\mu-1}{2})\cdot B(2\delta,\gamma-2\delta)$. Setting $C$ to be the reciprocal of this product of Beta constants establishes the identity.
The integrability of the kernel (i.e., that the exponents $B=(\mu-3)/2>-1$ and $Q=(\mu-d-3)/2-\kappa+l>-1$) is guaranteed respectively by $\mu>1$ and $\mu+2l>1+d+2\kappa$, which are precisely the hypotheses of the theorem. ∎
A.7 Proof of Theorem 7
Proof.
If $d=1$ and $\kappa=k$ is a nonnegative integer, using formula 7.3.1.70 of Prudnikov et al. (1990) and applying the identity
$P_{a}^{b}(x)=P^{b}_{-a-1}(x),$
where $P_{a}^{b}(x)$ denotes the associated Legendre function (of the first kind) with degree $a$ and order $b$ (see Abramowitz and Stegun (1970, 8.3.1)), one can rewrite (17) as
$\mathcal{H}_{k,\mu,a,1}(x)$ $=$ $\frac{\Gamma(\frac{\mu}{2}+2k+1)}{\Gamma(k+\frac{1}{2})}\sqrt{\pi}\left(1-\frac{x^{2}}{a^{2}}\right)^{\frac{\mu}{2}+k}\left(\frac{x}{a}\right)^{k}P^{-\frac{\mu}{2}-k}_{k}\left(\frac{a^{2}+x^{2}}{2ax}\right).$
The claim of the Theorem then derives from the following result, which is valid for $a$ being a nonnegative integer and $z>1$:
$P^{-b}_{a}(z)=\frac{(z-1)^{b/2}}{(z+1)^{b/2}}\sum_{j=0}^{a}\frac{(j+a)!(0.5(z-1))^{j}}{j!\Gamma(j+b+1)(a-j)!}.$ (39)
A.8 Proof of Theorem 8
Proof.
For $\mu\geq 1$ and $\kappa=0$, applying formula 7.3.1.41 from Prudnikov et al. (1990) to Equation (17), one gets
$\mathcal{H}_{0,\mu,a,d}(x)=\frac{\Gamma(\frac{1+\mu}{2})\Gamma(\frac{1+\mu+d}{2})}{\sqrt{\pi}}2^{\mu+\frac{d-1}{2}}\left(1-\frac{x^{2}}{a^{2}}\right)_{+}^{\frac{\mu}{2}+\frac{d-1}{4}}\,P_{\frac{d-1}{2}}^{\frac{1-d}{2}-\mu}\left(\frac{x}{a}\right).$
The claim of the Theorem stems from the fact that, for $m$ being an integer, the associated Legendre function of degree $m$ is a terminating series:
$P_{m}^{\eta}(z)=\left(\frac{1+z}{1-z}\right)^{\frac{\eta}{2}}\sum\limits_{n=0}^{m}\frac{(-m)_{n}(m+1)_{n}}{n!\Gamma(n-\eta+1)}\left(\frac{1-z}{2}\right)^{n},\quad 0<z<1.$ (40)
A.9 Proof of Theorem 9
Proof.
For notational convenience, we use the representation ${\cal GH}_{\delta,\beta,\gamma,a}$ for the model $\mathcal{H}_{\kappa,\mu,a,d}$. Recall (see 16) that if
$\delta=\frac{d+1}{2}+\kappa,\qquad\beta=\delta+\frac{\mu}{2},\qquad\gamma=\delta+\frac{\mu}{2}+\frac{d}{2}+\kappa,$ (41)
then
$\mathcal{H}_{\kappa,\mu,a,d}=\mathcal{GH}_{\delta,\beta,\gamma,a}.$
To prove Theorem 9, we proceed as in the proof of Theorem 1. Using (30), we seek conditions such that, for some positive and finite $c$,
$\int_{c}^{\infty}z^{d-1}\left(\frac{\sigma_{1}^{2}\widehat{{\cal GH}}_{\delta,\beta,\gamma,a}(z)-\sigma_{0}^{2}\widehat{{\cal MT}}_{\nu,\alpha}(z)}{\sigma_{0}^{2}\widehat{{\cal MT}}_{\nu,\alpha}(z)}\right)^{2}{\rm d}z<\infty.$ (42)
It is known that $\sigma_{0}^{2}\widehat{{\cal MT}}_{\nu,\alpha}(z)z^{\tau}$ is bounded away from 0 and $\infty$ as $z\to\infty$ for some $\tau>0$ (Zhang, 2004). Specifically, from (3), as $z\to\infty$,
$\widehat{\mathcal{MT}}_{\nu,\alpha}(z)\sim K_{\mathrm{MT}}\,z^{-2\nu-d},\qquad K_{\mathrm{MT}}=\frac{\Gamma(\nu+d/2)}{\pi^{d/2}\Gamma(\nu)}\,\alpha^{-2\nu},$
so that $\tau=2\nu+d$.
Let $\lambda=\frac{1}{2}+\delta-\beta-\gamma.$ Under parameterization (41),
$\lambda$ $=\frac{1}{2}+\delta-\!\left(\delta+\frac{\mu}{2}\right)-\!\left(\delta+\frac{\mu}{2}+\frac{d}{2}+\kappa\right)$
$=\frac{1}{2}-\delta-\mu-\frac{d}{2}-\kappa$
$=-d-2\kappa-\mu\;<\;0,$
and also
$\beta+\gamma\;=\;2\delta+\mu+\frac{d}{2}+\kappa\;=\;3\delta+\mu-\frac{1}{2},$
so that $\beta+\gamma\geq 3\delta+\frac{1}{2}$ holds whenever $\mu\geq 1$, which is exactly the validity condition of the $\mathcal{H}$ model. Hence Lemma 1 applies.
Using Point 1 of Lemma 1, as $z\to\infty$,
$\left|\frac{\sigma_{1}^{2}\widehat{\mathcal{GH}}_{\delta,\beta,\gamma,a}(z)-\sigma_{0}^{2}\widehat{\mathcal{MT}}_{\nu,\alpha}(z)}{\sigma_{0}^{2}\widehat{\mathcal{MT}}_{\nu,\alpha}(z)}\right|$ $=\Bigg|C\Big[c_{3}(za)^{-2\delta}\{1+\mathcal{O}(z^{-2})\}$
$\qquad+c_{4}(za)^{\lambda}\{\cos(za+c_{5})+\mathcal{O}(z^{-1})\}\Big]$
$\qquad\times z^{2\nu+d}\Big[1+(\nu+d/2)(\alpha z)^{-2}+\mathcal{O}(z^{-2})\Big]-1\Bigg|$
$=\Bigg|w_{1}z^{2\nu+d-2\delta}\{1+\mathcal{O}(z^{-2})\}$
$\qquad\times\Big[1+(\nu+d/2)(\alpha z)^{-2}+\mathcal{O}(z^{-2})\Big]$
$\qquad+w_{2}z^{\lambda+2\nu+d}\Big[1+(\nu+d/2)(\alpha z)^{-2}+\mathcal{O}(z^{-2})\Big]$
$\qquad\times\{\cos(az-c_{5})+\mathcal{O}(z^{-1})\}-1\Bigg|,$
where
$C=\frac{\sigma_{1}^{2}Ka^{d}\Gamma(\nu)\pi^{d/2}}{\Gamma(\nu+d/2)\sigma_{0}^{2}\alpha^{-2\nu}},\qquad w_{1}=\frac{\sigma_{1}^{2}Ka^{d-2\delta}\Gamma(\nu)c_{3}}{\Gamma(\nu+d/2)\sigma_{0}^{2}\alpha^{-2\nu}\pi^{-d/2}},\qquad w_{2}=\frac{w_{1}c_{4}a^{\lambda}}{c_{3}}.$
Since
$z^{2\nu+d}\Big[(\nu+d/2)(\alpha z)^{-2}+\mathcal{O}(z^{-2})\Big]=\mathcal{O}(z^{2\nu+d-2}),$
we obtain
$\int_{c}^{\infty}z^{d-1}\Bigg|\frac{\sigma_{1}^{2}\widehat{{\cal GH}}_{\delta,\beta,\gamma,a}(z)-\sigma_{0}^{2}\widehat{{\cal MT}}_{\nu,\alpha}(z)}{\sigma_{0}^{2}\widehat{{\cal MT}}_{\nu,\alpha}(z)}\Bigg|^{2}{\rm d}z$ $=\int_{c}^{\infty}z^{d-1}\Big|T_{1}(z)+T_{2}(z)+T_{3}(z)\Big|^{2}{\rm d}z,$
where
$T_{1}(z)=w_{1}z^{-2\delta}\mathcal{O}(z^{2\nu+d-2}),\qquad T_{2}(z)=w_{1}z^{2\nu+d-2\delta}-1,$
$T_{3}(z)=w_{2}z^{\lambda}\{\mathcal{O}(z^{2\nu+d-2})+z^{2\nu+d}\}\{\cos(az-c_{5})+\mathcal{O}(z^{-1})\}.$
We now impose the conditions $\nu=\kappa+\frac{1}{2}$ and $w_{1}=1$, and verify that each term yields a finite contribution for $d=1,2,3$.
Condition $\nu=\kappa+\frac{1}{2}$ implies $2\nu+d=2\delta$. Since $\delta=\frac{d+1}{2}+\kappa$, we have $2\delta=d+1+2\kappa=2\nu+d$. Hence $2\nu+d-2\delta=0$.
Under $w_{1}=1$ and $2\nu+d=2\delta$, we have $w_{1}z^{2\nu+d-2\delta}=1$, so $T_{2}(z)\equiv 0$. With $2\nu+d=2\delta$,
$T_{1}(z)=w_{1}\,z^{-2\delta}\,\mathcal{O}(z^{2\nu+d-2})=\mathcal{O}(z^{-2}).$
Therefore
$\int_{c}^{\infty}z^{d-1}T_{1}(z)^{2}\,\mathrm{d}z=\mathcal{O}\!\left(\int_{c}^{\infty}z^{d-5}\,\mathrm{d}z\right),$
which converges for d5<1d-5<-1, i.e. $d<4$. Hence the integral is finite for $d=1,2,3$.
For $T_{3}$, since $\lambda<0$, the dominant contribution comes from the $z^{2\nu+d}$ factor:
$T_{3}(z)=w_{2}\,z^{\lambda+2\nu+d}\{\cos(az-c_{5})+\mathcal{O}(z^{-1})\}+\mathcal{O}(z^{\lambda+2\nu+d-2}).$
Under $2\nu+d=2\delta$,
$\lambda+2\nu+d\;=\;\lambda+2\delta\;=\;(-d-2\kappa-\mu)+(d+1+2\kappa)\;=\;1-\mu.$
Since $\mu\geq 1+\frac{d}{2}$ (as required by Theorems 10), we have $\lambda+2\delta=1-\mu<0$. By the Riemann–Lebesgue argument for oscillatory integrals,
$\int_{c}^{\infty}z^{d-1}T_{3}(z)^{2}\,\mathrm{d}z=\mathcal{O}\!\left(\int_{c}^{\infty}z^{d-1+2(1-\mu)}\,\mathrm{d}z\right),$
which converges provided $d-1+2(1-\mu)<-1$, i.e. $\mu>\frac{d}{2}+1$, which holds under the stated assumption.
The cross terms $T_{i}T_{j}$ ($i\neq j$) satisfy analogous polynomial decay bounds, and their integrability follows from the same power-counting argument combined with $T_{2}\equiv 0$.
Combining the above, condition (42) holds whenever $\nu=\kappa+\frac{1}{2}$, $w_{1}=1$, $\mu\geq 1+\frac{d}{2}$, and $d=1,2,3$. Hence
$P(\sigma_{0}^{2}{\cal MT}_{\nu,\alpha})\equiv P(\sigma_{1}^{2}{\cal GH}_{\delta,\beta,\gamma,a}).$
Condition $w_{1}=1$ is equivalent to
$\frac{\sigma_{0}^{2}}{\alpha^{2\nu}}=\frac{\sigma_{1}^{2}A}{a^{2\delta-d}},$
where
$A=\frac{2^{2\delta-d}\Gamma(\beta-d/2)\Gamma(\gamma-d/2)}{\Gamma(\gamma-\delta)\Gamma(\beta-\delta)}.$
Using parameterization (41), this becomes
$\frac{\sigma_{0}^{2}}{\alpha^{2\nu}}=\frac{\sigma_{1}^{2}}{a^{2\kappa+1}}F(\kappa,\mu,d),$
where $F(\kappa,\mu,d)$ is defined in (20). Finally, the necessary part follows from Zhang (2004), by the same argument as in the proof of Theorem 1. ∎
A.10 Proof of Theorem 11
A preliminary lemma is required for the proof of Theorem 11. It establishes a uniform stochastic control of the plug-in microergodic estimator $\hat{\theta}_{n}(\mu,a)$ over $a\in I$, which is the key ingredient for showing that replacing a fixed $a$ by its ML estimator does not affect the fixed-domain limit distribution in Theorem 11.
Lemma 2.
Assume the setting of Theorem 10 with fixed $\mu>1+\frac{d}{2}$ and $a_{0}\in I$. Let
$\hat{\theta}_{n}(\mu,a)=\frac{\hat{\sigma}_{n}^{2}(\mu,a)\,F(\kappa,\mu,d)}{a^{2\kappa+1}},\qquad\hat{\sigma}_{n}^{2}(\mu,a)=\frac{\mathbf{Z}_{n}^{\top}R_{n}(\mu,a)^{-1}\mathbf{Z}_{n}}{n},$
and $\theta_{0}=\sigma_{0}^{2}F(\kappa,\mu,d)/a_{0}^{2\kappa+1}$. Then, as $n\to\infty$,
  1. (i)
    $\sup_{a\in I}\bigl|\hat{\theta}_{n}(\mu,a)-\theta_{0}\bigr|\xrightarrow{p}0$;
  2. (ii)
    for any fixed $a^{\star}\in I$, $\sup_{a\in I}\sqrt{n}\,\bigl|\hat{\theta}_{n}(\mu,a)-\hat{\theta}_{n}(\mu,a^{\star})\bigr|=o_{p}(1)$.
Proof.
Fix $\mu>1+\frac{d}{2}$ and $I=[a_{L},a_{U}]$. For $a\in I$ set
$G_{n}(a):=\frac{a^{2\kappa+1}}{F(\kappa,\mu,d)}\,R_{n}(\mu,a),\qquad\hat{\theta}_{n}(\mu,a)=\frac{1}{n}\,\mathbf{Z}_{n}^{\top}G_{n}(a)^{-1}\mathbf{Z}_{n}.$
Let $G_{0n}:=G_{n}(a_{0})$ and define the vector
$\mathbf{Y}_{n}:=\theta_{0}^{-1/2}\,G_{0n}^{-1/2}\mathbf{Z}_{n}\sim\mathcal{N}(0,I_{n}),\qquad\theta_{0}=\sigma_{0}^{2}F(\kappa,\mu,d)/a_{0}^{2\kappa+1}.$
Then
$\frac{\hat{\theta}_{n}(\mu,a)}{\theta_{0}}=\frac{1}{n}\,\mathbf{Y}_{n}^{\top}M_{n}(a)\mathbf{Y}_{n},\qquad M_{n}(a):=G_{0n}^{1/2}G_{n}(a)^{-1}G_{0n}^{1/2}.$ (43)
Fix $a^{\star}\in I$ and set $\Delta_{n}(a):=M_{n}(a)-M_{n}(a^{\star})$. The decomposition
$\frac{\sqrt{n}\,\bigl(\hat{\theta}_{n}(\mu,a)-\hat{\theta}_{n}(\mu,a^{\star})\bigr)}{\theta_{0}}=\underbrace{\frac{1}{\sqrt{n}}\operatorname{tr}(\Delta_{n}(a))}_{\mathrm{(I)}}+\underbrace{\frac{1}{\sqrt{n}}\bigl(\mathbf{Y}_{n}^{\top}\Delta_{n}(a)\mathbf{Y}_{n}-\operatorname{tr}(\Delta_{n}(a))\bigr)}_{\mathrm{(II)}}$ (44)
will be used to prove (ii), and then (i) follows from (ii) by taking $a^{\star}=a_{0}$.
Hereafter, we use $\|A\|_{\mathrm{HS}}^{2}:=\mathrm{tr}(A^{*}A)$ for Hilbert–Schmidt operators $A$, where $A^{*}$ denotes the adjoint operator. For matrices $B$, we use $\|B\|_{F}^{2}:=\mathrm{tr}(B^{\top}B)$.
Let $g_{a}$ denote the isotropic spectral density of the stationary covariance function $G(\cdot\,;a):=\frac{a^{2\kappa+1}}{F(\kappa,\mu,d)}\mathcal{H}_{\kappa,\mu,a,d}(\cdot)$. By Lemma 1 specialised to the $\mathcal{H}$ model, the high-frequency expansion of $g_{a}$ has the form
$g_{a}(z)=C_{\kappa,\mu,d}\,z^{-2\delta}\bigl(1+\mathcal{O}(z^{-2})\bigr)+\mathcal{O}(z^{1-\mu-2\delta}),\qquad z\to\infty,$
with $\delta=\frac{d+1}{2}+\kappa$ and with $\mathcal{O}$-constants uniform over $a\in I$ (the leading term is independent of $a$ because of the prefactor $a^{2\kappa+1}$). Hence, for any fixed $c>0$,
$\sup_{a\in I}\int_{c}^{\infty}z^{d-1}\Bigl(\frac{g_{a}(z)-g_{a_{0}}(z)}{g_{a_{0}}(z)}\Bigr)^{2}\,dz<\infty,$ (45)
since $\mu>1+d/2$ and $d\in\{1,2,3\}$ make the tail square-integrable. By the fixed-domain equivalence criterion of Stein (2004) and its connection to covariance-ratio operators (see Skorokhod and Yadrenko (1973, Assertion B)), (45) implies that the operator
$\mathcal{M}(a):=\mathcal{G}(a_{0})^{1/2}\mathcal{G}(a)^{-1}\mathcal{G}(a_{0})^{1/2}-\mathcal{I}\quad\text{on }L^{2}(D)$
is Hilbert–Schmidt, with $\sup_{a\in I}\|\mathcal{M}(a)\|_{\mathrm{HS}}<\infty$. Let $\Pi_{n}$ denote the orthogonal projector onto the finite-dimensional subspace generated by the observations $\{Z(\mathbf{s}_{1}),\ldots,Z(\mathbf{s}_{n})\}$ (identified with its image in $L^{2}(D)$ through the covariance-operator embedding). Then $M_{n}(a)-I_{n}$ represents the restriction of $\Pi_{n}\mathcal{M}(a)\Pi_{n}$ to that subspace, hence
$\sup_{a\in I}\|M_{n}(a)-I_{n}\|_{F}^{2}=\sup_{a\in I}\|\Pi_{n}\mathcal{M}(a)\Pi_{n}\|_{\mathrm{HS}}^{2}\leq\sup_{a\in I}\|\mathcal{M}(a)\|_{\mathrm{HS}}^{2}=:C_{\mathrm{HS}}<\infty,\qquad\forall n.$ (46)
In particular,
$\sup_{a\in I}\|M_{n}(a)\|$ $\leq 1+\sup_{a\in I}\|M_{n}(a)-I_{n}\|\leq 1+\sup_{a\in I}\|M_{n}(a)-I_{n}\|_{F}$ (47)
$\leq 1+\sqrt{C_{\mathrm{HS}}},\qquad\sup_{a\in I}\|\Delta_{n}(a)\|_{F}^{2}\leq 4C_{\mathrm{HS}},\qquad\forall n.$
Let $p:=\mu+\frac{d-1}{2}+2\kappa$ and $\alpha:=\min(1,p)$. Since $\mu>1+\frac{d}{2}$ and $\kappa>-1/2$, one has $\alpha>1/2$. From the explicit form of $\mathcal{H}_{\kappa,\mu,a,d}$ and its compact support in $[0,a]$, the map $a\mapsto\mathcal{H}_{\kappa,\mu,a,d}(r)$ is uniformly $\alpha$-Hölder on $I$, namely
$\sup_{r\geq 0}\bigl|\mathcal{H}_{\kappa,\mu,a,d}(r)-\mathcal{H}_{\kappa,\mu,a^{\prime},d}(r)\bigr|\leq C_{\alpha}|a-a^{\prime}|^{\alpha},\qquad a,a^{\prime}\in I,$ (48)
where the exponent $\alpha=\min(1,p)$ reflects the boundary behaviour at $r=a$. Moreover, using (48), continuity in $(a,z)$ on compact sets, positivity of $g_{a_{0}}$, and the common high-frequency expansion above, one obtains that for any fixed $c>0$,
$\int_{c}^{\infty}z^{d-1}\Bigl(\frac{g_{a}(z)-g_{a^{\prime}}(z)}{g_{a_{0}}(z)}\Bigr)^{2}dz\leq C|a-a^{\prime}|^{2\alpha},\qquad a,a^{\prime}\in I,$
with $C$ independent of $(a,a^{\prime})$. Consequently, by the same covariance-ratio/operator correspondence (Stein, 2004, Appendix A) (see also Skorokhod and Yadrenko (1973, Assertion B)),
$\|\mathcal{M}(a)-\mathcal{M}(a^{\prime})\|_{\mathrm{HS}}\leq C|a-a^{\prime}|^{\alpha},\qquad a,a^{\prime}\in I.$
Passing to the discretised matrices and using again that orthogonal projections do not increase HS norms yields the Frobenius increment bound
$\|\Delta_{n}(a)-\Delta_{n}(a^{\prime})\|_{F}\leq C|a-a^{\prime}|^{\alpha},\qquad a,a^{\prime}\in I,\ \forall n.$ (49)
By Hanson–Wright (Hanson and Wright, 1971; Rudelson and Vershynin, 2013), for any symmetric $A$ and $\mathbf{Y}\sim\mathcal{N}(0,I_{n})$,
$P\!\Bigl(\bigl|\mathbf{Y}^{\top}A\mathbf{Y}-\operatorname{tr}(A)\bigr|>t\Bigr)\leq 2\exp\!\Bigl(-c\min\!\Bigl\{\frac{t^{2}}{\|A\|_{F}^{2}},\frac{t}{\|A\|}\Bigr\}\Bigr).$
With $A=\Delta_{n}(a)$ and $t=\varepsilon\sqrt{n}$, using (47) and $\|A\|\leq\|A\|_{F}$,
$P\!\Bigl(\Bigl|\frac{1}{\sqrt{n}}\bigl(\mathbf{Y}_{n}^{\top}\Delta_{n}(a)\mathbf{Y}_{n}-\mathrm{tr}\Delta_{n}(a)\bigr)\Bigr|>\varepsilon\Bigr)\leq 2e^{-c^{\prime}n},$
uniformly in $a\in I$. Let $\delta_{n}:=n^{-4}$ and let $\mathcal{G}_{n}\subset I$ be a regular grid of mesh $\delta_{n}$ (so $|\mathcal{G}_{n}|=\mathcal{O}(n^{4})$). A union bound gives
$\sup_{a\in\mathcal{G}_{n}}\Bigl|\frac{1}{\sqrt{n}}\bigl(\mathbf{Y}_{n}^{\top}\Delta_{n}(a)\mathbf{Y}_{n}-\mathrm{tr}\Delta_{n}(a)\bigr)\Bigr|=o_{p}(1).$ (50)
For arbitrary $a\in I$, choose $a^{(n)}\in\mathcal{G}_{n}$ with $|a-a^{(n)}|\leq\delta_{n}$. Then by (49) and $\|\mathbf{Y}_{n}\|^{2}=O_{p}(n)$,
$\frac{1}{\sqrt{n}}\bigl|\mathbf{Y}_{n}^{\top}(\Delta_{n}(a)-\Delta_{n}(a^{(n)}))\mathbf{Y}_{n}\bigr|$ $\leq\frac{\|\Delta_{n}(a)-\Delta_{n}(a^{(n)})\|_{F}}{\sqrt{n}}\,\|\mathbf{Y}_{n}\|^{2}$
$\leq C\delta_{n}^{\alpha}\,O_{p}(\sqrt{n})=O_{p}(n^{1/2-4\alpha})=o_{p}(1).$
since $\alpha>1/2$. Combining with (50) yields
$\sup_{a\in I}\Bigl|\frac{1}{\sqrt{n}}\bigl(\mathbf{Y}_{n}^{\top}\Delta_{n}(a)\mathbf{Y}_{n}-\mathrm{tr}\Delta_{n}(a)\bigr)\Bigr|=o_{p}(1).$ (51)
Now, define
$r_{n}(a):=\frac{1}{n}\mathrm{tr}(M_{n}(a))-1=\frac{E[\hat{\theta}_{n}(\mu,a)]}{\theta_{0}}-1.$
Then $\frac{1}{\sqrt{n}}\mathrm{tr}(\Delta_{n}(a))=\sqrt{n}\bigl(r_{n}(a)-r_{n}(a^{\star})\bigr)$. First, by (49) and Cauchy–Schwarz,
$\sqrt{n}\,|r_{n}(a)-r_{n}(a^{\prime})|=\frac{|\mathrm{tr}(M_{n}(a)-M_{n}(a^{\prime}))|}{\sqrt{n}}\leq\|M_{n}(a)-M_{n}(a^{\prime})\|_{F}\leq C|a-a^{\prime}|^{\alpha},$
uniformly in $n$, so $\{\sqrt{n}\,r_{n}(\cdot)\}_{n\geq 1}$ is equicontinuous on $I$. Second, for each fixed $a$, Theorem 10(ii) gives $\sqrt{n}(\hat{\theta}_{n}(\mu,a)-\theta_{0})\Rightarrow\mathcal{N}(0,2\theta_{0}^{2})$. Moreover, for Gaussian quadratic forms,
$\rm Var\!\Big(\mathbf{Y}_{n}^{\top}M_{n}(a)\mathbf{Y}_{n}-\operatorname{tr}(M_{n}(a))\Big)=2\,\operatorname{tr}(M_{n}(a)^{2})\leq 2\,\|M_{n}(a)\|_{F}^{2},$
so (46) yields a uniform second-moment bound and hence uniform integrability. Therefore
[…]
By equicontinuity and compactness of $I$ (finite-cover argument),
$\sup_{a\in I}\frac{1}{\sqrt{n}}|\mathrm{tr}(\Delta_{n}(a))|=\sup_{a\in I}\sqrt{n}\,|r_{n}(a)-r_{n}(a^{\star})|\leq 2\sup_{a\in I}\sqrt{n}\,|r_{n}(a)|\longrightarrow 0.$ (52)
Finally, combining (44), (51), and (52) gives
$\sup_{a\in I}\sqrt{n}\,\bigl|\hat{\theta}_{n}(\mu,a)-\hat{\theta}_{n}(\mu,a^{\star})\bigr|=o_{p}(1),$
which proves (ii).
To show (i), take $a^{\star}=a_{0}$ in (ii) to get
$\sup_{a\in I}|\hat{\theta}_{n}(\mu,a)-\theta_{0}|\leq\sup_{a\in I}|\hat{\theta}_{n}(\mu,a)-\hat{\theta}_{n}(\mu,a_{0})|+|\hat{\theta}_{n}(\mu,a_{0})-\theta_{0}|.$
The first term is $o_{p}(1)$ by (ii), and the second converges to 0 almost surely by Theorem 10(i). Hence $\sup_{a\in I}|\hat{\theta}_{n}(\mu,a)-\theta_{0}|\to 0$ in probability, proving (i). ∎
A.11 Proof of Theorem 12
The proof of Theorem 12 requires a uniform stochastic control of the plug-in microergodic estimator over the compact parameter set $J\times I$, since the ML estimators $(\hat{\mu}_{n},\hat{a}_{n})$ are random. The next lemma extends Lemma 2 to the case where $\mu$ is unknown and varies in $J$.
Lemma 3.
Assume the setting of Theorem 10, with $(\sigma_{0}^{2},\mu_{0},a_{0})\in(0,\infty)\times J\times I$ and $\mu_{0}>1+\frac{d}{2}$, where $J=[\mu_{L},\mu_{U}]$ and $I=[a_{L},a_{U}]$ with $\mu_{L}>1+\frac{d}{2}$. Let
$\hat{\theta}_{n}(\mu,a)=\frac{\hat{\sigma}_{n}^{2}(\mu,a)\,F(\kappa,\mu,d)}{a^{2\kappa+1}},\qquad\hat{\sigma}_{n}^{2}(\mu,a)=\frac{\mathbf{Z}_{n}^{\top}R_{n}(\mu,a)^{-1}\mathbf{Z}_{n}}{n},$
and $\theta_{0}=\sigma_{0}^{2}F(\kappa,\mu_{0},d)/a_{0}^{2\kappa+1}$. Then, as $n\to\infty$,
  1. (i)
    $\sup_{(\mu,a)\in J\times I}\big|\hat{\theta}_{n}(\mu,a)-\theta_{0}\big|\xrightarrow{p}0;$
  2. (ii)
    for any fixed $(\mu^{\star},a^{\star})\in J\times I$,
    $\sup_{(\mu,a)\in J\times I}\sqrt{n}\,\big|\hat{\theta}_{n}(\mu,a)-\hat{\theta}_{n}(\mu^{\star},a^{\star})\big|=o_{p}(1).$
Proof.
The argument follows closely the proof of Lemma 2. We keep the same notation and decomposition, replacing $a\in I$ by $(\mu,a)\in J\times I$, and we record only the modifications needed to obtain uniformity over the two-parameter compact set.
All bounds that in Lemma 2 were uniform over $a\in I$ remain uniform over $(\mu,a)\in J\times I$, because $J\times I$ is compact and $\mu_{L}>1+\frac{d}{2}$. In particular, the high-frequency expansion of the normalised spectral density and Stein’s integrability condition hold uniformly on $J\times I$, yielding uniform Hilbert–Schmidt bounds for the covariance-ratio operators and hence
$\sup_{(\mu,a)\in J\times I}\|M_{n}(\mu,a)-I_{n}\|_{F}^{2}\leq C,\qquad\sup_{(\mu,a)\in J\times I}\|M_{n}(\mu,a)\|\leq 1+\sqrt{C},$
for some finite constant $C$ and all $n$, where $M_{n}(\mu,a)$ is defined as in Lemma 2 with $G_{n}(\mu,a):=\frac{a^{2\kappa+1}}{F(\kappa,\mu,d)}\,R_{n}(\mu,a)$. Moreover, on $J\times I$ the map $\mu\mapsto\mathcal{H}_{\kappa,\mu,a,d}(r)$ is $C^{1}$ and thus uniformly Lipschitz in $\mu$, while the map $a\mapsto\mathcal{H}_{\kappa,\mu,a,d}(r)$ is uniformly $\alpha$–Hölder in $a$ with $\alpha>1/2$, implying the two-parameter increment bound
$\|\Delta_{n}(\mu,a)-\Delta_{n}(\mu^{\prime},a^{\prime})\|_{F}\leq C\bigl(|\mu-\mu^{\prime}|+|a-a^{\prime}|^{\alpha}\bigr),\qquad(\mu,a),(\mu^{\prime},a^{\prime})\in J\times I,$
where $\Delta_{n}(\mu,a)$ denotes the corresponding increment matrix.
Finally, the Hanson–Wright and grid argument is applied on a two-dimensional grid of mesh $\delta_{n}=n^{-4}$ in $J\times I$ (of cardinality $\mathcal{O}(n^{8})$), and the remainder of the proof proceeds exactly as in Lemma 2, yielding (ii). Statement (i) then follows by taking $(\mu^{\star},a^{\star})=(\mu_{0},a_{0})$ and using Theorem 10(i). ∎
Proof of Theorem 12 .
By definition, $(\hat{\mu}_{n},\hat{a}_{n})\in J\times I$ for all $n$.
Proof of (i). Since $(\hat{\mu}_{n},\hat{a}_{n})\in J\times I$,
$\big|\hat{\theta}_{n}(\hat{\mu}_{n},\hat{a}_{n})-\theta_{0}\big|\leq\sup_{(\mu,a)\in J\times I}\big|\hat{\theta}_{n}(\mu,a)-\theta_{0}\big|.$
By Lemma 3(i), the right-hand side converges to 0 in probability, hence $\hat{\theta}_{n}(\hat{\mu}_{n},\hat{a}_{n})\xrightarrow{p}\theta_{0}$.
Proof of (ii). Fix the nonrandom point $(\mu^{\star},a^{\star})=(\mu_{0},a_{0})\in J\times I$. Then
$\sqrt{n}\big(\hat{\theta}_{n}(\hat{\mu}_{n},\hat{a}_{n})-\theta_{0}\big)$ $=\sqrt{n}\big(\hat{\theta}_{n}(\mu_{0},a_{0})-\theta_{0}\big)$
$\quad+\sqrt{n}\big(\hat{\theta}_{n}(\hat{\mu}_{n},\hat{a}_{n})-\hat{\theta}_{n}(\mu_{0},a_{0})\big).$
By Lemma 3(ii), the second term is $o_{p}(1)$. By Theorem 10(ii) (applied with $\mu=\mu_{0}$ and $a=a_{0}$ fixed),
$\sqrt{n}\big(\hat{\theta}_{n}(\mu_{0},a_{0})-\theta_{0}\big)\xrightarrow{\mathcal{D}}\mathcal{N}(0,2\theta_{0}^{2}).$
Therefore, Slutsky’s theorem yields
$\sqrt{n}\big(\hat{\theta}_{n}(\hat{\mu}_{n},\hat{a}_{n})-\theta_{0}\big)\xrightarrow{\mathcal{D}}\mathcal{N}(0,2\theta_{0}^{2}),$
which proves (ii). ∎
Appendix B Closed-form expressions of the hypergeometric model
In Euclidean spaces of odd dimension. When $d$ is an odd integer, $\kappa=k$ is a nonnegative integer and $\mu$ is a positive integer, the hypergeometric model yields a truncated polynomial (Emery and Alegría, 2022, Section 2.5.3). Specifically, if $\mu$ is odd, then
$\begin{split}\mathcal{H}_{k,\mu,a,d}(x)=\begin{cases}\frac{\Gamma(\frac{1}{2}-k)\Gamma(\frac{\mu+1}{2}+k)}{\Gamma(\frac{1-\mu-d}{2}-2k)}\bigg[\sum_{n=0}^{k+\frac{\mu-1}{2}}\frac{(-1)^{n}\Gamma(\frac{1-\mu-d}{2}-2k+n)}{\Gamma(\frac{1}{2}-k+n)\Gamma(\frac{\mu+1}{2}+k-n)n!}\left(\frac{x}{a}\right)^{2n}\\ +\sum_{n=0}^{k-1+\frac{\mu+d}{2}}\frac{(-1)^{n}\Gamma(n+1-\frac{\mu}{2})}{\Gamma(\frac{3}{2}+k+n)\Gamma(\frac{\mu+d}{2}+k-n)n!}\left(\frac{x}{a}\right)^{2n+2k+1}\bigg]\text{ if $x<a$}\\ 0\text{ if $x\geq a$,}\end{cases}\end{split}$
while, if $\mu$ is even,
$\begin{split}\mathcal{H}_{k,\mu,a,d}(x)=\begin{cases}\frac{\Gamma(\frac{1}{2}-k)\Gamma(\frac{\mu+d+1}{2}+2k)}{\Gamma(\frac{1-\mu}{2}-k)}\bigg[\sum_{n=0}^{2k+\frac{\mu+d-1}{2}}\frac{(-1)^{n}\Gamma(\frac{1-\mu}{2}-k+n)}{\Gamma(\frac{1}{2}-k+n)\Gamma(\frac{\mu+d+1}{2}+2k-n)n!}\left(\frac{x}{a}\right)^{2n}\\ +\sum_{n=0}^{\frac{\mu}{2}-1}\frac{(-1)^{n}\Gamma(n+1-\frac{\mu+d}{2}-k)}{\Gamma(\frac{3}{2}+k+n)\Gamma(\frac{\mu}{2}-n)n!}\left(\frac{x}{a}\right)^{2n+2k+1}\bigg]\text{ if $x<a$}\\ 0\text{ if $x\geq a$.}\end{cases}\end{split}$
The cases $k=0$, $\mu=1$ and $d=1$, 33 and $5$ correspond to the well-known triangular, spherical and pentaspherical models (Matérn, 1986), which behave linearly near the origin:
$\mathcal{H}_{0,1,a,1}(x)=1-\frac{x}{a},\quad x\leq a,$
$\mathcal{H}_{0,1,a,3}(x)=1-\frac{3x}{2a}+\frac{x^{3}}{2a^{3}},\quad x\leq a,$
$\mathcal{H}_{0,1,a,5}(x)=1-\frac{15x}{8a}+\frac{5x^{3}}{4a^{3}}-\frac{3x^{5}}{8a^{5}},\quad x\leq a,$
whereas the cases $d=3$, $\mu=1$ and $k=1$ or $2$ correspond to the so-called cubic and penta models (Chilès and Delfiner, 2012), which are associated with Gaussian RFs that are once and twice mean-squared differentiable, respectively:
$\mathcal{H}_{1,1,a,3}(x)=1-\frac{7x^{2}}{a^{2}}+\frac{35x^{3}}{4a^{3}}-\frac{7x^{5}}{2a^{5}}+\frac{3x^{7}}{4a^{7}},\quad x\leq a,$
$\mathcal{H}_{2,1,a,3}(x)=1-\frac{22x^{2}}{3a^{2}}+\frac{33x^{4}}{a^{4}}-\frac{77x^{5}}{2a^{5}}+\frac{33x^{7}}{2a^{7}}-\frac{11x^{9}}{2a^{9}}+\frac{5x^{11}}{6a^{11}},\quad x\leq a.$
In the 2D plane. If $d=2$ and $\mu=1$, the following closed-form expressions are found.
  1. 1.
    When $\kappa=0$, formula 7.3.2.92 of Prudnikov et al. (1990) gives:
    $\mathcal{H}_{0,1,a,2}(x)=\frac{2}{\pi}\left(\arcsin\sqrt{\left(1-\frac{x^{2}}{a^{2}}\right)_{+}}-\frac{x}{a}\left(1-\frac{x^{2}}{a^{2}}\right)_{+}^{\frac{1}{2}}\right),$ (53)
    which corresponds to the circular correlation model (Chilès and Delfiner, 2012, 2.46), which behaves linearly near the origin.
  2. 2.
    When $\kappa=\frac{1}{2}$, formula 7.3.2.99 of Prudnikov et al. (1990) gives:
    $\mathcal{H}_{\frac{1}{2},1,a,2}(x)=\left(2-\frac{x^{2}}{a^{2}}\right)\operatorname{arctanh}\sqrt{\left(1-\frac{x^{2}}{a^{2}}\right)_{+}}-\sqrt{\left(1-\frac{x^{2}}{a^{2}}\right)_{+}},$
    which is the montée (upgrading) (Matheron, 1965, I.4.18) of order $1$ of the spherical correlation model in $\mathbb{R}^{3}$.
  3. 3.
    When $\kappa=1$, one obtains (Weisstein, 2025):
    $\mathcal{H}_{1,1,a,2}(x)$ $=$ $\frac{2}{3\pi}\Bigg\{\frac{x}{a}\left(1-\frac{x^{2}}{a^{2}}\right)_{+}^{\frac{1}{2}}\left[15-4\left(1-\frac{x^{2}}{a^{2}}\right)_{+}\left(3-\frac{x^{2}}{a^{2}}\right)\right]$ (54)
    $+3\left[6\left(1-\frac{x^{2}}{a^{2}}\right)_{+}-5\right]\arcsin\sqrt{\left(1-\frac{x^{2}}{a^{2}}\right)_{+}}\Bigg\},$
    which is the montée of order $2$ of the Euclid’s hat model in $\mathbb{R}^{4}$ and is associated with a Gaussian RF that is mean-squared differentiable.